### v1

# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

Abstract
Recent work has made non-interactive privacy-preserving inference more practical by running deep Convolution Neural Network (CNN) with Fully Homomorphic Encryption (FHE). However, these methods remain limited by their reliance on bootstrapping, a costly FHE operation applied across multiple layers, severely slowing inference. They also depend on high-degree polynomial approximations of non-linear activations, which increase multiplicative depth and reduce accuracy by 2–5% compared to plaintext ReLU models. In this work, we focus on ResNets, a widely adopted benchmark architecture in privacy-preserving inference, and close the accuracy gap between their FHE-based non-interactive models and plaintext counterparts, while also achieving faster inference than existing methods. We use a quadratic polynomial approximation of ReLU, which achieves the theoretical minimum multiplicative depth for non-linear activations, along with a penalty-based training strategy. We further introduce structural optimizations such as node fusing, weight redistribution, and tower reuse. These optimizations reduce the required FHE levels in CNNs by nearly a factor of five compared to prior work, allowing us to run ResNet models under leveled FHE without bootstrapping. To further accelerate inference and recover accuracy typically lost with polynomial approximations, we introduce parameter clustering along with a joint strategy of data encoding layout and ensemble techniques. Experiments with ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 show that our approach achieves up to $4\times$ faster private inference than prior work with comparable accuracy to plaintext ReLU models.
1 Introduction
Machine Learning as a Service (MLaaS) is increasingly adopted across industries as it provides access to powerful models without the need for in-house development or maintenance (Ribeiro et al., 2015). However, it raises serious privacy concerns since models are often trained on proprietary or sensitive data (Anwar et al., 2018; Özbayoglu et al., 2020), and inference requires clients to share private information (e.g., medical or financial records) with external service providers. Privacy-Preserving Machine Learning (PPML) addresses these risks using cryptographic methods, mainly secure Multi-Party Computation (MPC) (Zhou et al., 2024) and Fully Homomorphic Encryption (FHE) (Podschwadt et al., 2022), which protect both client data and provider models. Broadly, PPML methods fall into interactive and non-interactive categories (Lee et al., 2022). Interactive PPML, based on MPC, requires the client and the service provider to jointly perform inference (Liu et al., 2017; Juvekar et al., 2018; Mishra et al., 2020; Lou et al., 2021; Huang et al., 2022; Diaa et al., 2024). It keeps computational costs relatively low but requires multiple communication rounds, leading to higher communication and bandwidth requirements that limit applicability. Non-interactive PPML, typically using FHE, adopts a fire-and-forget paradigm (Gilad-Bachrach et al., 2016; Brutzkus et al., 2019; Lou & Jiang, 2021; Lee et al., 2022; Sarkar et al., 2023; Ao & Boddeti, 2024), where the client encrypts inputs, the provider computes on ciphertexts, and the client decrypts the output. This reduces communication to a single round at the expense of heavy server-side computation. Despite the costs, non-interactive PPML is well-suited to third-party services as it eliminates client involvement during inference, supports low-resource devices, and works in low-bandwidth settings. Motivated by these advantages, this work focuses on non-interactive PPML.
A key challenge in non-interactive PPML is implementing activation functions with cryptographic primitives. Activations are non-polynomial, while FHE supports only polynomial operations like addition and multiplication. Prior work addresses this by approximating activations with single polynomials (Diaa et al., 2024) or piecewise polynomials (Liu et al., 2017). These approximations can be introduced before or after training. In post-training approximation, models are trained with standard activations (e.g., ReLU) and later replaced with high-precision polynomials. For example, MPCNN (Lee et al., 2022) uses minimax polynomials of orders $\{15,15,27\}$. In pre-training approximation, ReLU is replaced with a polynomial before training, allowing lower-degree approximations that reduce multiplicative depth and speed up inference. For example, shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), similar to LeNet-5 (LeCun et al., 1998), achieve over 98% accuracy on MNIST with degree-2 polynomials. However, Garimella et al. (2021) showed that training larger models (beyond AlexNet (Krizhevsky et al., 2012) or VGG-11 (Simonyan & Zisserman, 2015)) with such low-degree polynomials causes large approximation errors that destabilize training. To date, PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial, is the lowest-degree approximation that generalizes to deep neural networks.
Implementing non-interactive PPML becomes increasingly difficult as model depth increases. While bootstrapping enables evaluation of arbitrarily deep models by refreshing ciphertexts, it is an expensive operation in FHE and remains the main scalability bottleneck. To avoid this cost, prior work has relied on Leveled Fully Homomorphic Encryption (LFHE) (see §A for details). LFHE eliminates bootstrapping and allows faster private inference, but its computation is bounded by a fixed multiplicative depth determined by encryption parameters. As a result, the size and complexity of neural networks that can be evaluated under LFHE are strictly limited. The deepest CNNs shown to run with LFHE alone are AlexNet and VGG-11, as demonstrated by Garimella et al. (2021). Larger models such as ResNet-20 or ResNet-32 have so far required bootstrapping. For example, AutoFHE executes ResNet-20 inference with at least five bootstrapping layers (Ao & Boddeti, 2024), which leads to heavy computational costs. In this work, we focus on ResNet (He et al., 2016), as it is widely adopted in state-of-the-art evaluations and is well-suited for efficient and low-latency deployment.
Contributions: This work advances non-interactive PPML by enabling, for the first time, large CNNs like ResNet-18, ResNet-20, and ResNet-32 to run entirely under LFHE without bootstrapping, achieving accuracy comparable to plaintext ReLU models, while outperforming prior work in both accuracy and inference time. Our contributions are threefold:
  • To the best of our knowledge, this is the first work to train large CNNs, like ResNet-18, ResNet-20, and ResNet-32 effectively with degree-2 polynomial activations. We introduce a novel penalty function that ensures stable training and accuracy on par with ReLU-based models. This achieves the theoretical minimum multiplicative depth of one for non-linear activations, whereas the lowest previously known stable alternative, PILLAR, required depth three (see §2).
  • We introduce three novel structural optimization techniques that reduce the overall multiplicative depth of a model while preserving functional equivalence: (i) Node Fusing, which merges consecutive operations such as convolution and batch normalization by folding normalization parameters into convolution weights, eliminating separate multiplications; (ii) Weight Redistribution, which adjusts parameters so that coefficients of highest-order terms in polynomial activations and batch normalization, as well as divisors in pooling layers, normalize to one, removing redundant multiplications by constants; and (iii) Tower Reuse, which allows multiple multiplications within the same FHE level before rescaling. In the CKKS scheme used in this work (Cheon et al., 2017), ciphertexts carry a scale factor that grows with multiplications. Rescaling keeps ciphertext values within range but consumes a level. Since each level requires a modulus and homomorphic operations apply to all moduli, fewer rescaling steps reduce the number of moduli, leading to faster homomorphic operations. Moreover, they permit using larger moduli for higher accuracy (see §3).
  • We propose an encoding strategy that utilizes unused ciphertext slots in FHE to pack multiple model instances, enabling simultaneous ensemble inference to improve accuracy at some additional inference cost. To offset this overhead, we introduce a parameter clustering method for convolutions that recovers much of the lost speed. To the best of our knowledge, this is the first work to exploit unallocated ciphertext slots for ensemble inference in FHE (see §4).
We evaluate our method on CIFAR-10 and CIFAR-100 datasets (Krizhevsky & Hinton, 2009), and show that it achieves up to $4\times$ faster private inference than prior work while achieving accuracy comparable to plaintext ReLU models (see §5).
2 Model Training with Polynomial Approximation
2.1 Problem Setting
Dataset. We consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$ for multi-class classification, where $x_{i}\in\mathbb{R}^{n}$ is the feature vector and $y_{i}\in\{1,\dots,K\}$ is the class label. $\mathcal{D}$ has $M$ samples and $K$ classes.
ReLU Network. Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ be an $L$-layer feed-forward network with ReLU activations, parameterized by weight matrices $W^{(1)},\dots,W^{(L)}$, where each $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $n_{l}$ is the number of neurons in layer $l$. The network output is given by $f(x)=h^{(L)}(x)$, where $h^{(0)}(x)=x$ and $h^{(l)}(x)=ReLU(W^{(l)}h^{(l-1)}(x))$ for all $1\leq l\leq L$. The activation function $ReLU(z)$ is applied coordinate-wise and defined as $ReLU(z)=\max(0,z)$.
Polynomial Network. We define $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with the same architecture as $f$, but replace ReLU with a degree-$d$ polynomial $p_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$ that approximates ReLU over a bounded interval [c,c][-c,c]. The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are chosen so that the maximum approximation error satisfies: $\max_{z\in[-c,c]}|ReLU(z)-p_{d}(z)|\leq\varepsilon$, where $\varepsilon>0$ is a small constant. The network output is given by $g(x)=h^{(L)}_{p_{d}}(x)$, where $h^{(0)}_{p_{d}}(x)=x$ and $h^{(l)}_{p_{d}}(x)=p_{d}(W^{(l)}h^{(l-1)}_{p_{d}}(x))$ for all $1\leq l\leq L$.
Challenges. Using polynomial networks ensures FHE compatibility but introduces training challenges such as escaping activations and coefficient truncation (see §B for more details). Escaping activations occur when intermediate outputs drift outside the interval [c,c][-c,c], leading to large approximation errors. Coefficient truncation occurs when polynomial coefficients are stored with limited fixed-point precision $b$, which changes the polynomial’s shape and causes deviation from ReLU even inside the approximation interval. These issues require careful selection of $d$, [c,c][-c,c], and $b$, together with regularization during training to keep activations within the valid approximation range.
2.2 Training Strategy
Quantization-Aware Polynomial Fitting. To avoid coefficient truncation, we integrate fixed-point constraints directly into coefficient estimation using quantization-aware polynomial fitting (Diaa et al., 2024). Given an approximation interval [c,c][-c,c] and fixed-point precision with $b$ fractional bits, we define a quantized input domain for polynomial fitting as: $X=\{x\in[-c,c]\mid x=k\cdot 2^{-b},\ k\in\mathbb{Z}\}.$ For each input $x_{i}\in X$, we compute scaled ReLU outputs $Y_{i}=2^{b}\cdot ReLU(x_{i})$ so that regression targets are integers, which reduces risk of precision loss during optimization. We build a matrix $B\in\mathbb{R}^{M\times(d+1)}$ with entries $B_{i,k}=x_{i}^{k}\ \text{for }0\leq k\leq d.$ The coefficients $A=[a_{0},\ldots,a_{d}]^{\top}\in\mathbb{Z}^{d+1}$ are obtained by solving bounded integer least-squares problem: $\min_{A\in\mathbb{Z}^{d+1}}\|BA-Y\|_{2}^{2}\ \text{subject to }a_{k}\in[-2^{b-1},2^{b-1}-1].$ The resulting fixed-point polynomial is $p_{d}(x)=\sum_{k=0}^{d}(\frac{a_{k}}{2^{b}})x^{k}.$
Activation Regularization. To avoid escaping activations, we introduce a regularization strategy that constrains the inputs to polynomial activation functions (i.e., the pre-activations) to remain within the valid approximation interval [c,c][-c,c]. The classification loss supervises only the final output and provides no mechanism to limit intermediate pre-activation values. Hence, minimizing it alone does not prevent pre-activations from drifting outside the interval [c,c][-c,c], where the polynomial diverges from ReLU and destabilizes training. We introduce a layer-wise penalty that penalizes out-of-range pre-activations. For a mini-batch $\mathcal{B}\subset\mathcal{D}$, the following training loss is minimized:
$\mathcal{L}_{\mathcal{B}}\;=\;\;\underbrace{\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell_{\mathrm{CE}}\left(g\left(x\right),y\right)}_{\text{classification loss}}\;+\;\zeta\;\underbrace{\frac{1}{L}\sum_{l=1}^{L}\left\|z_{p}^{(l)}\;-\;\mathrm{clip}\left(z_{p}^{(l)};[-c,c]\right)\right\|_{2}}_{\text{clip-range penalty}}$ (1)
where $\ell_{\mathrm{CE}}$ is the cross-entropy loss, $z_{p}^{(l)}$ are pre-activations at layer $l$, i.e., outputs of affine transformations $W^{(l)}h^{(l-1)}_{p_{d}}(x)$ before the polynomial activation. The function $\mathrm{clip}(z;[-c,c])$ is applied element-wise and defined as: $\mathrm{clip}(z;[-c,c])_{i}=\max(-c,\min(z_{i},c))$. The second loss term penalizes pre-activations that lie outside the interval [c,c][-c,c], with strength controlled by the regularization parameter $\zeta>0$. See §C for a proof of the penalty function’s correctness. Training can still become unstable for two reasons: (i) in early epochs pre-activations can grow unbounded before the model learns to contain them and (ii) using full regularization weight $\zeta$ from the start can let the penalty dominate and destabilize optimization. As additional strategies, we also consider clipping pre-activations during training and introducing a warm-up schedule for $\zeta$ (see §D for more details).
3 Structural Optimizations
3.1 Node Fusing
Our first optimization is node fusing, which merges consecutive computational nodes into a single equivalent function. This reduces both number of operations and multiplicative depth. Formally, two sequential nodes $F(G(x))$ are replaced with $H(x)$ such that $F(G(x))=H(x)$, with $H(x)$ more efficient to evaluate. Since our method uses batch normalization, recall its polynomial form $B(x)=b_{1}x+b_{0},\;b_{1}=\frac{\gamma}{\sigma},\;b_{0}=\beta_{b}-b_{1}\mu$, where $\mu,\sigma$ are the input mean and standard deviation, and $\gamma,\beta_{b}$ are learnable parameters. We outline four common fusion cases. See §E.1 for full derivations.
Case 1: $P(B(x))\mapsto P(x)$. When a polynomial activation $P(\cdot)$ follows batch normalization, the two can be fused into a single quadratic polynomial. If $P(x)=c_{2}x^{2}+c_{1}x+c_{0}$, then $P(B(x))=p_{2}x^{2}+p_{1}x+p_{0}$, with fused coefficients $p_{2}=b_{1}^{2}c_{2},\;p_{1}=b_{1}(2b_{0}c_{2}+c_{1}),\;p_{0}=b_{0}^{2}c_{2}+b_{0}c_{1}+c_{0}.$
Case 2: $B(C(x))\mapsto C(x)$. When batch normalization follows a convolution $C(\cdot)$, the two can be fused (Markuš, 2018). A convolution is given by $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$, where $w_{i}$ are the convolution weights and $\beta_{c}$ is the bias. The fused form becomes a single convolution $B(C(x))=\sum_{i}\omega_{i}x_{i}+\alpha$, with fused coefficients $\omega_{i}=b_{1}w_{i},\;\alpha=b_{1}\beta_{c}+b_{0}$.
Case 3: $P(B_{X}(x)+B_{Y}(y))\mapsto S(x,y)$. In residual networks, skip connections often merge two batch-normalized branches by summing them before applying an activation $P(\cdot)$. Specifically, $P(B_{X}(x)+B_{Y}(y))$. Here, $B_{X}$ and $B_{Y}$ are independent batch normalization layers to input $x$ and $y$. This structure can be fused into a single quadratic bivariate polynomial, which we call polyskip: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2}b_{Y1}^{2},\;d_{XY}=2c_{2}b_{X1}b_{Y1}$, $d_{X}=b_{X1}(2c_{2}(b_{X0}+b_{Y0})+c_{1}),\;d_{Y}=b_{Y1}(2c_{2}(b_{X0}+b_{Y0})+c_{1})$, $d_{0}=c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}$.
Case 4: $P(B_{X}(x)+y)\mapsto S(x,y)$. Some skip connections use identity shortcuts, where the input $y$ is added directly to the batch-normalized branch $B_{X}(x)$ before applying the activation. This structure can be fused into a quadratic bivariate polynomial: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$,with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2},\;d_{XY}=2c_{2}b_{X1}$, $d_{X}=b_{X1}(2c_{2}b_{X0}+c_{1}),\;d_{Y}=2c_{2}b_{X0}+c_{1}$, $d_{0}=c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}$.
Node fusing collapses batch normalization, activations, convolutions, and skip connections into single fused polynomials, eliminating redundant nodes and reducing multiplicative depth.
3.2 Weight Redistribution
Our second optimization is weight redistribution, a technique that redistributes weights across the network while maintaining functional equivalence. The goal is to reduce the multiplicative depth of certain functions by one. It specifically targets average pooling, batch normalization, and polynomial activations. For average pooling, redistribution eliminates the normalization step by setting the divisor to one. For polynomial functions, the highest-order coefficient is set to one. For example, the polynomial activation $c_{2}x^{2}+c_{1}x+c_{0}$ is transformed into $x^{2}+p_{1}x+p_{0}$, reducing its multiplicative depth from two to one, the theoretical minimum for an activation function. These transformations alone break model equivalence. To maintain it, other nodes in the network must compensate for the change. We refer to the nodes initiating weight redistribution as donors, and the nodes adjusting their parameters to preserve equivalence as receivers. We represent the network as a directed graph and traverse it, identifying all eligible donor nodes. For each donor, redistribution can be applied either forward or backward, affecting receivers among the donor’s successors or predecessors, respectively.
Update Forward. Donors. We first consider donor updates in the forward direction. Our goal is to construct a normalized function $\bar{F}(x)$ such that $\upsilon\bar{F}(x)=F(x)$, where $F(x)$ is the original donor function and $\upsilon$ is the update term to be propagated to receivers. The average pooling operation is given by $\mu(x)=k^{-1}\sum_{i}{x_{i}}$, where $k$ is the kernel size. In the normalized function we set $k=1$, reducing the operation to $\bar{\mu}(x)=\sum_{i}x_{i}$. Thus, for the equality to be valid, we have $\upsilon=k^{-1}$.
Batch normalization and polynomial activation are polynomial functions. For a univariate polynomial $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$ of degree $d$, and a normalized polynomial $\bar{P}(x)=x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}$, where $\bar{c}_{i}$ are the normalized coefficients, we have $\upsilon=c_{d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-1}$ for $\upsilon\bar{P}(x)=P(x)$ to hold. The extension to the bivariate polynomial activation is analogous and deferred to §E.2.
Receivers. After a donor update, receivers must be adjusted to maintain model equivalence. The original composed function is $G(F(\cdot))$, where $F(\cdot)$ is the donor and $G(\cdot)$ is the receiver. Updating the donor to $\bar{F}(\cdot)$ without compensating the receiver results in $G(\upsilon^{-1}F(\cdot))$, breaking model equivalence. To restore it, we modify the receiver so that $\bar{G}(\bar{F}(\cdot))=\bar{G}(\upsilon^{-1}F(\cdot))=G(F(\cdot))$.
Receivers fall into two categories: kernel functions and polynomial functions. Kernel functions share the form $K(x)=\sum_{i}w_{i}x_{i}+\beta$. To maintain model equivalence, we must update the kernel function such that $\bar{K}(\upsilon^{-1}x)=K(x)$. For that, we must set $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta$.
Polynomial functions update as $\bar{P}(\upsilon^{-1}x)=P(x)$, which implies that $\bar{c}_{i}=c_{i}\upsilon^{i}$. For the bivariate case, in $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$ the index $i$ corresponds to the exponent of $x$, while in $\bar{S}(x,\upsilon^{-1}y)=S(x,y)$ it corresponds to the exponent of $y$.
Update Backward. Donors. Backward updates are essentially the inverse of forward updates. For average pooling, the update is identical, since $\bar{\mu}(\upsilon x)=\upsilon\bar{\mu}(x)$, while for polynomial donors, it must hold that $\bar{P}(\upsilon x)=P(x)$; hence, $\upsilon=c_{d}^{1/d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-i}$.
Receivers. Receivers must satisfy $\bar{G}(x)=\upsilon G(x)$ to preserve model equivalence. Kernel receivers are updated as $\bar{K}(x)=\upsilon K(x)$, which implies that $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta\upsilon$. And polynomial receivers update according to $\bar{P}(x)=\upsilon P(x)$; therefore, $\bar{c}_{i}=c_{i}\upsilon$.
3.3 Tower Reuse
In CKKS, the Residue Number System (RNS) moduli are typically chosen as primes close to the scaling factor $\Delta$ (see §A for an overview of LFHE). This choice is motivated by the fact that, after each homomorphic multiplication, the scaling factors of the operands are multiplied, producing a ciphertext with scale $\Delta^{2}$. To control this growth, the rescale operation reduces the scale back to $\Delta$ by removing one modulus from the modulus chain. However, rescaling introduces approximation errors: the further the dropped modulus deviates from $\Delta$, the larger the error. These errors may accumulate and amplify through subsequent homomorphic operations. A larger $\Delta$ is desirable for two reasons: (i) it provides higher precision for CKKS encoding and (ii) the RNS moduli are relatively closer to the scaling factor, reducing approximation errors.
The drawback is that a larger $\Delta$ requires larger RNS moduli, which in turn increases the ciphertext modulus $Q=\prod_{i=0}^{\mathcal{L}}q_{i}$, where $\mathcal{L}$ is the number of levels and $q_{i}$ are the individual RNS primes. Since the security level depends on the ratio $N/Q$ (with $N$ the polynomial degree), a larger $Q$ reduces security. One way to increase security is to increase $N$, but this slows computations due to larger polynomials. Thus, for a fixed security level and polynomial degree, there is a maximum allowable $Q$. Consequently, for an application-defined multiplicative depth $\delta$, this imposes a bound on the maximum scaling factor $\Delta$. In practice, when $\delta$ is large, as is common in deep learning models, the resulting $\Delta$ is small. This has two adverse effects, opposite to the benefits of a large $\Delta$: lower numerical precision and larger approximation errors, as moduli $q_{i}$ are relatively farther from $\Delta$.
Proposed Solution. We introduce a more general method for determining the RNS moduli by introducing sublevels within each level. Instead of enforcing $q_{i}\approx\Delta$, we allow $q_{i}\approx\Delta^{\ell}$, where $\ell$ denotes the number of sublevels. Rescaling is performed only when a ciphertext scale exceeds the sublevel capacity of the modulus to be dropped. Formally, $\lambda(x)>\lambda(q_{i})\implies\downarrow x$, where $\downarrow$ denotes the rescale operation, and $\lambda(x)=\left\lfloor\log_{\Delta}{\Delta_{x}}\right\rceil$ returns the sublevel of a ciphertext, plaintext, or modulus: with $\Delta_{x}$ referring to the scale of $x$ and $\Delta$ being the default scale defined by the encryption parameters. See §E.3 for an example of this technique.
This method allows smaller $\Delta$ values while maintaining larger $q_{i}$, mitigating rescaling errors. Moreover, fewer effective levels (moduli) reduce error amplification, and since homomorphic operations are applied across all moduli, reducing their number improves computational efficiency.
3.4 Impact on RNS Levels
The combination of our polynomial activation function with the structural optimizations proposed in this section reduces the number of RNS levels $\mathcal{L}$ required for multiplication compared to PILLAR (Diaa et al., 2024), which previously achieved the lowest $\mathcal{L}$: from 87 to 18 for ResNet-18, from 97 to 20 for ResNet-20, and from 157 to 32 for ResNet-32. A complete analysis of the contribution of each technique to the reduction in required levels for ResNet models is provided in §E.4.
4 Codesign Techniques
4.1 Data Layout
Data layout defines how model inputs and weights are mapped into FHE ciphertext and plaintext slots during encoding. In this work, we adopt the HW layout (Dathathri et al., 2019), which assigns
Figure 1: Illustration of HW layout for convolution on a $3\times 3\times 3$ input and two $3\times 2\times 2$ filters, configured with padding 0 and stride 1.
each input channel to a separate ciphertext and fills its slots with spatial values indexed by two–dimensional coordinates. Unlike the naive layout (Gilad-Bachrach et al., 2016), where each ciphertext holds only one value and incurs significant overhead, HW layout is more efficient, packing $h\cdot w$ values per ciphertext, with $h$ and $w$ denoting input height and width. Figure 1 shows this for a $3\times 3\times 3$ input with two $3\times 2\times 2$ filters (stride 1, no padding), illustrating slot arrangements and gaps that may appear after convolution or other kernel operations. These gaps could be removed by remapping, but that requires additional multiplications and increases multiplicative depth. Instead, we use an adaptive (lazy) mapping strategy: some slots are left unused, and subsequent weights, biases, and coefficients are aligned with the slot arrangement of the preceding layer’s output. HW layout does not always fully utilize slots, especially when input channels have far fewer values than available slots. More compact layouts such as CHW layout (Dathathri et al., 2019) and its variants (Lee et al., 2022) address this by mapping multiple channels into one ciphertext, which reduces ciphertext counts and decreases the number of multiplications and additions, ultimately leading to faster inference. However, PPML relies on polynomial approximations that reduce accuracy compared to plaintext ReLU models. For this reason, we retain the less efficient HW layout, repurposing unused slots to improve accuracy instead of minimizing inference time. To offset the added latency, we introduce a parameter clustering technique.
4.2 Clustering of Convolution Parameters
Convolution layers account for most parameters in deep neural networks, and their weight handling under FHE is costly since each weight must be repeatedly encoded to match varying ciphertext layouts and levels. As these depend on kernel position and layer structure, the same scalar weight is often redundantly encoded. Pre-encoding all weight–layout combinations would lead to prohibitive memory usage, while on-demand encoding reduces memory but adds heavy computational overhead. To mitigate this, weights are limited to a small fixed set of representative values. Each representative is encoded once into a plaintext codebook, and during inference, the appropriate plaintext is retrieved rather than recomputed. Time and memory then scale with codebook size instead of the number of weights. Clustering provides a way to build this representative set by grouping similar weights and replacing each with its closest centroid, bounding the number of unique encodings.
Formally, let a polynomial network contain $L$ convolution layers. The weight tensor of layer $l$ is $W^{(l)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$, where $O_{l}$ is the number of output channels, $I_{l}$ the number of input channels, and $H_{l}\times W_{l}$ the kernel size. Flattening all convolution weights gives $\theta=(\theta_{1},\dots,\theta_{P})\in\mathbb{R}^{P},$ where $P$ is the total number of parameters. To compress the model, each $\theta_{j}$ is replaced by its nearest codebook value from $\mathcal{C}=\{c_{1},\dots,c_{k}\}\subset\mathbb{R}$, where $k\ll P$ is the number of representatives in the codebook. The codebook entries act as centroids approximating the original weights. The choice of $k$ determines the efficiency-accuracy tradeoff: smaller $k$ reduces the number of encodings but increases approximation error, while larger $k$ produces finer approximation at higher computational and memory cost. Given a distance function $d:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ (assumed to be non-negative), each parameter is quantized as: $\tilde{\theta}_{j}=\arg\min_{c\in\mathcal{C}}d(\theta_{j},c)\text{ for all }j=1,\dots,P,$ producing a quantized parameter vector $\tilde{\theta}\in\mathcal{C}^{P}$. Quantization is applied elementwise, so each convolutional kernel retains its shape while its values are drawn from the codebook. The choice of clustering strategy determines how the codebook is constructed and how assignments are made.
Full Clustering. A straightforward strategy is to cluster the entire parameter vector $\theta\in\mathbb{R}^{P}$ using a single global codebook $\mathcal{C}$. All convolution parameters, regardless of layer or kernel position, are quantized to the same shared set of representative values. This produces a uniform quantization scheme with easy implementation and a small codebook size.
Limitation. While full clustering reduces the number of distinct weight values, it does not eliminate repeated plaintext encodings. Each weight must still match the ciphertext layout and level of its layer, so the same scalar value appearing in different layers must be re-encoded. Even within a single convolution, all weights share the same level, but weights from different filter columns map to different layouts and require separate encodings. Only weights in the same column across layer filters align in layout and level. We define this group as a slice, which can share plaintext encoding. Thus, full clustering reduces codebook size but not redundancy from mismatched levels and layouts.
Slice Clustering. To address the redundancy remaining under full clustering, we refine quantization to the slice level. Since only weights within the same slice can share plaintext encodings, clustering is applied per slice rather than across all parameters. For a given layer $l$ and spatial position $s\in\{1,\dots,W_{l}\}$, the slice is $S^{(l)}_{s}=\{W^{(l)}_{o,i,h,s}|o=1,\dots,O_{l};\ i=1,\dots,I_{l};\ h=1,\dots,H_{l}\},$ which can be viewed as a vector in $\mathbb{R}^{O_{l}I_{l}H_{l}}$ (see §F.1 for an illustration). Each slice $S^{(l)}_{s}$ is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}\subset\mathbb{R}$, and quantization is applied elementwise: $\tilde{W}^{(l)}_{o,i,h,s}=\arg\min_{c\in\mathcal{C}^{(l)}_{s}}d(W^{(l)}_{o,i,h,s},c).$ This ensures clustering respects FHE structure while adapting to the local weight distribution of each slice. The additional storage from slice-specific codebooks is modest compared to the resulting gains in accuracy and efficiency.
4.3 Ensemble of Polynomial Network
Inference with a single polynomial network $g(x)$ often shows high variance across training runs. Let $g_{(m)}(x)\in\mathbb{R}^{K}$ be the logits from the $m$-th independently trained instance. For a fixed input, these logits can differ significantly, especially near decision boundaries, due to (1) approximation error from the fixed-point polynomial activation $p_{d}(\cdot)$ and (2) training stochasticity such as random weight initialization and mini-batch ordering. As a result, different $g_{(m)}$ may predict different classes for the same input. To reduce this variance, we use an ensemble of $M$ polynomial networks $\{g_{(m)}\}_{m=1}^{M}$. Each has the same architecture and activation $p_{d}(\cdot)$ but is trained independently with different seeds and mini-batch orders. For input $x$, the ensemble output is the average of logits: $\overline{g}(x)=\frac{1}{M}\sum_{m=1}^{M}g_{(m)}(x)$ with the predicted class $\hat{y}=\arg\max_{k\in\{1,\dots,K\}}\overline{g}(x)_{k}.$ Averaging smooths training noise and polynomial approximation errors, producing more stable and accurate predictions. Each $g_{(m)}$ is trained with the regularized loss $\mathcal{L}_{\mathcal{B}}$ (Eq. 1), ensuring compatibility with fixed-point polynomial inference under FHE. Crucially, this ensemble design adds no extra computation or memory overhead. As noted in §4.1, our HW layout leaves some ciphertext slots unused. We fill these slots with weights from different models, while all ensemble members share the same ciphertexts for inputs and activations. This reuse of ciphertext–plaintext structures avoids redundant ciphertexts and repeated encodings, making ensemble inference practical in the FHE framework.
Limits of Ensemble with Clustering. While ensemble inference reuses unused ciphertext slots, its combination with parameter clustering creates a challenge. In slice clustering, each slice $S_{s}^{(l)}$ with $O_{l}\times I_{l}\times H_{l}$ weights is represented by $k$ centroids in codebook $\mathcal{C}_{s}^{(l)}$, reducing plaintext encodings to $k$ per slice. In an ensemble, however, each plaintext must encode parameters from all $M$ models. As centroids are unlikely to align across models, the count of unique encodings per slice grows as $\mathcal{O}(k^{M})$, canceling the gains of clustering and making the naive combination impractical.
Slice Ensemble Clustering. To address the inefficiency of independent clustering, we extend slice clustering to ensembles by enforcing shared representatives across models. We cluster weights jointly at the same kernel position so that all models use a common codebook. Formally, the ensemble has $M$ polynomial networks. For convolution layer $l$, model $m$ has weights $W^{(l,m)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$. For kernel position $s\in\{1,\dots,W_{l}\}$, we extract slices: $S^{(l,m)}_{s}=\{W^{(l,m)}_{o,i,h,s}|o=1,\dots,O_{l};i=1,\dots,I_{l};h=1,\dots,H_{l}\}.$ Stacking slices across $M$ models produces $X^{(l)}_{s}=[S^{(l,1)}_{s}\;S^{(l,2)}_{s}\cdots S^{(l,M)}_{s}]\in\mathbb{R}^{N_{s}\times M},$ where $N_{s}=O_{l}I_{l}H_{l}$. Each row is an $M$-dimensional vector for the same weight coordinate across models (see §F.2 for an illustration). We cluster rows in $\mathbb{R}^{M}$: $\min_{\mathcal{C}^{(l)}_{s}\subset\mathbb{R}^{M}}\sum_{j=1}^{N_{s}}\min_{c\in\mathcal{C}^{(l)}_{s}}d(X^{(l)}_{s,j},c),$ with codebook $\mathcal{C}^{(l)}_{s}=\{c_{1},\dots,c_{k}\}$. Each $X^{(l)}_{s,j}$ is replaced by its nearest centroid, producing quantized slices $\tilde{X}^{(l)}_{s,j}\in\mathcal{C}^{(l)}_{s}$. Quantized weights $\tilde{W}^{(l,m)}$ are reconstructed by reshaping slices. This shared clustering aligns weights across models, mapping small variations at the same coordinate to a common centroid. It thus requires fewer distinct encodings, reducing codebook size and inference time while retaining the accuracy gains of ensemble.
5 Experimental Results
5.1 Experimental Setup
We trained ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using PyTorch 2.4.1. Each experiment was repeated 10 times, and we report mean accuracy and private inference time with standard deviations. Private inference used a C++20 framework we developed, built on Microsoft SEAL 4.1 (SEAL, ) for leveled FHE operations and GMP 6.2.1 for multiprecision arithmetic. The framework includes a tool that automatically converts trained models for C++ inference. We use a machine with two AMD EPYC 64-core processors, 2 TB memory across 32 DDR4 DIMMs at 2933 MT/s, on Ubuntu 22.04.5 LTS with kernel 5.15.0-151-generic (x86_64) to run all experiments, including related work. For model training (see §2), we used hyperparameters $c=2$ and $\zeta=0.001$. An ablation study for both is provided in §G. We set $b=10$, aligned with PILLAR (Diaa et al., 2024), as it provides accurate polynomial approximation after quantization-aware fitting. For all clustering strategies, we used the $\ell_{2}$ norm as the distance metric, with k-means as the clustering algorithm (Lloyd, 1982). Details on FHE encryption parameters are in §H.1. Additional results, including peak memory, are reported in §H.3.
We compare our approach against non-interactive PPML baselines, MPCNN (Lee et al., 2022) and AutoFHE (Ao & Boddeti, 2024), in terms of accuracy and inference time. For fairness, we use their publicly available implementations (git, 2024a; 2022) with specified encryption parameters. Both baselines implement ResNet-20 on CIFAR-10 and ResNet-32 on CIFAR-10 and CIFAR-100, all with $N=2^{16}$. We also trained plaintext models using standard ReLU activations to serve as a reference for accuracy. In this section, we mainly report results for ResNet-20 and ResNet-32, since these allow direct comparison with baselines. Results for ResNet-18 are provided in §H.2.
5.2 Summary of Results
Accuracy of Polynomial Approximation. Training with the polynomial approximation described in §2 shows that, on average across CIFAR-10 and CIFAR-100, our degree-2 polynomial achieves accuracy only about 1.3% lower than ReLU, while PILLAR with degree-4 polynomial is about 5.4% lower. A degree-2 adaptation of PILLAR shows an average drop of 6.5%. These results demonstrate that our method can use low-degree polynomials effectively, while PILLAR does not reach comparable accuracy even with higher degrees. A detailed comparison is provided in §G.3.
[…]
Figure 2: Accuracy vs. private inference time for ResNet-20/32 on CIFAR-10/100. Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model). We compare Standard ensembles (no clustering), Slice ensembles (slice-wise clustering), and Full Clustering. Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering. Baselines include AutoFHE (numbers = bootstrapping layers) and MPCNN. The green dashed line shows plaintext ReLU accuracy.
Ensemble Evaluation. We evaluate ensembles with $M=\{1,2,4\}$ models. Figure 2 (middle row) shows accuracy and inference time for two approaches: Standard-$M$ (ensembles without clustering) and Slice-$M$ (ensembles with slice clustering as in §4.3). Inference time does not increase with ensemble size since all models fit in one ciphertext, requiring the same number of homomorphic operations. For Standard-MM\!, accuracy consistently improves as $M$ increases. Slice-MM\!, however, does not always scale the same way. Since each centroid represents an $M$-dimensional space, higher $M$ can require more centroids to capture model parameters. Thus, Slice-4 can underperform Slice-2 at a fixed $k$, and Slice-2 needs more centroids to match Standard-2 accuracy. Despite this, Slice-$M$ provides notable advantages. Slice-2 mostly outperforms Standard-2 and reaches accuracy comparable to Standard-4. This suggests clustering acts as an implicit regularizer, mitigating overfitting.
Comparison with Related Work. Figure 2 (bottom row) presents Pareto fronts of Slice Clustering in terms of accuracy and private inference time, compared with AutoFHE and MPCNN. For reference, we also report Standard-4 and plaintext ReLU accuracy. For ResNet-20, where our models operate with $N=2^{15}$, Slice Clustering achieves consistently faster inference than related work (up to $4\times$). For ResNet-32, where all approaches use $N=2^{16}$, our method is comparable in inference time to AutoFHE on CIFAR-10 and $2\times$ faster on CIFAR-100. Inference time does not change for our approach for different datasets, whereas AutoFHE slows down because it requires a high number of bootstrapping operations to maintain accuracy. The main factors affecting our inference time relative to related work are: The main factors affecting our inference time relative to related work are: (i) the polynomial activation (§2) and structural optimizations (§3) reduce both multiplicative depth and number of RNS moduli, allowing execution under LFHE without bootstrapping; (ii) our clustering technique (§4.2) further reduces the number of unique parameters, thereby limiting plaintext encodings per convolution; (iii) for ResNet-20 (and ResNet-18), we operate at a smaller polynomial degree, which directly accelerates homomorphic operations; (iv) our choice of data layout (§4.1) deliberately sacrifices some latency to better exploit ciphertext slots for accuracy, but above optimizations offset this overhead. In terms of accuracy, our approach consistently outperforms related work thanks to our training strategy (§2). In fact, it closely matches plaintext ReLU models, effectively eliminating the accuracy gap typically observed in polynomial approximations of ResNets.
6 Conclusion
[…]
Appendix A Leveled Fully Homomorphic Encryption
Homomorphic Encryption is a special type of encryption that allows meaningful computations to be performed directly on encrypted data. In this work, we employ the Cheon-Kim-Kim-Song (CKKS) (Cheon et al., 2017) encryption scheme, specifically the Residue Number System (RNS) variant (Cheon et al., 2018), which supports approximate arithmetic over encrypted real or complex numbers. CKKS is an FHE scheme based on the Ring-Learning With Errors (RLWE) problem (Lyubashevsky et al., 2013). The scheme operates across three domains: message $\mathcal{M}=\mathbb{C}$, plaintext $\mathcal{P}=R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$, and ciphertext $\mathcal{C}=R_{Q}\times R_{Q}$ spaces. The encoding function $encode(\cdot)=\mathcal{M}^{N/2}\mapsto\mathcal{P}$ maps a vector of $N/2$ complex numbers into a degree-$N$ polynomial in the plaintext space. The encryption function $encrypt(\cdot)=\mathcal{P}\mapsto\mathcal{C}$ transforms a plaintext into a ciphertext. In RNS-CKKS, the ciphertext modulus $Q$ is decomposed into $\mathcal{L}+1$ smaller primes $Q=\prod_{i=0}^{\mathcal{L}}{q_{i}}$, where $\mathcal{L}$ denotes the level. A polynomial in $R_{Q}$ is represented as $\mathcal{L}+1$ polynomials in $R_{q_{i}}\forall i\in[0,\mathcal{L}]\cap\mathbb{Z}$, which allows computations using native 64-bit integer arithmetic rather than costly arbitrary-precision operations.
CKKS supports three homomorphic operations: addition, multiplication, and rotation. Additions and multiplications act element-wise on the encoded vector, similar to SIMD vector operations. The rotation operation rotates the vector of $N/2$ encoded messages by a step $s\in\mathbb{Z}_{N/2}$. During encoding, coefficients are scaled by a factor $\Delta$, rounded to the nearest integers, and reduced modulo $Q$, making CKKS a fixed-point FHE scheme. Multiplying two encoded messages scales the result by $\Delta^{2}$, requiring a rescale operation to restore the scale to $\Delta$. Rescaling removes one RNS modulus, reducing the level by 1. It also accumulates approximation errors, since the RNS modulus being discarded and the scaling factor $\Delta$ are not equal. When the level reaches zero, bootstrapping can reset the ciphertext to a higher level, enabling unbounded computation. However, bootstrapping is computationally expensive and avoided when possible. FHE without bootstrapping is called Levelled Fully Homomorphic Encryption (LFHE). In LFHE, the encryption parameters $(N,Q)$ are chosen to provide the desired security and sufficient levels for a predefined arithmetic circuit with known multiplicative depth. To the best of our knowledge, this work is the first to apply LFHE to deep CNN models like ResNet-20 and ResNet-32.
Appendix B Challenges of Using Polynomial Activation
ReLU, defined as $ReLU(x)=\max(0,x)$, is one of the most widely used activation functions in deep learning due to its simplicity and computational efficiency. Its piecewise linear structure introduces the non-linearity necessary for neural networks to capture complex patterns in data. While ReLU is straightforward to implement in plaintext settings, it poses significant challenges in FHE environments due to its reliance on conditional branching. FHE can evaluate only polynomial functions as it supports only additions and multiplications, thus, any FHE representation of ReLU must be a polynomial approximation. Evidently, the higher degree the polynomial, the more precise the approximation. However, the multiplicative depth of the activation functions grows logarithmically to the degree of the polynomial used for approximation. Precisely, the multiplicative depth can be computed as $\delta=\lceil\log_{2}{d}+1\rceil$, where $d$ denotes the polynomial degree and the +1+1 refers to the coefficient multiplication. In the literature, the lowest multiplicative depth for a polynomial approximation of ReLU is achieved by PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial of the form $\sum_{i=0}^{4}{c_{i}x^{i}}$ with $\delta=3$. The theoretical minimum multiplicative depth for an activation function is one, achieved with a polynomial of the form $x^{2}+c_{1}x+c_{0}$. Polynomials like this have been effectively used in shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), but failed to work for deeper models due to the escaping activation problem (Garimella et al., 2021). In addition, polynomial approximations are inherently limited to bounded intervals of the input domain. When substituted directly for ReLU during model training, these approximations often lead to a significant drop in model accuracy (Garimella et al., 2021; Hussain et al., 2021; Diaa et al., 2024). This section analyzes two primary causes of this degradation: escaping activations, where intermediate layer outputs fall outside the polynomial’s approximation interval and coefficient truncation, which results from representing polynomial coefficients in fixed-point format to ensure compatibility with FHE arithmetic.
B.1 Escaping Activation Problem
Let $p_{d}(x)=\sum_{k=0}^{d}a_{k}x^{k}$ denote a degree-$d$ polynomial approximation of the ReLU function, where each coefficient $a_{k}\in\mathbb{R}$ for all $k\in\{0,\dots,d\}$. These coefficients are typically obtained by minimizing the least-squares error between the polynomial and the ReLU function over a finite set of real-valued sample points $\{x_{i}\}_{i=1}^{N}\subset[-c,c]$, for some parameter $c>0$:
$\min_{a_{0},\ldots,a_{d}}\;\sum_{i=1}^{N}\left(ReLU(x_{i})\;-\;p_{d}(x_{i})\right)^{2}$
The resulting polynomial $p_{d}(x)$ provides a close approximation to ReLU within the interval [c,c][-c,c]. However, outside the interval, the polynomial behavior differs significantly from that of ReLU. ReLU exhibits piecewise linear growth:
$ReLU(x)\;=\;\begin{cases}0,&\text{if }x\leq 0\\ x,&\text{if }x>0\end{cases}\quad\Rightarrow\quad ReLU(x)\;=\;\mathcal{O}(x)$
In contrast, the growth of $p_{d}(x)$ for $d\geq 2$ is polynomial, dominated asymptotically by the highest-degree term:
$p_{d}(x)\;=\;a_{d}x^{d}\;+\;a_{d-1}x^{d-1}\;+\;\cdots\;+\;a_{1}x\;+\;a_{0}\;=\;\mathcal{O}(x^{d})$
This fundamental mismatch in growth rates gives rise to the problem of escaping activations, first identified by Garimella et al. (2021). As inputs propagate through the layers of a neural network, the intermediate values passed into the polynomial activation can escape the intended interval [c,c][-c,c], entering regions where $p_{d}(x)$ no longer approximates ReLU accurately. As a consequence, the network can produce excessively large activations, which in turn cause exploding weights and rapid degradation of model performance unless specific modifications are made to the training procedure to contain them.
B.2 Coefficient Truncation
In privacy-preserving inference based on FHE, all computations are carried out using fixed-point arithmetic with limited precision. This constraint restricts the range and resolution of values that can be accurately represented, affecting both the domain of activation function inputs and the magnitudes of polynomial approximation coefficients. The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are typically derived via floating-point least-squares fitting. To enable fixed-point evaluation under FHE, these coefficients are quantized using:
$\tilde{a}_{k}\;=\;\frac{\lfloor a_{k}\cdot 2^{b}\rceil}{2^{b}}$
where $b\in\mathbb{N}$ represents the number of fractional bits in the fixed-point format and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. The least-squares fitting procedure often produces small-magnitude coefficients, particularly for higher-degree monomials. If any coefficient satisfies $\left|a_{k}\right|<2^{-(b+1)}$, the quantized value $\tilde{a}_{k}$ becomes zero. This effectively discards the corresponding monomial $x^{k}$ from the approximation. This phenomenon, referred to as coefficient truncation (Diaa et al., 2024), changes the shape of the polynomial and can cause significant deviation from the intended ReLU behavior, even within the designated approximation interval [c,c][-c,c].
Appendix C Why the Penalty Function Works?
Lemma 1 (Pre-activation Update Decomposition).
Let $z_{p}^{(l)}\;=\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)\in\mathbb{R}^{n_{l}}$ denote the pre-activation vector at layer $l$ for input $x$. We define two quantities based on this vector: the clipping residual
$d^{(l)}\;=\;z_{p}^{(l)}\;-\;\mathrm{clip}\left(z_{p}^{(l)};[-c,c]\right)$
which measures the amount by which $z_{p}^{(l)}$ exceeds the clipping range, and the gradient of the cross-entropy loss with respect to the pre-activation
$g^{(l)}\;=\;\frac{\partial}{\partial z_{p}^{(l)}}\ell_{\mathrm{CE}}\left(g(x),y\right)$
which captures the sensitivity of the loss to changes in $z_{p}^{(l)}$. After a single gradient-descent update with learning rate $\eta$ on the minibatch loss $\mathcal{L}_{\mathcal{B}}$, the change in $z_{p}^{(l)}$ decomposes as
$\boxed{\Delta z_{p}^{(l)}=\underbrace{\left(-\eta\right)\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}g^{(l)}}_{\displaystyle\Delta z_{\rm CE}^{(l)}}+\underbrace{\left(-\eta\right)\zeta\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\frac{d^{(l)}}{\|d^{(l)}\|_{2}}}_{\displaystyle\Delta z_{\rm pen}^{(l)}}}$
Here, $\Delta z_{\rm CE}^{(l)}$ is the component induced by the cross‐entropy loss, while $\Delta z_{\rm pen}^{(l)}$ is the component induced by the clip‐range penalty.
Consider a single training example $(x,y)$, where $x$ is the input and $y$ the corresponding true label. The total per-sample loss includes two terms:
$\ell(x)=\ell_{\mathrm{CE}}\left(g(x),y\right)\;+\;\zeta\;R\left(z_{p}^{(l)}\right),\quad R\left(z_{p}^{(l)}\right)\;=\;\left\|d^{(l)}\right\|_{2}$
The first term is the standard cross‑entropy loss computed from the network output $g(x)$. The second term is a layer-specific penalty that acts only on the pre-activations at layer $l$, penalizing values that fall outside the clipping interval. A gradient-descent step with learning rate $\eta$ updates the weight matrix by
$\Delta W^{(l)}$ $\;=\;-\eta\;\frac{\partial}{\partial W^{(l)}}\;\ell(x)$
$\;=\;-\eta\left(\frac{\partial}{\partial W^{(l)}}\;\ell_{\mathrm{CE}}\left(g(x),y\right)\;+\;\zeta\;\frac{\partial}{\partial W^{(l)}}\;R\left(z_{p}^{(l)}\right)\right)$
Since $z_{p}^{(l)}$ depends linearly on the weights, its Jacobian with respect to $W^{(l)}$ is
$\frac{\partial}{\partial W^{(l)}}\;z_{p}^{(l)}=\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}.$
Using the chain rule, the gradient of the cross-entropy term with respect to the weights becomes
$\frac{\partial}{\partial W^{(l)}}\;\ell_{\mathrm{CE}}\left(g(x),y\right)$ $\;=\;\frac{\partial}{\partial z_{p}^{(l)}}\;\ell_{\mathrm{CE}}\left(g(x),y\right)\;\frac{\partial}{\partial W^{(l)}}\;z_{p}^{(l)}$
$\;=\;g^{(l)}\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}$
The penalty function $R\left(z_{p}^{(l)}\right)\;=\;\left\|d^{(l)}\right\|_{2}$ is nonzero only when some elements lie outside [c,c][-c,c]. Its gradient with respect to $z_{p}^{(l)}$ is
$\frac{\partial}{\partial z_{p}^{(l)}}\,R\left(z_{p}^{(l)}\right)\;=\;\begin{cases}\dfrac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}&\text{if }\left\|d^{(l)}\right\|_{2}>0,\\ 0&\text{if }\left\|d^{(l)}\right\|_{2}=0.\end{cases}$
This expression accounts for the case where the clipping has no effect; that is, when $z_{p}^{(l)}\in[-c,c]$ elementwise, the clipping residual $d^{(l)}$ becomes zero, and the gradient vanishes accordingly. Applying the chain rule again, we compute
$\frac{\partial}{\partial W^{(l)}}\;R\left(z_{p}^{(l)}\right)$ $\;=\;\frac{\partial}{\partial z_{p}^{(l)}}\;R\left(z_{p}^{(l)}\right)\;\frac{\partial}{\partial W^{(l)}}\;z_{p}^{(l)}$
$\;=\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}$
Putting it all together, we obtain the total gradient-based update for $W^{(l)}$
$\Delta W^{(l)}\;=\;-\eta\left(g^{(l)}\;+\;\zeta\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\right)\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}$
By definition, the change in pre-activation $z_{p}^{(l)}$ after one gradient descent step is given by
$\Delta z_{p}^{(l)}$ $\;=\;\overline{z}_{p}^{(l)}\;-\;z_{p}^{(l)}$
$\;=\;\overline{W}^{(l)}\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
$\;=\;\left(W^{(l)}\;+\;\Delta W^{(l)}\right)\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
$\;=\;\Delta W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
where $\overline{z}_{p}^{(l)}$ and $\overline{W}^{(l)}$ are updated pre-activations and weight matrix after one gradient-descent step. Substituting the previously derived expression for $\Delta W^{(l)}$
$\Delta z_{p}^{(l)}$ $\;=\;-\eta\left(g^{(l)}\;+\;\zeta\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\right)\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}\;h^{(l-1)}_{p_{d}}(x)$
$\;=\;-\eta\left(g^{(l)}\;+\;\zeta\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\right)\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$
$\;=\;\underbrace{\left(-\eta\right)\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;g^{(l)}}_{\displaystyle\Delta z_{\rm CE}^{(l)}}+\underbrace{\left(-\eta\right)\zeta\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\frac{d^{(l)}}{\|d^{(l)}\|_{2}}}_{\displaystyle\Delta z_{\rm pen}^{(l)}}$
we obtain the penalty components $\Delta z_{\rm CE}^{(l)}$ and $\Delta z_{\rm pen}^{(l)}$ as mentioned in the lemma. ∎
Lemma 2 (Clipping Gradient Pullback).
Let $\Delta z_{\rm pen}^{(l)}$ be the penalty-induced component of the pre-activation update from Lemma 1, and $d^{(l)}$ be the clipping residual. Then the inner product between $\Delta z_{\rm pen}^{(l)}$ and $d^{(l)}$ satisfies
$\boxed{\left\langle\Delta z_{\rm pen}^{(l)},\;d^{(l)}\right\rangle=-\;\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\left\|d^{(l)}\right\|_{2}\;<\;0}$
whenever $\left\|d^{(l)}\right\|_{2}\neq 0$. Therefore, $\Delta z_{\mathrm{pen}}^{(l)}$ “pulls back” each element of $z_{p}^{(l)}$ lying outside [c,c][-c,c] towards the clipping interval.
The clipping residual $d^{(l)}$ is defined as the difference between the current pre‑activation $z_{p}^{(l)}$ and its clipped version. $\Delta z_{\rm pen}^{(l)}$ is the component in $\Delta z_{p}^{(l)}$ induced by the clip-range penalty. Taking the inner product of $\Delta z_{\rm pen}^{(l)}$ with the clipping residual $d^{(l)}$ gives
$\left\langle\Delta z_{\rm pen},\;d^{(l)}\right\rangle$ $\;=\;\left\langle-\;\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}},\;d^{(l)}\right\rangle$
$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\left\langle\frac{d^{(l)}}{\|d^{(l)}\|_{2}},\;d^{(l)}\right\rangle$
$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{1}{\left\|d^{(l)}\right\|_{2}}\;\langle d^{(l)},\,d^{(l)}\rangle$
$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{1}{\left\|d^{(l)}\right\|_{2}}\,\left\|d^{(l)}\right\|_{2}^{2}$
$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\left\|d^{(l)}\right\|_{2}$
Each factor on the right-hand side: $\eta$ (learning rate), $\zeta$ (regularization parameter), $\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$ (squared norm of the preceding hidden representation) and $\left\|d^{(l)}\right\|_{2}$ (magnitude of the clipping residual) is strictly positive whenever any pre‑activation lies outside [c,c][-c,c]. Hence the entire product is strictly negative
$\left\langle\Delta z_{\rm pen}^{(l)},\;d^{(l)}\right\rangle\;<\;0$
A negative inner product implies that $\Delta z_{\mathrm{pen}}^{(l)}$ points in the exact opposite direction of the clipping residual $d^{(l)}$. Consequently, every out‑of‑range component of pre‑activation $z_{p}^{(l)}$ is pulled back towards the clipping boundary, while pre‑activation already within [c,c][-c,c] experience no change. ∎
Appendix D Training Stability during Polynomial Training
D.1 Pre-Activation Clipping
During early training stages, network weights remain close to their random initialization, and the optimization process does not immediately constrain the pre-activation values to lie within the target approximation interval [c,c][-c,c]. As a result, unbounded pre-activations may arise before the model learns to keep them within range, leading to training instability that can degrade model behavior beyond recovery. To prevent this, we restrict pre-activations to the interval [c,c][-c,c] before evaluating the polynomial activation, which is achieved through a clipping strategy $\mathrm{clip}(z;[-c,c])$ (Diaa et al., 2024):
$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(\mathrm{clip}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x);[-c,c]\right)\right).$
Importantly, this clipping operation is applied only during training to stabilize learning. It is performed after computing the clip-range penalty to ensure that gradients from the regularization term, which encourages the network to keep pre-activations within the approximation interval [c,c][-c,c], are preserved and not masked by the clipping. At inference time, the clipping function is removed:
$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x)\right).$
The model, having learned to constrain pre-activations during training, is expected to remain within the approximation interval without explicit clipping at inference.
D.2 Regularization Warm-up
While activation regularization and pre-activation clipping are both necessary for training stability, applying the full regularization strength from the outset can lead to numerical instability, particularly in larger models. In the early epochs, many pre-activation values lie outside the target interval [c,c][-c,c] across several layers. The clip-range penalty adds a contribution from every such layer, so the penalty term becomes very large. In extreme cases, this can cause the total loss to become numerically undefined.
To address this issue, the regularization strength $\zeta$ is progressively increased over the initial training epochs, which is implemented through a regularization warm up schedule (Diaa et al., 2024). Let $T_{warm}$ denote the total number of warm-up epochs. For each epoch $t$, we define a time-dependent regularization weight $\zeta_{t}$ as:
$\zeta_{t}\;=\;\begin{cases}\alpha_{t}\cdot\zeta&\text{if }t\leq T_{warm}\\ \zeta&\text{if }t>T_{warm}\end{cases},$
where $\{\alpha_{t}\}_{t=1}^{T_{warm}}$ is a fixed sequence of scaling factors satisfying $0<\alpha_{1}<\cdots<\alpha_{T_{warm}}<1$. In practice, we empirically find that setting $T_{warm}=4$ provides a stable convergence. We use a simple predefined sequence of scaling factors: $\alpha=\left\{\frac{1}{100},\frac{1}{50},\frac{1}{10},\frac{1}{5}\right\}$, which produces the epoch-wise regularization strength: $\zeta_{t}\;\in\;\left\{\frac{\zeta}{100},\frac{\zeta}{50},\frac{\zeta}{10},\frac{\zeta}{5},\zeta,\zeta,\dots\right\}$. This warm-up schedule ensures that the regularization penalty is introduced progressively in the initial training epochs without suffering from exploding loss values.
Appendix E Additional Details on Structural Optimizations
E.1 Detailed Derivations of Node Fusing
Batch normalization in linear form is given by
$B(x)=b_{1}x+b_{0},\quad b_{1}=\frac{\gamma}{\sigma},\quad b_{0}=\beta_{b}-b_{1}\mu.$
A quadratic activation $P(x)$ serves as a polynomial substitute for ReLU, where $P(x)$ is given by
$P(x)=c_{2}x^{2}+c_{1}x+c_{0}.$
Case 1: $P(B(x))\mapsto P(x)$. Start with $B(x)=b_{1}x+b_{0}$.
$P(B(x))$ $=c_{2}\big(B(x)\big)^{2}+c_{1}B(x)+c_{0}$
$=c_{2}(b_{1}x+b_{0})^{2}+c_{1}(b_{1}x+b_{0})+c_{0}$
$=c_{2}\big(b_{1}^{2}x^{2}+2b_{1}b_{0}x+b_{0}^{2}\big)+c_{1}b_{1}x+c_{1}b_{0}+c_{0}$
$=(c_{2}b_{1}^{2})x^{2}+\big(2c_{2}b_{1}b_{0}+c_{1}b_{1}\big)x+\big(c_{2}b_{0}^{2}+c_{1}b_{0}+c_{0}\big).$
Resulting polynomial coefficients
$p_{2}=b_{1}^{2}c_{2},\quad p_{1}=b_{1}(2b_{0}c_{2}+c_{1}),\quad p_{0}=b_{0}^{2}c_{2}+b_{0}c_{1}+c_{0}.$
Case 2: $B(C(x))\mapsto C(x)$. Consider a convolution $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$. When batch normalization is applied, it can be rewritten as a rescaled convolution with updated weights and bias.
$B(C(x))$ $=b_{1}C(x)+b_{0}$
$=b_{1}\Big(\sum_{i}w_{i}x_{i}+\beta_{c}\Big)+b_{0}$
$=\sum_{i}(b_{1}w_{i})x_{i}+(b_{1}\beta_{c}+b_{0}).$
The equivalent convolution $\sum_{i}\omega_{i}x_{i}+\alpha$ uses the following updated parameters
$\omega_{i}=b_{1}w_{i},\quad\alpha=b_{1}\beta_{c}+b_{0}.$
Case 3: $P(B_{X}(x)+B_{Y}(y))\mapsto S(x,y)$. Define $B_{X}(x)=b_{X1}x+b_{X0}$ and $B_{Y}(y)=b_{Y1}y+b_{Y0}$. Each of these represents a linearized batch normalization applied to one input variable. Let $z=B_{X}(x)+B_{Y}(y)=b_{X1}x+b_{Y1}y+(b_{X0}+b_{Y0}).$ Applying the activation
$P(z)$ $=c_{2}z^{2}+c_{1}z+c_{0}$
$=c_{2}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)^{2}+c_{1}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)+c_{0}.$
Expanding the square term
$z^{2}$ $=b_{X1}^{2}x^{2}+b_{Y1}^{2}y^{2}+2b_{X1}b_{Y1}xy+2b_{X1}(b_{X0}+b_{Y0})x+2b_{Y1}(b_{X0}+b_{Y0})y+(b_{X0}+b_{Y0})^{2}.$
Substituting and grouping terms by monomials
$P(z)$ $=\underbrace{c_{2}b_{X1}^{2}}_{d_{X2}}x^{2}+\underbrace{c_{2}b_{Y1}^{2}}_{d_{Y2}}y^{2}+\underbrace{2c_{2}b_{X1}b_{Y1}}_{d_{XY}}xy$
$\quad+\underbrace{\big[b_{X1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{X}}x+\underbrace{\big[b_{Y1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{Y}}y$
$\quad+\underbrace{\big[c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}\big]}_{d_{0}}.$
Hence $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$ with the coefficients above.
Case 4: $P(B_{X}(x)+y)\mapsto S(x,y)$. This is the identity shortcut case. Set $B_{X}(x)=b_{X1}x+b_{X0}$ and define the combined input as $z=B_{X}(x)+y=b_{X1}x+y+b_{X0}.$ Applying the activation
$P(z)=c_{2}z^{2}+c_{1}z+c_{0}.$
Expanding the square term
$z^{2}=b_{X1}^{2}x^{2}+y^{2}+2b_{X1}xy+2b_{X1}b_{X0}x+2b_{X0}y+b_{X0}^{2}.$
Substituting and grouping terms by monomials
$P(z)$ $=\underbrace{c_{2}b_{X1}^{2}}_{d_{X2}}x^{2}+\underbrace{c_{2}}_{d_{Y2}}y^{2}+\underbrace{2c_{2}b_{X1}}_{d_{XY}}xy+\underbrace{\big[b_{X1}(2c_{2}b_{X0}+c_{1})\big]}_{d_{X}}x$
$\quad+\underbrace{\big[2c_{2}b_{X0}+c_{1}\big]}_{d_{Y}}y+\underbrace{\big[c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}\big]}_{d_{0}}.$
Hence $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$ with the coefficients above.
E.2 Detailed Derivations of Weight Redistribution
E.2.1 Update Forward
Donors. Average Pooling. Start with $\mu(x)=k^{-1}\sum_{i}{x_{i}}$ and its normalized form $\bar{\mu}(x)=\sum_{i}x_{i}$. We must find $\upsilon$ such that $\upsilon\bar{\mu}(x)=\mu(x)$ is valid
$\upsilon\bar{\mu}(x)$ $=\mu(x)$
$\upsilon\sum_{i}x_{i}$ $=k^{-1}\sum_{i}x_{i}$
$\upsilon$ $=k^{-1}.$
Polynomial Functions. Let $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$ and $\bar{P}(x)=x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}$ be a degree-$d$ polynomial and its normalization, respectively. For $\upsilon\bar{P}(x)=P(x)$ to hold we have $\upsilon$ given by
$\upsilon\bar{P}(x)$ $=P(x)$
$\upsilon\Big(x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}\Big)$ $=\sum_{i=0}^{d}{c_{i}x^{i}}$
$\upsilon x^{d}$ $=c_{d}x^{d}$
$\upsilon$ $=c_{d}.$
The normalized coefficients must satisfy
$\upsilon\bar{P}(x)$ $=P(x)$
$\upsilon\Big(x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}\Big)$ $=\sum_{i=0}^{d}{c_{i}x^{i}}$
$\upsilon\bar{c}_{i}x^{i}$ $=c_{i}x^{i}$
$\bar{c}_{i}$ $=c_{i}\upsilon^{-1}\quad\forall i\in\{0,\dots,d-1\}.$
Bivariate Polynomial. Consider the degree-$d$ bivariate polynomial $S(x,y)=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$, and let $\bar{S}(x,y)=x^{d}+\sum_{i=0}^{d-1}\sum_{j=0}^{d-i}{\bar{c}_{ij}x^{i}y^{j}}$ be its normalization on $x$ (normalizing on $y$ is equivalent). As in the univariate case, we require $\upsilon\bar{S}(x,y)=S(x,y)$. Thus, $\upsilon$ is determined by
$\upsilon\bar{S}(x,y)$ $=S(x,y)$
$\upsilon\Big(x^{d}+\sum_{i=0}^{d-1}\sum_{j=0}^{d-i}{\bar{c}_{ij}x^{i}y^{j}}\Big)$ $=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\upsilon x^{d}$ $=c_{d0}x^{d}$
$\upsilon$ $=c_{d0}.$
The normalized coefficients are obtained through
$\upsilon\bar{S}(x,y)$ $=S(x,y)$
$\upsilon\Big(x^{d}+\sum_{i=0}^{d-1}\sum_{j=0}^{d-i}{\bar{c}_{ij}x^{i}y^{j}}\Big)$ $=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\upsilon\bar{c}_{ij}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-1}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
Receivers. Kernel Functions. Let $K(x)=\sum_{i\in\mathcal{I}}w_{i}x_{i}+\beta$ be the original kernel function and $\bar{K}(x)=\sum_{i\in\mathcal{I}}\bar{w}_{i}x_{i}+\bar{\beta}$ its updated version. To ensure model equivalence, it must hold that $\bar{K}(\upsilon^{-1}x)=K(x)$. Therefore, the updated parameters are given as follows
$\bar{K}(\upsilon^{-1}x)$ $=K(x)$
$\sum_{i\in\mathcal{I}}{\bar{w}_{i}\upsilon^{-1}x_{i}}+\bar{\beta}$ $=\sum_{i\in\mathcal{I}}w_{i}x_{i}+\beta$
$\bar{w}_{i}\upsilon^{-1}x_{i}$ $=w_{i}x_{i}$
$\bar{w}_{i}$ $=w_{i}\upsilon\quad\forall i\in\mathcal{I},\quad\bar{\beta}=\beta.$
Polynomial Functions. For a degree-$d$ polynomial receiver $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$, the coefficients of its updated version $\bar{P}(x)=\sum_{i=0}^{d}{\bar{c}_{i}x^{i}}$ must be chosen such that $\bar{P}(\upsilon^{-1}x)=P(x)$ holds
$\bar{P}(\upsilon^{-1}x)$ $=P(x)$
$\sum_{i=0}^{d}{\bar{c}_{i}(\upsilon^{-1}x)^{i}}$ $=\sum_{i=0}^{d}{c_{i}x^{i}}$
$\bar{c}_{i}\upsilon^{-i}x^{i}$ $=c_{i}x^{i}$
$\bar{c}_{i}$ $=c_{i}\upsilon^{i}\quad\forall i\in\{0,\dots,d\}.$
Bivariate Polynomial. For a degree-$d$ bivariate polynomial $S(x,y)=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$ and its updated version $\bar{S}(x,y)=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{\bar{c}_{ij}x^{i}y^{j}}$,, normalizing with respect to $x$ requires that $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$. Consequently, the coefficients of $\bar{S}(\cdot)$ are
$\bar{S}(\upsilon^{-1}x,y)$ $=S(x,y)$
$\sum_{i=0}^{d}\sum_{j=0}^{d-i}{\bar{c}_{ij}(\upsilon^{-1}x)^{i}y^{j}}$ $=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\bar{c}_{ij}\upsilon^{-i}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{i}\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
E.2.2 Update Backward
Donors. Average Pooling. Start with $\mu(x)=k^{-1}\sum_{i}{x_{i}}$ and its normalization $\bar{\mu}(x)=\sum_{i}x_{i}$. We must determine $\upsilon$ such that $\bar{\mu}(\upsilon x)=\mu(x)$ holds
$\bar{\mu}(\upsilon x)$ $=\mu(x)$
$\sum_{i}\upsilon x_{i}$ $=k^{-1}\sum_{i}x_{i}$
$\upsilon\sum_{i}x_{i}$ $=k^{-1}\sum_{i}x_{i}$
$\upsilon$ $=k^{-1}.$
Polynomial Functions. For a degree-$d$ polynomial donor $P(\cdot)$, the equality $\bar{P}(\upsilon x)=P(x)$ must be satisfied, where $\bar{P}(\cdot)$ is the normalized version. The appropriate $\upsilon$ is found from
$\bar{P}(\upsilon x)$ $=P(x)$
$(\upsilon x)^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}(\upsilon x)^{i}}$ $=\sum_{i=0}^{d}{c_{i}x^{i}}$
$\upsilon^{d}x^{d}$ $=c_{d}x^{d}$
$\upsilon$ $=c_{d}^{1/d}.$
and the normalized coefficients are
$\bar{P}(\upsilon x)$ $=P(x)$
$(\upsilon x)^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}(\upsilon x)^{i}}$ $=\sum_{i=0}^{d}{c_{i}x^{i}}$
$\bar{c}_{i}\upsilon^{i}x^{i}$ $=c_{i}x^{i}$
$\bar{c}_{i}$ $=c_{i}\upsilon^{-i}\quad\forall i\in\{0,\dots,d-1\}.$
Bivariate Polynomial. Assuming normalization along $x$, we require $\bar{S}(\upsilon x,y)=S(x,y)$ for the degree-$d$ bivariate polynomial $S(x,y)$ and its normalization $\bar{S}(x,y)$. For that, $\upsilon$ is determined from
$\bar{S}(\upsilon x,y)$ $=S(x,y)$
$(\upsilon x)^{d}+\sum_{i=0}^{d-1}\sum_{j=0}^{d-i}{\bar{c}_{ij}(\upsilon x)^{i}y^{j}}$ $=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\upsilon^{d}x^{d}$ $=c_{d0}x^{d}$
$\upsilon$ $=c_{d0}^{1/d}.$
with the normalized coefficients being
$\bar{S}(\upsilon x,y)$ $=S(x,y)$
$(\upsilon x)^{d}+\sum_{i=0}^{d-1}\sum_{j=0}^{d-i}{\bar{c}_{ij}(\upsilon x)^{i}y^{j}}$ $=\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\bar{c}_{ij}\upsilon^{i}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-i}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
Receivers. Kernel Functions. Let $K(x)=\sum_{i\in\mathcal{I}}w_{i}x_{i}+\beta$ and $\bar{K}(x)=\sum_{i\in\mathcal{I}}\bar{w}_{i}x_{i}+\bar{\beta}$ be a kernel function and its updated version. To preserve model equivalence, we require $\bar{K}(x)=\upsilon K(x)$, thus
$\bar{K}(x)$ $=\upsilon K(x)$
$\sum_{i\in\mathcal{I}}{\bar{w}_{i}x_{i}}+\bar{\beta}$ $=\upsilon\Big(\sum_{i\in\mathcal{I}}w_{i}x_{i}+\beta\Big)$
$\bar{w}_{i}x_{i}$ $=\upsilon w_{i}x_{i}$
$\bar{w}_{i}$ $=w_{i}\upsilon\quad\forall i\in\mathcal{I},\quad\bar{\beta}=\beta\upsilon.$
Polynomial Functions. For a degree-$d$ polynomial receiver $P(x)=\sum_{i=0}^{d}c_{i}x^{i}$, the updated polynomial $\bar{P}(x)=\sum_{i=0}^{d}\bar{c}_{i}x^{i}$ must satisfy $\bar{P}(x)=\upsilon P(x)$. This yields
$\bar{P}(x)$ $=\upsilon P(x)$
$\sum_{i=0}^{d}{\bar{c}_{i}x^{i}}$ $=\upsilon\sum_{i=0}^{d}{c_{i}x^{i}}$
$\bar{c}_{i}x^{i}$ $=\upsilon c_{i}x^{i}$
$\bar{c}_{i}$ $=c_{i}\upsilon\quad\forall i\in\{0,\dots,d\}.$
Bivariate Polynomial. For a degree-$d$ bivariate polynomial $S(x,y)=\sum_{i=0}^{d}\sum_{j=0}^{d-i}c_{ij}x^{i}y^{j}$ and its updated version $\bar{S}(x,y)=\sum_{i=0}^{d}\sum_{j=0}^{d-i}\bar{c}_{ij}x^{i}y^{j}$, we require $\bar{S}(x,y)=\upsilon S(x,y)$. Therefore
$\bar{S}(x,y)$ $=\upsilon S(x,y)$
$\sum_{i=0}^{d}\sum_{j=0}^{d-i}{\bar{c}_{ij}x^{i}y^{j}}$ $=\upsilon\sum_{i=0}^{d}\sum_{j=0}^{d-i}{c_{ij}x^{i}y^{j}}$
$\bar{c}_{ij}x^{i}y^{j}$ $=\upsilon c_{ij}x^{i}y^{j}$
$\bar{c}_{ij}$ $=c_{ij}\upsilon\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
E.3 Example of Tower Reuse
Suppose following an activation function, there is a convolution. Let modulus $q_{i}\approx\Delta^{2}$, thus $\lambda(q_{i})=2$. Consider an input $x$ with scaling factor $\Delta_{x}=\Delta$, giving $\lambda(x)=1$. Applying a quadratic activation $y=x^{2}+c_{1}x+c_{0}$ yields $\lambda(y)=2$, with $\lambda(c_{1})=1$ and $\lambda(c_{0})=2$. Next, a convolution $z=\sum_{i}\omega_{i}y_{i}+\alpha$ is performed. Since $\lambda(y)=2$, by setting $\lambda(\omega)=1$, we obtain $\lambda(z)=3$. As $\lambda(z)>\lambda(q_{i})$, a rescale is triggered. The rescale operation drops $q_{i}$ from the modulus chain, reducing the level $\Lambda(\cdot)$ by one and resetting the sublevel to one:
$\Lambda(\downarrow z)=\Lambda(z)-1,\qquad\lambda(\downarrow z)=\lambda(z)-\lambda(q_{i})=1.$
In terms of $\Delta$, a ciphertext $z$ with $\Delta_{z}=\Delta^{3}$ rescaled by $q_{i}\approx\Delta^{2}$ results in output scale $\Delta_{\downarrow z}\approx\Delta$.
E.4 Level Analysis
Table 1: Number of levels required for multiplications in ResNet models. $P_{4}$: Degree-4 polynomial activation used in PILLAR (Diaa et al., 2024). $P_{2}$: Degree-2 polynomial activation (§2). $P_{2}F$: $P_{2}$ with node fusing (§3.1). $P_{2}R$: $P_{2}$ with weight redistribution (§3.2). $P_{2}FR$: $P_{2}$ with node fusing and weight redistribution. $P_{2}FRT$: $P_{2}FR$ with tower reuse (§3.3).
Model $P_{4}$ $P_{2}$ $P_{2}F$ $P_{2}R$ $P_{2}FR$ $P_{2}FRT$
ResNet-18 87 70 53 35 35 18
ResNet-20 97 78 59 39 39 20
ResNet-32 157 126 95 63 63 32
Table 1 summarizes the number of levels $\mathcal{L}$ required for multiplication across different ResNet variants under various optimization techniques. To illustrate, we focus on ResNet-18, which comprises 17 convolutional layers, batch normalization layers, and activation functions, along with a single average pooling and linear layer:
  • $P_{4}$: Prior work achieving the lowest $\mathcal{L}$ (Diaa et al., 2024) uses a degree-4 polynomial as activation function. This results in $\mathcal{L}=87$, with each convolution, batch normalization, average pooling, and linear layer contributing $\mathcal{L}=1$, and the polynomial activation $\mathcal{L}=3$.
  • $P_{2}$: By replacing the activation function with a degree-2 polynomial, as described in §2, the required level for activations drops to 2, reducing the model $\mathcal{L}$ to 70.
  • $P_{2}F$: Applying node fusing (§3.1) to $P_{2}$ removes all batch normalizations, lowering $\mathcal{L}$ to 53.
  • $P_{2}R$: Alternatively, applying weight redistribution (§3.2) to $P_{2}$ yields $\mathcal{L}=35$. This is achieved because the highest-order coefficients in batch normalization and polynomial activations are normalized to one, simplifying them to $x+b_{0}$ ($\mathcal{L}=0$) and $x^{2}+c_{1}x+c_{0}$ ($\mathcal{L}=1$), respectively. Moreover, the divisor in average pooling is also normalized to one, giving it $\mathcal{L}=0$. These simplifications collectively result in $\mathcal{L}=35$.
  • $P_{2}FR$: Combining node fusing and weight redistribution retains $\mathcal{L}=35$. However, it requires fewer homomorphic operations overall, since node fusing eliminates batch normalizations entirely.
  • $P_{2}FRT$: Finally, applying tower reuse (§3.3) to $P_{2}FR$ allows a non-linear polynomial activation followed by a linear kernel function to share the same level. This halves the total levels to $\mathcal{L}=18$, following $\mathcal{L}_{P_{2}FRT}=\lceil\frac{\mathcal{L}_{P_{2}FR}}{2}\rceil$ for ResNet models.
In summary, our techniques reduce levels by nearly a factor of five compared to prior work.
Appendix F Illustration of Slice-wise Clustering
F.1 Single Model Slice Clustering
Figure 3 illustrates how convolution filters are divided into slices along the width axis $W_{l}$, with each colored slice $S^{(l)}_{s}$ clustered independently using its own codebook. Each 3D block in the figure corresponds to a convolution filter associated with one output channel. The three axes correspond to kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth). The collection of such filters forms the output channels $O_{l}$. Slices are taken as vertical strips along the width axis $W_{l}$, and the figure highlights them with distinct colors. The blue strip corresponds to slice $S^{(l)}_{1}$, the yellow strip to slice $S^{(l)}_{2}$, and the pink strip to slice $S^{(l)}_{3}$. Each slice $S^{(l)}_{s}$ groups together all weights $W^{(l)}_{o,i,h,s}$} across output channels $O_{l}$, input channels $I_{l}$, and kernel height $H_{l}$ at a fixed width position $s$, and is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}$. This illustration emphasizes that clustering is performed slice by slice, rather than across the entire filter volume.
Figure 3: Illustration of single model slice-wise clustering.
F.2 Ensemble Slice Clustering
Figure 4 illustrates how slice-wise clustering is extended from a single model to an ensemble of polynomial networks.
Figure 4: Illustration of ensemble slice-wise clustering.
Each 3D block corresponds to the convolution filters of one model instance, spanning kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth). The collection of such filters across output channels $O_{l}$ forms the layer weights $W^{(l,m)}$ for model $m$. For a fixed layer $l$, slices are again taken as vertical strips along the width axis $W_{l}$, with colors indicating different slices $S^{(l,m)}_{s}$ within each model. Unlike the single-model case, however, we now align slices from multiple models at the same width position $s$ and stack them together. This produces a slice matrix $X^{(l)}_{s}\in\mathbb{R}^{N_{s}\times M}$, where each row corresponds to one weight coordinate across all $M$ models, and $N_{s}=O_{l}I_{l}H_{l}$. Clustering is performed jointly in $\mathbb{R}^{M}$, producing a shared codebook $\mathcal{C}^{(l)}_{s}$ for slice $s$. The illustration emphasizes that ensemble clustering is not applied to each model independently. Instead, slices are stacked across models at the same spatial position, and clustering is performed jointly.
Appendix G Ablation Study
G.1 Ablation on Clipping Range (Value of $c$)
We conduct an ablation study to evaluate the effect of the clipping interval [c,c][-c,c] on model performance. The choice of $c$ determines how much of the activation distribution is preserved versus truncated: if $c$ is too small, informative activations are excessively clipped, which can harm representational capacity, while if $c$ is too large, the clipping mechanism provides little stabilization benefit, allowing unstable activations to propagate. Figure 5 shows the results for CIFAR-10 with ResNet-18.
Figure 5: Results of the clipping range ablation study on CIFAR-10 with ResNet-18. (a) Distribution of pre-activation inputs to ReLU across the entire network, showing concentration within [2,2][-2,2]. (b) Test accuracy under different clipping intervals, where [2,2][-2,2] produces the best performance.
Figure 5(a) presents the distribution of pre-activation inputs to ReLU across the entire network, where the values are concentrated within the interval [2,2][-2,2] and only a small fraction lies outside this range, supporting the choice of restricting pre-activations to [2,2][-2,2]. Figure 5(b) reports the accuracy obtained under different clipping intervals. While several ranges were tested, [2,2][-2,2] consistently achieved the highest mean accuracy across runs. Smaller intervals degraded performance by removing too many activations, whereas larger intervals reduced stability and resulted in lower accuracy. The same trend holds across other dataset–architecture combinations. Based on these observations, we adopt [2,2][-2,2] as the clipping range in all our experiments.
G.2 Ablation on Penalty Strength (Value of $\zeta$)
We conduct an ablation study to evaluate the effect of the penalty strength $\zeta$ on model performance. The parameter $\zeta$ controls the magnitude of the regularization term that penalizes pre-activations outside the target interval [c,c][-c,c]. Figure 6 shows the results for CIFAR-10 with ResNet-18.
Figure 6: Results of the penalty strength ablation study on CIFAR-10 with ResNet-18, showing test accuracy across different values of $\zeta$. The best performance is achieved at $\zeta=10^{-3}$; smaller values fail to regularize effectively, while larger values start to decrease accuracy.
The highest accuracy is obtained for $\zeta=10^{-3}$. Values smaller than $10^{-3}$ do not provide sufficient regularization and result in poor accuracy, while values greater than $10^{-3}$ start decreasing performance, indicating that excessive penalization constrains the model too strongly. The same trend holds across other dataset–architecture combinations. Based on these results, we set $\zeta=10^{-3}$ in all our experiments.
G.3 Comparison with PILLAR
We evaluate PILLAR using its open-source implementation (git, 2024b). PILLAR is designed for interactive PPML, which is a different use case from our non-interactive setting, and its repository does not provide code for private inference. For these reasons, we compare only in terms of accuracy and do not evaluate inference time. Table 2 reports accuracies (mean $\pm$ standard deviation) for ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using ReLU, PAPER (our method), and two versions of PILLAR. The original PILLAR is defined with a degree-4 polynomial. For a fair comparison, we additionally adapted this implementation to use a degree-2 polynomial, allowing a direct comparison to PAPER. On average, the adapted degree-2 version shows an accuracy drop of 6.5% with very high variance, particularly on CIFAR-100. The original degree-4 version is more stable but still falls short of ReLU by about 5.4% on average. In contrast, PAPER with degree-2 consistently stays much closer to ReLU, with only about a 1.3% drop on average. Importantly, PAPER remains stable across both datasets and all ResNet variants, avoiding the large fluctuations seen in PILLAR. These results demonstrate that while PAPER can exploit low-degree polynomials effectively, PILLAR either collapses when reduced to degree 2 or, in its intended degree-4 form, does not reach comparable accuracy.
Table 2: Classification accuracies (mean $\pm$ standard deviation) on CIFAR-10 and CIFAR-100 for ResNet-18, ResNet-20, and ResNet-32 using ReLU, PAPER (degree-2), and two versions of PILLAR. The original PILLAR method is degree-4 (Diaa et al., 2024), while the degree-2 variant is our adaptation of its open-source implementation for comparison.
Model Dataset ReLU PAPER PILLAR ($d=2$) PILLAR ($d=4$)
ResNet-18 CIFAR-10 94.57 $\pm$ 0.23 93.73 $\pm$ 0.27 91.78 $\pm$ 0.13 93.26 $\pm$ 0.15
CIFAR-100 76.22 $\pm$ 0.44 74.96 $\pm$ 0.51 71.92 $\pm$ 0.28 74.81 $\pm$ 0.22
ResNet-20 CIFAR-10 94.84 $\pm$ 0.25 94.06 $\pm$ 0.26 90.43 $\pm$ 0.40 90.98 $\pm$ 0.29
CIFAR-100 74.96 $\pm$ 0.36 73.89 $\pm$ 0.37 62.30 $\pm$ 3.96 62.55 $\pm$ 7.51
ResNet-32 CIFAR-10 95.23 $\pm$ 0.30 93.86 $\pm$ 0.24 90.40 $\pm$ 1.73 92.24 $\pm$ 0.27
CIFAR-100 76.58 $\pm$ 0.40 73.93 $\pm$ 0.53 66.38 $\pm$ 1.78 66.10 $\pm$ 1.12
Appendix H Additional Experimental Results
H.1 Additional Details of FHE Parameters
For each model, we detail the encryption parameters used in our experiments. Our framework automatically determines the RNS moduli according to the specified sizes $|q_{i}|$. We chose the smallest values of $\Delta$ and $|q_{i}|$ that preserve accuracy during private inference and adjust $N$ for $\approx$128-bit security. The moduli are ordered as $q_{0},\dots,q_{L},P$, where $q_{0}$ holds the output, $q_{1},\ldots,q_{L}$ are for rescaling, and $P$ is for modulus expansion in relinearization. FHE terminology is provided in §A.
ResNet-18: $N=2^{15}$, $\Delta=2^{22}$, $\log_{2}{Q}=18\cdot 44+23+54=869$. Moduli: {0x100000020001, 0x100000050001, 0x100000090001, 0x1000000b0001, 0xfffffcf0001, 0x100000180001, 0x1000001a0001, 0xfffffc60001, 0x1000002c0001, 0xfffffb70001, 0x1000002d0001, 0x1000003c0001, 0xfffffb50001, 0x1000003e0001, 0xfffffaf0001, 0x100000480001, 0xfffffac0001, 0x100000570001, 0x820001, 0x3fffffffd60001}.
ResNet-20: $N=2^{15}$, $\Delta=2^{21}$, $\log_{2}{Q}=20\cdot 42+22+44=906$. Moduli: {0x400000b0001, 0x3ffffe80001, 0x400002f0001, 0x3ffffd20001, 0x40000330001, 0x3ffffca0001, 0x40000390001, 0x3ffffc30001, 0x400003b0001, 0x3ffffbe0001, 0x400004d0001, 0x3ffff850001, 0x40000560001, 0x400005c0001, 0x3ffff7b0001, 0x400006c0001, 0x3ffff550001, 0x40000770001, 0x3ffff4f0001, 0x400007a0001, 0x390001, 0x100000020001}.
ResNet-32: $N=2^{16}$, $\Delta=2^{26}$, $\log_{2}{Q}=32\cdot 52+27+54=1745$. Moduli: {0x10000000060001, 0xffffffff00001, 0x10000000180001, 0xfffffffe40001, 0x10000000200001, 0xfffffffe20001, 0x100000003e0001, 0xfffffffbe0001, 0x10000000500001, 0xfffffffa60001, 0x100000006e0001, 0xfffffff820001, 0x100000007e0001, 0xfffffff480001, 0x10000000960001, 0xfffffff280001, 0x10000000c80001, 0x10000000d80001, 0xffffffed60001, 0x10000000ec0001, 0xffffffec40001, 0x10000000fc0001, 0xffffffeb00001, 0x100000010e0001, 0xffffffe9e0001, 0x10000001380001, 0xffffffe9a0001, 0x100000016a0001, 0xffffffe940001, 0x10000001bc0001, 0xffffffe6a0001, 0x10000001be0001, 0x8020001, 0x3fffffffd60001}.
H.2 Results on ResNet18
Figure 7: Accuracy vs. private inference time for ResNet-18 on CIFAR-10/100. Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model). We compare Standard ensembles (no clustering), Slice ensembles (slice-wise clustering), and Full Clustering. Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering. The green dashed line shows plaintext ReLU accuracy.
Figure 7 shows the trade-offs between accuracy and inference time for ResNet-18 on CIFAR-10 and CIFAR-100. Results show similar traits to ResNet-20, leading to similar conclusions. In the clustering plots (left column), both Full and Slice Clustering reduce inference time compared to Standard. At small centroid counts ($k\leq 32$), accuracy drops noticeably due to a lack of representative centroids. Slice Clustering consistently achieves higher accuracy than Full Clustering, benefiting from its finer-grained parameter representation. When the centroid count is sufficiently large ($k\geq 64$), Slice Clustering reaches the accuracy of Standard while still reducing latency substantially.
The middle column plots compare Standard and Slice Clustering ensembles with $M\in\{1,2,4\}$. Standard gains accuracy as $M$ increases, while Slice Clustering shows more variable behavior because each centroid encodes an $M$-dimensional parameter space; e.g. Slice-4 underperforms relative to Slice-2 for small $k$. Slice Clustering provides a regularization benefit and, with larger centroid counts, Slice-2 achieves higher accuracy than Standard-2 and nearly matches Standard-4.
The right column presents the Pareto front of Slice Clustering and compares it to Standard-4. Slice Clustering shifts the curve toward lower latency while staying close to Standard in accuracy. On both datasets, Slice Clustering fully recovers plaintext ReLU accuracy and sometimes even surpasses it, demonstrating that accuracy can be preserved while reducing inference time.
H.3 Experimental Results
Tables 3, 4, and 5 report the accuracy, inference time, and peak memory usage of ResNet-18, ResNet-20, and ResNet-32, respectively, on the CIFAR-10 and CIFAR-100 datasets using the Standard, Full Clustering, and Slice Clustering methods. For Full Clustering and Slice Clustering, we varied the number of centroids as $k=2^{i}\ \forall i\in\{1,\dots,10\}$. For Standard and Slice Clustering, we additionally evaluated ensemble models with $M\in{1,2,4}$. In the case of Slice Clustering with $M\geq 2$, we extended the centroid range to $k=2^{i}\ \forall i\in\{1,\dots,12\}$ in ResNet-18 and ResNet-20. We did not extend this configuration to ResNet-32 due to system memory limitations (§5.1).
Table 3: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-18.
Standard Standard
M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
1 93.73 0.27 3925 51 292 1 74.96 0.51 3935 79 338
2 94.34 0.16 3930 54 314 2 76.48 0.28 3874 40 355
4 94.69 0.12 3840 44 347 4 77.43 0.24 3929 54 390
Full Clustering (M = 1) Full Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 10.16 0.72 404 16 197 2 1.00 0.07 414 25 240
4 10.14 0.89 406 18 200 4 1.02 0.12 413 17 242
8 9.91 1.48 420 31 202 8 1.08 0.17 426 23 244
16 15.89 7.51 427 27 204 16 2.01 1.10 430 25 244
32 70.18 14.19 439 28 204 32 32.33 13.66 441 23 246
64 91.42 1.25 481 31 204 64 68.01 4.00 484 34 246
128 93.24 0.29 530 34 205 128 72.62 1.04 554 41 247
256 93.53 0.26 619 36 209 256 73.98 0.68 646 44 250
512 93.65 0.25 719 40 221 512 74.50 0.70 773 53 263
1024 93.69 0.28 856 34 245 1024 74.65 0.69 926 58 286
Slice Clustering (M = 1) Slice Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 9.79 0.76 418 32 198 2 1.07 0.17 426 16 239
4 12.66 3.02 409 23 200 4 2.76 1.17 413 13 242
8 81.28 6.04 412 23 203 8 42.21 13.71 420 24 244
16 92.25 0.68 420 19 204 16 69.28 2.37 443 25 246
32 93.36 0.33 456 28 204 32 72.85 0.92 466 27 245
64 93.62 0.27 504 30 204 64 74.31 0.77 520 31 246
128 93.70 0.26 586 30 205 128 74.61 0.73 602 36 248
256 93.70 0.28 696 34 214 256 74.68 0.71 724 39 256
512 93.73 0.27 855 58 232 512 74.69 0.73 874 49 274
1024 93.72 0.27 1016 57 266 1024 74.71 0.74 1066 66 308
Slice Clustering (M = 2) Slice Clustering (M = 2)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 92.59 0.45 520 43 222 64 68.93 1.94 524 10 263
128 93.85 0.15 597 38 222 128 74.47 0.59 622 23 263
256 94.35 0.10 695 22 230 256 75.79 0.43 734 51 273
512 94.50 0.09 854 28 250 512 76.35 0.18 871 37 292
1024 94.60 0.08 1045 61 284 1024 76.67 0.12 1049 23 326
2048 94.73 0.07 1164 36 394 2048 76.82 0.14 1202 72 435
4096 94.78 0.04 1268 42 643 4096 76.90 0.08 1308 18 681
8192 94.77 0.05 1482 59 1148 8192 76.97 0.07 1440 13 1174
Slice Clustering (M = 4) Slice Clustering (M = 4)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 18.38 4.80 526 27 256 64 1.73 0.41 538 26 297
128 35.79 8.02 621 45 258 128 5.31 0.94 615 27 299
256 66.88 5.19 728 51 265 256 29.69 2.16 750 47 308
512 84.28 2.38 866 28 285 512 57.88 2.20 885 16 326
1024 91.47 0.46 1035 69 319 1024 68.98 1.13 1057 46 360
2048 93.10 0.21 1211 57 428 2048 73.96 0.38 1195 64 469
4096 94.11 0.14 1298 17 679 4096 75.93 0.22 1263 24 716
8192 94.55 0.11 1509 22 1183 8192 76.66 0.12 1355 17 1208
Table 4: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-20.
Standard Standard
M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
1 94.06 0.26 1886 30 279 1 73.89 0.37 1902 26 305
2 94.71 0.11 1884 41 300 2 76.59 0.25 1874 27 320
4 95.04 0.10 1886 32 330 4 77.97 0.13 1904 30 351
Full Clustering (M = 1) Full Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 10.08 0.26 242 10 175 2 1.03 0.07 246 6 208
4 10.08 0.32 246 6 190 4 1.08 0.23 246 9 211
8 10.17 0.89 250 12 192 8 1.17 0.24 252 7 212
16 14.44 6.34 260 13 192 16 2.05 1.58 258 8 213
32 78.18 8.99 264 7 193 32 32.59 14.07 272 16 214
64 92.46 0.95 288 10 193 64 68.18 3.24 290 11 214
128 93.77 0.37 334 12 194 128 71.56 1.19 334 15 214
256 94.00 0.26 407 22 198 256 73.39 0.49 392 9 218
512 94.02 0.26 475 18 212 512 73.75 0.35 473 20 232
1024 94.05 0.26 553 25 248 1024 73.85 0.35 563 28 264
Slice Clustering (M = 1) Slice Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 10.05 0.41 249 18 186 2 1.12 0.22 254 10 207
4 14.59 4.33 244 7 190 4 2.11 0.62 255 13 210
8 85.17 2.66 247 7 192 8 49.94 5.99 264 21 212
16 93.20 0.45 262 6 193 16 69.60 1.71 262 7 214
32 93.87 0.32 276 9 194 32 72.98 0.45 292 12 214
64 94.01 0.28 314 9 194 64 73.68 0.38 323 13 214
128 94.01 0.28 366 14 194 128 73.85 0.38 377 18 215
256 94.04 0.27 442 10 205 256 73.89 0.34 452 17 226
512 94.05 0.26 536 26 226 512 73.89 0.36 538 10 247
1024 94.05 0.26 612 19 288 1024 73.88 0.36 626 21 309
Slice Clustering (M = 2) Slice Clustering (M = 2)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 92.52 0.56 312 9 209 64 68.98 0.97 321 14 230
128 94.34 0.10 370 8 211 128 74.31 0.44 380 16 230
256 94.81 0.08 443 20 219 256 76.08 0.25 464 7 241
512 94.93 0.08 531 31 242 512 76.65 0.19 541 22 263
1024 95.08 0.07 611 14 304 1024 76.99 0.11 614 22 324
2048 95.11 0.06 679 7 458 2048 77.16 0.13 690 23 476
4096 95.08 0.04 814 60 768 4096 77.12 0.12 746 17 781
8192 95.08 0.03 885 37 1389 8192 77.23 0.09 897 26 1390
Slice Clustering (M = 4) Slice Clustering (M = 4)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 12.79 2.88 332 14 240 64 1.72 0.48 341 13 261
128 25.95 4.81 389 7 242 128 3.86 0.96 407 10 262
256 57.23 3.93 455 21 251 256 19.97 2.42 463 19 271
512 83.18 1.59 534 10 274 512 52.26 2.81 546 8 294
1024 91.45 0.46 603 18 335 1024 68.37 0.63 631 14 356
2048 93.89 0.26 648 9 489 2048 73.73 0.53 673 11 508
4096 94.65 0.16 722 12 799 4096 76.44 0.31 765 9 813
8192 94.90 0.10 815 17 1421 8192 77.62 0.17 855 16 1421
Table 5: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-32.
Standard Standard
M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
1 93.86 0.24 9405 109 1332 1 73.93 0.53 8865 132 1401
2 94.69 0.18 8854 80 1394 2 76.79 0.34 8902 152 1418
4 94.99 0.13 8813 63 1441 4 78.34 0.23 8907 141 1465
Full Clustering (M = 1) Full Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 10.12 0.22 1578 63 842 2 0.98 0.07 1544 5 889
4 10.13 0.21 1538 3 867 4 0.98 0.07 1577 20 908
8 10.12 0.22 1588 79 879 8 0.97 0.07 1540 24 914
16 10.35 1.35 1570 8 893 16 1.03 0.21 1547 27 929
32 37.13 16.87 1578 3 898 32 3.91 2.63 1565 13 942
64 86.61 3.69 1703 80 901 64 52.92 7.49 1686 44 945
128 92.82 0.47 2131 27 902 128 70.26 1.24 1819 5 947
256 93.70 0.32 2164 151 916 256 73.18 0.38 2013 33 954
512 93.84 0.27 2495 164 987 512 73.75 0.37 2381 70 1020
1024 93.86 0.26 2898 174 1155 1024 73.89 0.33 2822 53 1159
Slice Clustering (M = 1) Slice Clustering (M = 1)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
2 10.12 1.07 1605 57 857 2 0.99 0.08 1559 11 894
4 10.00 0.99 1573 24 879 4 1.00 0.09 1526 27 920
8 34.61 12.19 1601 65 893 8 9.67 5.64 1586 11 936
16 89.08 2.44 1559 6 901 16 60.50 4.61 1576 8 942
32 93.12 0.49 1761 90 902 32 71.75 0.80 1679 19 946
64 93.75 0.29 1833 54 905 64 73.49 0.36 1803 9 947
128 93.84 0.26 2077 111 905 128 73.83 0.33 2027 30 946
256 93.86 0.27 2488 185 950 256 73.91 0.33 2379 83 993
512 93.85 0.25 2849 128 1060 512 73.93 0.33 2815 60 1104
1024 93.86 0.24 3317 192 1373 1024 73.94 0.53 3313 57 1417
Slice Clustering (M = 2) Slice Clustering (M = 2)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 85.12 4.61 1831 69 927 64 52.29 6.60 1791 17 968
128 92.75 0.58 2069 79 933 128 70.20 1.26 2014 44 972
256 94.06 0.21 2388 156 977 256 74.51 0.42 2340 52 1019
512 94.59 0.12 2788 167 1087 512 76.30 0.25 2809 38 1128
1024 94.76 0.12 3093 11 1399 1024 76.77 0.24 3238 64 1440
Slice Clustering (M = 4) Slice Clustering (M = 4)
k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
64 9.27 0.65 1845 89 979 64 0.92 0.00 1817 22 1019
128 14.20 2.55 2096 100 983 128 0.92 0.01 2031 26 1022
256 30.01 9.08 2409 149 1030 256 1.79 0.74 2367 50 1068
512 69.18 5.72 2842 193 1136 512 16.68 6.66 2823 47 1178
1024 88.24 1.02 3379 64 1450 1024 47.96 5.14 3495 15 1492

### v2

@@
-PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE
+PAPER: Privacy-Preserving Convolutional Neural Networks using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE
 Abstract
-Recent work has made non-interactive privacy-preserving inference more practical by running deep Convolution Neural Network (CNN) with Fully Homomorphic Encryption (FHE).
-However, these methods remain limited by their reliance on bootstrapping, a costly FHE operation applied across multiple layers, severely slowing inference.
-They also depend on high-degree polynomial approximations of non-linear activations, which increase multiplicative depth and reduce accuracy by 2–5% compared to plaintext ReLU models.
-In this work, we focus on ResNets, a widely adopted benchmark architecture in privacy-preserving inference, and close the accuracy gap between their FHE-based non-interactive models and plaintext counterparts, while also achieving faster inference than existing methods.
-We use a quadratic polynomial approximation of ReLU, which achieves the theoretical minimum multiplicative depth for non-linear activations, along with a penalty-based training strategy.
-We further introduce structural optimizations such as node fusing, weight redistribution, and tower reuse.
-These optimizations reduce the required FHE levels in CNNs by nearly a factor of five compared to prior work, allowing us to run ResNet models under leveled FHE without bootstrapping.
-To further accelerate inference and recover accuracy typically lost with polynomial approximations, we introduce parameter clustering along with a joint strategy of data encoding layout and ensemble techniques.
-Experiments with ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 show that our approach achieves up to $4\times$ faster private inference than prior work with comparable accuracy to plaintext ReLU models.
+Recent work using Fully Homomorphic Encryption (FHE) has made non-interactive privacy-preserving inference of deep Convolutional Neural Networks (CNN) possible.
+However, the performance of these methods remain limited by their heavy reliance on bootstrapping, a costly FHE operation applied across multiple layers, severely slowing inference.
+Moreover, they depend on high-degree polynomial approximations of non-linear activations, which increase multiplicative depth and reduce accuracy by 2–5% compared to plaintext ReLU models.
+In this work, we close the accuracy gap between FHE-based non-interactive CNNs and their plaintext counterparts, while also achieving faster inference than existing methods.
+We propose a quadratic polynomial approximation of ReLU, which achieves the theoretical minimum multiplicative depth for non-linear activations, together with a penalty-based training strategy.
+We further introduce structural optimizations that reduce the required FHE levels in CNNs by a factor of five compared to prior work, allowing us to run deep CNN models under leveled FHE without bootstrapping.
+To further accelerate inference and recover accuracy typically lost with polynomial approximations, we introduce parameter clustering along with a joint strategy of data layout and ensemble techniques.
+Experiments with VGG and ResNet models on CIFAR and Tiny-ImageNet datasets show that our approach achieves up to $4\times$ faster private inference than prior work, with accuracy comparable to plaintext ReLU models.
 1 Introduction
-Machine Learning as a Service (MLaaS) is increasingly adopted across industries as it provides access to powerful models without the need for in-house development or maintenance (Ribeiro et al., 2015).
-However, it raises serious privacy concerns since models are often trained on proprietary or sensitive data (Anwar et al., 2018; Özbayoglu et al., 2020), and inference requires clients to share private information (e.g., medical or financial records) with external service providers.
-Privacy-Preserving Machine Learning (PPML) addresses these risks using cryptographic methods, mainly secure Multi-Party Computation (MPC) (Zhou et al., 2024) and Fully Homomorphic Encryption (FHE) (Podschwadt et al., 2022), which protect both client data and provider models.
-Broadly, PPML methods fall into interactive and non-interactive categories (Lee et al., 2022).
-Interactive PPML, based on MPC, requires the client and the service provider to jointly perform inference (Liu et al., 2017; Juvekar et al., 2018; Mishra et al., 2020; Lou et al., 2021; Huang et al., 2022; Diaa et al., 2024).
-It keeps computational costs relatively low but requires multiple communication rounds, leading to higher communication and bandwidth requirements that limit applicability.
-Non-interactive PPML, typically using FHE, adopts a fire-and-forget paradigm (Gilad-Bachrach et al., 2016; Brutzkus et al., 2019; Lou & Jiang, 2021; Lee et al., 2022; Sarkar et al., 2023; Ao & Boddeti, 2024), where the client encrypts inputs, the provider computes on ciphertexts, and the client decrypts the output.
+Machine Learning as a Service (MLaaS) is increasingly adopted across industries because it provides access to powerful models without requiring in-house development or maintenance [30].
+However, it raises serious privacy concerns since models are often trained on proprietary or sensitive data [4, 28], and inference requires clients to share private information (e.g., medical or financial records) with external service providers.
+Privacy-Preserving Machine Learning (PPML) addresses these risks using predominantly cryptographic methods, mainly secure Multi-Party Computation (MPC) [34] and Fully Homomorphic Encryption (FHE) [29], which protect both client data and provider models.
+Broadly, PPML methods fall into interactive and non-interactive categories [20].
+Interactive PPML, based on MPC, requires the client and the service provider to jointly perform inference [21, 16, 27, 24, 14, 11].
+It keeps computational costs relatively low but requires multiple communication rounds, which leads to high communication and bandwidth requirements.
+Non-interactive PPML, typically using FHE, adopts a fire-and-forget paradigm [13, 6, 23, 20, 31, 5], where the client encrypts inputs, the provider computes on ciphertexts, and the client decrypts the output.
 This reduces communication to a single round at the expense of heavy server-side computation.
-Despite the costs, non-interactive PPML is well-suited to third-party services as it eliminates client involvement during inference, supports low-resource devices, and works in low-bandwidth settings.
+Despite the costs, non-interactive PPML is well-suited to third-party services because it eliminates client involvement during inference, supports clients with low-resource devices, and works in low-bandwidth settings.
 Motivated by these advantages, this work focuses on non-interactive PPML.
 A key challenge in non-interactive PPML is implementing activation functions with cryptographic primitives.
-Activations are non-polynomial, while FHE supports only polynomial operations like addition and multiplication.
-Prior work addresses this by approximating activations with single polynomials (Diaa et al., 2024) or piecewise polynomials (Liu et al., 2017).
+Activations are non-polynomial, and FHE natively supports only polynomial operations such as addition and multiplication.
+Existing work addresses this by approximating activations with single polynomials [11] or piecewise polynomials [21].
 These approximations can be introduced before or after training.
 In post-training approximation, models are trained with standard activations (e.g., ReLU) and later replaced with high-precision polynomials.
-For example, MPCNN (Lee et al., 2022) uses minimax polynomials of orders $\{15,15,27\}$.
-In pre-training approximation, ReLU is replaced with a polynomial before training, allowing lower-degree approximations that reduce multiplicative depth and speed up inference.
-For example, shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), similar to LeNet-5 (LeCun et al., 1998), achieve over 98% accuracy on MNIST with degree-2 polynomials.
-However, Garimella et al.
-(2021) showed that training larger models (beyond AlexNet (Krizhevsky et al., 2012) or VGG-11 (Simonyan & Zisserman, 2015)) with such low-degree polynomials causes large approximation errors that destabilize training.
-To date, PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial, is the lowest-degree approximation that generalizes to deep neural networks.
-Implementing non-interactive PPML becomes increasingly difficult as model depth increases.
-While bootstrapping enables evaluation of arbitrarily deep models by refreshing ciphertexts, it is an expensive operation in FHE and remains the main scalability bottleneck.
-To avoid this cost, prior work has relied on Leveled Fully Homomorphic Encryption (LFHE) (see §A for details).
+For example, MPCNN [20] uses minimax polynomials of orders $\{15,15,27\}$.
+In pre-training approximation, ReLU is replaced with a polynomial before training, which allows lower-degree approximations that reduce multiplicative depth and speed up inference.
+For example, shallow CNNs like CryptoNets [13] and LoLa [6], similar to LeNet-5 [19], achieve over 98% accuracy on MNIST with degree-2 polynomials.
+However, prior work [12] showed that training larger models (beyond AlexNet [18] or VGG-11 [33]) with such low-degree polynomials causes large approximation errors that destabilize training.
+To date, PILLAR [11], which uses a degree-4 polynomial, is the lowest-degree approximation that generalizes to deep neural networks.
+Implementing non-interactive PPML becomes increasingly difficult with deeper models.
+While bootstrapping enables evaluation of arbitrarily deep models, it is an expensive operation in FHE and remains the main scalability bottleneck.
+To avoid this cost, the literature has relied on Leveled Fully Homomorphic Encryption (LFHE).
 LFHE eliminates bootstrapping and allows faster private inference, but its computation is bounded by a fixed multiplicative depth determined by encryption parameters.
 As a result, the size and complexity of neural networks that can be evaluated under LFHE are strictly limited.
-The deepest CNNs shown to run with LFHE alone are AlexNet and VGG-11, as demonstrated by Garimella et al.
-(2021).
-Larger models such as ResNet-20 or ResNet-32 have so far required bootstrapping.
-For example, AutoFHE executes ResNet-20 inference with at least five bootstrapping layers (Ao & Boddeti, 2024), which leads to heavy computational costs.
-In this work, we focus on ResNet (He et al., 2016), as it is widely adopted in state-of-the-art evaluations and is well-suited for efficient and low-latency deployment.
-Contributions: This work advances non-interactive PPML by enabling, for the first time, large CNNs like ResNet-18, ResNet-20, and ResNet-32 to run entirely under LFHE without bootstrapping, achieving accuracy comparable to plaintext ReLU models, while outperforming prior work in both accuracy and inference time.
-Our contributions are threefold:
-  • To the best of our knowledge, this is the first work to train large CNNs, like ResNet-18, ResNet-20, and ResNet-32 effectively with degree-2 polynomial activations.
-We introduce a novel penalty function that ensures stable training and accuracy on par with ReLU-based models.
-This achieves the theoretical minimum multiplicative depth of one for non-linear activations, whereas the lowest previously known stable alternative, PILLAR, required depth three (see §2).
-  • We introduce three novel structural optimization techniques that reduce the overall multiplicative depth of a model while preserving functional equivalence: (i) Node Fusing, which merges consecutive operations such as convolution and batch normalization by folding normalization parameters into convolution weights, eliminating separate multiplications; (ii) Weight Redistribution, which adjusts parameters so that coefficients of highest-order terms in polynomial activations and batch normalization, as well as divisors in pooling layers, normalize to one, removing redundant multiplications by constants; and (iii) Tower Reuse, which allows multiple multiplications within the same FHE level before rescaling.
-In the CKKS scheme used in this work (Cheon et al., 2017), ciphertexts carry a scale factor that grows with multiplications.
+The deepest CNNs shown to run with LFHE alone are AlexNet and VGG-11 [12].
+Larger models have so far required bootstrapping.
+Consequently, modern deep CNNs have largely been regarded as incompatible with LFHE, and bootstrapping has become the default mechanism to address this limitation despite its high computational cost.
+For example, AutoFHE executes ResNet-32 inference with 8 or 19 bootstrapping layers in its lower- or higher-accuracy configuration, respectively [5], resulting in heavy computational costs.
+This paper revisits the prevailing assumption that modern deep CNNs are incompatible with LFHE.
+We argue that heavy reliance on bootstrapping has obscured a large, underexplored optimization space.
+In FHE-based PPML, multiplicative depth directly determines the encryption parameters and the frequency of bootstrapping operations needed during evaluation, both of which have a dominant impact on inference time.
+Crucially, the multiplicative depth of CNNs is not an immutable property of the model architecture itself, but emerges from how its computation is realized.
+The network structure, degree of polynomial activations, linear transformations, data layout, function implementations, and placement of rescaling operations collectively govern depth.
+Careful analysis and optimization of these components and their interactions can substantially reduce multiplicative depth without changing model functionality.
+By aggressively minimizing multiplicative depth at every level of the model, rather than compensating for inefficiencies with bootstrapping, we show that deep CNNs can be evaluated efficiently under LFHE.
+Guided by this philosophy, our goal is to design the fastest non-interactive PPML framework that does not compromise accuracy, under LFHE.
+We achieve this by combining four techniques: (i) low-degree polynomial activations that attain the theoretical minimum multiplicative depth for non-linear activation functions, (ii) a suite of structural optimizations that remove redundant operations, (iii) an ensemble-based strategy that recovers accuracy to near-ReLU levels, and (iv) a clustering technique to speed up the computation of convolutional layers.
+Importantly, while this work focuses on LFHE, the proposed techniques are not limited to it.
+They can be applied with FHE to run very deep models while substantially reducing the number of required bootstrapping operations.
+Contributions: This work advances non-interactive PPML by enabling, for the first time, deep CNNs (e.g.
+ResNet-32) to run entirely under LFHE without bootstrapping, achieving accuracy comparable to plaintext ReLU models while outperforming prior work in both accuracy and inference time.
+Our contributions are fourfold:
+  • To the best of our knowledge, this is the first work to train deep CNNs effectively with degree-2 polynomial activations.
+We introduce a novel penalty function that ensures stable training and delivers accuracy on par with plaintext ReLU-based models.
+This achieves the theoretical minimum multiplicative depth of one for non-linear activations, whereas the lowest previously known stable alternative, PILLAR, requires a multiplicative depth of three (see §3).
+  • We propose three structural optimization techniques that reduce multiplicative depth of a model while preserving functional equivalence: (i) Node Fusing, which merges consecutive layers by folding normalization parameters into neighboring nodes, effectively eliminating the normalization layer; (ii) Weight Redistribution, which adjusts parameters so that coefficients of highest-order terms in polynomial activations and batch normalization, as well as divisors in pooling layers, normalize to one, removing redundant multiplications by constants; and (iii) Tower Reuse, which allows multiple multiplications within the same FHE level before rescaling.
+In the CKKS scheme used in this work [8], ciphertexts carry a scaling factor that grows with multiplications.
 Rescaling keeps ciphertext values within range but consumes a level.
-Since each level requires a modulus and homomorphic operations apply to all moduli, fewer rescaling steps reduce the number of moduli, leading to faster homomorphic operations.
-Moreover, they permit using larger moduli for higher accuracy (see §3).
-  • We propose an encoding strategy that utilizes unused ciphertext slots in FHE to pack multiple model instances, enabling simultaneous ensemble inference to improve accuracy at some additional inference cost.
-To offset this overhead, we introduce a parameter clustering method for convolutions that recovers much of the lost speed.
-To the best of our knowledge, this is the first work to exploit unallocated ciphertext slots for ensemble inference in FHE (see §4).
-We evaluate our method on CIFAR-10 and CIFAR-100 datasets (Krizhevsky & Hinton, 2009), and show that it achieves up to $4\times$ faster private inference than prior work while achieving accuracy comparable to plaintext ReLU models (see §5).
-2 Model Training with Polynomial Approximation
+Since each level is represented by a modulus and homomorphic operations apply to all moduli, requiring fewer rescaling steps reduce the number of moduli, leading to faster homomorphic operations (see §4).
+  • We introduce a novel encoding strategy that utilizes unused ciphertext slots in FHE to pack multiple CNN model instances, enabling simultaneous ensemble inference within a single homomorphic evaluation.
+This approach enables the use of multiple independently trained models to recover accuracy lost to low-degree polynomial approximations through careful reuse of ciphertext–plaintext structures, albeit with additional inference overhead (see §5).
+  • To offset the additional private inference overhead from the choice of data layout that enables ensemble inference, we introduce a parameter clustering technique for convolutional layers that groups similar parameters to reduce redundant plaintext encodings and homomorphic computation.
+This technique recovers much of the lost inference speed while preserving model accuracy (see §5).
+We evaluate our method on CIFAR [17] and Tiny-ImageNet [10] datasets, and show that it achieves up to $4\times$ faster private inference than prior work while achieving accuracy comparable to plaintext ReLU models (see §6).
+2 Background
+2.1 Leveled Fully Homomorphic Encryption
+Homomorphic Encryption allows meaningful computations to be performed directly on encrypted data.
+In this work, we employ Residue Number System (RNS) variant [7] of the Cheon-Kim-Kim-Song (CKKS) [8] encryption scheme, which supports approximate arithmetic over encrypted complex numbers.
+CKKS is an FHE scheme based on the Ring-Learning With Errors (RLWE) problem [25].
+The scheme operates across three spaces: message $\mathcal{M}=\mathbb{C}$, plaintext $\mathcal{P}=R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$, and ciphertext $\mathcal{C}=R_{Q}\times R_{Q}$.
+Encoding maps a vector of $N/2$ numbers $\in\mathcal{M}$ into a degree-$N$ polynomial $\in\mathcal{P}$, and encryption transforms a plaintext $\in\mathcal{P}$ into a ciphertext $\in\mathcal{C}$.
+In RNS-CKKS, the ciphertext modulus $Q$ is decomposed into $\mathcal{L}+1$ smaller primes $Q=\prod_{i=0}^{\mathcal{L}}{q_{i}}$, where $\mathcal{L}$ denotes the level.
+A polynomial in $R_{Q}$ is represented as $\mathcal{L}+1$ polynomials in $R_{q_{i}}\forall i\in[0,\mathcal{L}]\cap\mathbb{Z}$, which allows computations using native 64-bit integer arithmetic.
+CKKS supports homomorphic additions, multiplications, and rotations.
+Additions and multiplications act element-wise on the encoded vector, akin to SIMD operations.
+Rotations rotate the vector of $N/2$ encoded messages by a step $\in\mathbb{Z}_{N/2}$.
+During encoding, coefficients are scaled by a factor $\Delta$, rounded to the nearest integers, and reduced modulo $Q$.
+Multiplying two encoded messages scales the result to $\Delta^{2}$, requiring rescaling to restore the scale to $\Delta$.
+Rescaling removes one RNS modulus, reducing the level by 1.
+When the level reaches zero, bootstrapping can reset the ciphertext to a higher level, enabling unbounded computation.
+However, it is computationally expensive and avoided when possible.
+FHE without bootstrapping is called Levelled Fully Homomorphic Encryption (LFHE).
+In LFHE, the encryption parameters $(N,Q)$ are chosen to provide the desired security and sufficient levels for a circuit with known multiplicative depth.
+2.2 Challenges of Using Polynomial Activation
+ReLU, defined as $ReLU(x)=\max(0,x)$, is one of the most widely used activation functions in deep learning due to its simplicity and computational efficiency.
+Its piecewise linear form provides the non-linearity required for neural networks to model complex patterns.
+However, in FHE settings, ReLU is difficult to implement because it relies on conditional branching, whereas FHE natively supports only additions and multiplications.
+As a result, any FHE-compatible ReLU must be a polynomial approximation.
+Higher-degree polynomials produce more accurate approximations but incur larger multiplicative depth.
+For a degree-$d$ polynomial, the multiplicative depth is $\delta=\left\lceil\log_{2}d\right\rceil+1$, where the $(+1)$ accounts for coefficient multiplication.
+The lowest multiplicative depth for a ReLU approximation that generalizes to deep CNNs is achieved by PILLAR [11], which uses a degree-4 polynomial $\sum_{i=0}^{4}c_{i}x^{i}$ with $\delta=3$.
+The theoretical minimum multiplicative depth for an activation function is one, attainable with a quadratic polynomial $x^{2}+c_{1}x+c_{0}$.
+Such polynomials have been successfully used in shallow CNNs like CryptoNets [13] and LoLa [6], but they fail in deeper models due to the escaping activation problem [12].
+Moreover, polynomial approximations are accurate only over bounded input intervals.
+When used as direct substitutes for ReLU during training, they often cause a substantial drop in model accuracy [12, 15, 11].
+This section focuses on two main reasons for this degradation: escaping activations, where intermediate layer outputs leave the polynomial’s approximation range, and coefficient truncation, arising from representing polynomial coefficients in fixed-point format to comply with FHE arithmetic.
+Escaping Activations.
+Let $p_{d}(x)=\sum_{k=0}^{d}a_{k}x^{k}$ be a degree-$d$ polynomial approximation of ReLU, with coefficients $a_{k}\in\mathbb{R}$, $\forall k\in\{0,\dots,d\}$.
+These coefficients are typically obtained by minimizing the least-squares error between $p_{d}(x)$ and ReLU on a finite set of real-valued points $\{x_{i}\}_{i=1}^{N}\subset[-c,c]$, for some parameter $c>0$: $\min_{a_{0},\ldots,a_{d}}\;\sum_{i=1}^{N}\left(ReLU(x_{i})\;-\;p_{d}(x_{i})\right)^{2}.$ The resulting polynomial approximates ReLU well on [c,c][-c,c] but can deviate substantially outside this range.
+ReLU exhibits piecewise linear growth:
+$ReLU(x)\;=\;\begin{cases}0,&\text{if }x\leq 0\\ x,&\text{if }x>0\end{cases}\quad\Rightarrow\quad ReLU(x)\;=\;\mathcal{O}(x)$
+whereas any polynomial $p_{d}(x)$ with $d\geq 2$ grows as:
+$p_{d}(x)\;=\;a_{d}x^{d}\;+\;a_{d-1}x^{d-1}\;+\;\cdots\;+\;a_{1}x\;+\;a_{0}\;=\;\mathcal{O}(x^{d})$
+This mismatch in growth rates leads to the escaping activations phenomenon [12]: as inputs propagate through the network, intermediate values can move outside [c,c][-c,c], into regions where $p_{d}(x)$ no longer resembles ReLU.
+The resulting excessively large activations can cause exploding weights and rapid degradation of model performance, unless the training procedure is specifically adapted to control them.
+Coefficient Truncation.
+In privacy-preserving FHE inference, all computations use fixed-point arithmetic with limited precision.
+This restriction limits both the range and resolution of representable values, constraining the domain of activation function inputs as well as the magnitudes of their polynomial approximation coefficients.
+The coefficients $\{a_{k}\}_{k=0}^{d}$ are usually obtained via floating-point least-squares fitting and then quantized for fixed-point evaluation under FHE: $\tilde{a}_{k}\;=\;\frac{\lfloor a_{k}\cdot 2^{b}\rceil}{2^{b}},$ where $b\in\mathbb{N}$ is the number of fractional bits in the fixed-point format and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer.
+Least-squares fitting often produces small-magnitude coefficients, especially for higher-degree monomials.
+If $\left|a_{k}\right|<2^{-(b+1)}$, the quantized coefficient $\tilde{a}_{k}$ becomes zero, effectively removing the monomial $x^{k}$ from the approximation.
+This coefficient truncation [11] changes the polynomial’s shape and can cause substantial deviation from the intended ReLU behavior, even within the approximation interval [c,c][-c,c].
+3 Training with Polynomial Approximation
 3.1 Problem Setting
-Dataset.
-We consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$ for multi-class classification, where $x_{i}\in\mathbb{R}^{n}$ is the feature vector and $y_{i}\in\{1,\dots,K\}$ is the class label.
-$\mathcal{D}$ has $M$ samples and $K$ classes.
+Dataset Configuration.
+Consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$ for a multi-class classification task, where $x_{i}\in\mathbb{R}^{n}$ denotes the feature vector and $y_{i}\in\{1,\dots,K\}$ the corresponding label.
+The dataset contains $M$ samples distributed across $K$ classes.
 ReLU Network.
-Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ be an $L$-layer feed-forward network with ReLU activations, parameterized by weight matrices $W^{(1)},\dots,W^{(L)}$, where each $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $n_{l}$ is the number of neurons in layer $l$.
-The network output is given by $f(x)=h^{(L)}(x)$, where $h^{(0)}(x)=x$ and $h^{(l)}(x)=ReLU(W^{(l)}h^{(l-1)}(x))$ for all $1\leq l\leq L$.
-The activation function $ReLU(z)$ is applied coordinate-wise and defined as $ReLU(z)=\max(0,z)$.
+Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ denote an $L$-layer feed-forward neural network with ReLU activations.
+The network is parameterized by weight matrices $\{W^{(l)}\}_{l=1}^{L}$, where each $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $n_{l}$ denotes the number of neurons in layer $l$.
+The network output is defined as $f(x)=h^{(L)}(x)$, where $h^{(0)}(x)=x$ and $h^{(l)}(x)=ReLU(W^{(l)}h^{(l-1)}(x))$ for all $1\leq l\leq L$.
+The activation function $ReLU(z)$ is applied coordinate-wise as $ReLU(z)=\max(0,z)$.
 Polynomial Network.
-We define $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with the same architecture as $f$, but replace ReLU with a degree-$d$ polynomial $p_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$ that approximates ReLU over a bounded interval [c,c][-c,c].
-The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are chosen so that the maximum approximation error satisfies: $\max_{z\in[-c,c]}|ReLU(z)-p_{d}(z)|\leq\varepsilon$, where $\varepsilon>0$ is a small constant.
-The network output is given by $g(x)=h^{(L)}_{p_{d}}(x)$, where $h^{(0)}_{p_{d}}(x)=x$ and $h^{(l)}_{p_{d}}(x)=p_{d}(W^{(l)}h^{(l-1)}_{p_{d}}(x))$ for all $1\leq l\leq L$.
+We define a network $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with the same architecture as $f$, but replacing each ReLU activation with a degree-$d$ polynomial $p_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$ that approximates the ReLU function on a bounded interval [c,c][-c,c].
+The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are chosen so that the maximum approximation error satisfies: $\max_{z\in[-c,c]}|ReLU(z)-p_{d}(z)|\leq\varepsilon$, where $\varepsilon>0$ is a prescribed approximation tolerance.
+The resulting network output is given by $g(x)=h^{(L)}_{p_{d}}(x)$, where $h^{(0)}_{p_{d}}(x)=x$ and $h^{(l)}_{p_{d}}(x)=p_{d}(W^{(l)}h^{(l-1)}_{p_{d}}(x))$ for all $1\leq l\leq L$.
 Challenges.
-Using polynomial networks ensures FHE compatibility but introduces training challenges such as escaping activations and coefficient truncation (see §B for more details).
-Escaping activations occur when intermediate outputs drift outside the interval [c,c][-c,c], leading to large approximation errors.
-Coefficient truncation occurs when polynomial coefficients are stored with limited fixed-point precision $b$, which changes the polynomial’s shape and causes deviation from ReLU even inside the approximation interval.
-These issues require careful selection of $d$, [c,c][-c,c], and $b$, together with regularization during training to keep activations within the valid approximation range.
+While polynomial activations enable compatibility with FHE schemes, they introduce several non-trivial optimization challenges, most notably escaping activations and coefficient truncation, as discussed in §2.2.
+Addressing these issues typically requires careful selection of the polynomial degree $d$, the approximation interval [c,c][-c,c], and coefficient bounds $b$, together with regularization strategies designed to ensure that intermediate activations remain within the valid approximation range throughout training.
 3.2 Training Strategy
 Quantization-Aware Polynomial Fitting.
-To avoid coefficient truncation, we integrate fixed-point constraints directly into coefficient estimation using quantization-aware polynomial fitting (Diaa et al., 2024).
+To avoid precision loss due to coefficient truncation, we incorporate fixed-point constraints directly into coefficient estimation procedure using quantization-aware polynomial fitting [11].
 Given an approximation interval [c,c][-c,c] and fixed-point precision with $b$ fractional bits, we define a quantized input domain for polynomial fitting as: $X=\{x\in[-c,c]\mid x=k\cdot 2^{-b},\ k\in\mathbb{Z}\}.$ For each input $x_{i}\in X$, we compute scaled ReLU outputs $Y_{i}=2^{b}\cdot ReLU(x_{i})$ so that regression targets are integers, which reduces risk of precision loss during optimization.
-We build a matrix $B\in\mathbb{R}^{M\times(d+1)}$ with entries $B_{i,k}=x_{i}^{k}\ \text{for }0\leq k\leq d.$ The coefficients $A=[a_{0},\ldots,a_{d}]^{\top}\in\mathbb{Z}^{d+1}$ are obtained by solving bounded integer least-squares problem: $\min_{A\in\mathbb{Z}^{d+1}}\|BA-Y\|_{2}^{2}\ \text{subject to }a_{k}\in[-2^{b-1},2^{b-1}-1].$ The resulting fixed-point polynomial is $p_{d}(x)=\sum_{k=0}^{d}(\frac{a_{k}}{2^{b}})x^{k}.$
+We construct a matrix $B\in\mathbb{R}^{M\times(d+1)}$ with entries $B_{i,k}=x_{i}^{k}\ \text{for }0\leq k\leq d.$ The coefficients $A=[a_{0},\ldots,a_{d}]^{\top}\in\mathbb{Z}^{d+1}$ are obtained by solving bounded integer least-squares problem: $\min_{A\in\mathbb{Z}^{d+1}}\|BA-Y\|_{2}^{2}\ \text{subject to }a_{k}\in[-2^{b-1},2^{b-1}-1].$ The resulting fixed-point polynomial is $p_{d}(x)=\sum_{k=0}^{d}(a_{k}/2^{b})x^{k}.$
 Activation Regularization.
@@
 For a mini-batch $\mathcal{B}\subset\mathcal{D}$, the following training loss is minimized:
-$\mathcal{L}_{\mathcal{B}}\;=\;\;\underbrace{\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell_{\mathrm{CE}}\left(g\left(x\right),y\right)}_{\text{classification loss}}\;+\;\zeta\;\underbrace{\frac{1}{L}\sum_{l=1}^{L}\left\|z_{p}^{(l)}\;-\;\mathrm{clip}\left(z_{p}^{(l)};[-c,c]\right)\right\|_{2}}_{\text{clip-range penalty}}$ (1)
+$\begin{split}\mathcal{L}_{\mathcal{B}}&=\underbrace{\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\ell_{\mathrm{CE}}\!\left(g(x),y\right)}_{\text{classification loss}}\\ &\quad+\;\zeta\;\underbrace{\frac{1}{L}\sum_{l=1}^{L}\left\|z_{p}^{(l)}-\mathrm{clip}\!\left(z_{p}^{(l)};[-c,c]\right)\right\|_{2}}_{\text{clip-range penalty}}\end{split}$ (1)
 where $\ell_{\mathrm{CE}}$ is the cross-entropy loss, $z_{p}^{(l)}$ are pre-activations at layer $l$, i.e., outputs of affine transformations $W^{(l)}h^{(l-1)}_{p_{d}}(x)$ before the polynomial activation.
 The function $\mathrm{clip}(z;[-c,c])$ is applied element-wise and defined as: $\mathrm{clip}(z;[-c,c])_{i}=\max(-c,\min(z_{i},c))$.
-The second loss term penalizes pre-activations that lie outside the interval [c,c][-c,c], with strength controlled by the regularization parameter $\zeta>0$.
-See §C for a proof of the penalty function’s correctness.
+The second loss term penalizes pre-activations that lie outside [c,c][-c,c], with strength controlled by the regularization parameter $\zeta>0$.
+See §A for a proof of the penalty function’s correctness.
 Training can still become unstable for two reasons: (i) in early epochs pre-activations can grow unbounded before the model learns to contain them and (ii) using full regularization weight $\zeta$ from the start can let the penalty dominate and destabilize optimization.
-As additional strategies, we also consider clipping pre-activations during training and introducing a warm-up schedule for $\zeta$ (see §D for more details).
+As additional strategies, we also consider clipping pre-activations during training and introducing a warm-up schedule for $\zeta$.
+Pre-Activation Clipping.
+During early training stages, network weights remain close to their random initialization, and the optimization process does not immediately constrain the pre-activation values to lie within the target approximation interval [c,c][-c,c].
+As a result, unbounded pre-activations may arise before the model learns to keep them within range, causing training instability that can degrade model behavior.
+To prevent this, we restrict pre-activations to the interval [c,c][-c,c] before evaluating the polynomial activation, which is achieved through a clipping strategy $\mathrm{clip}(z;[-c,c])$ [11]:
+$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(\mathrm{clip}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x);[-c,c]\right)\right).$
+Importantly, this clipping operation is applied only during training to stabilize learning.
+It is performed after computing the clip-range penalty to ensure that gradients from the regularization term, which encourages the network to keep pre-activations within the approximation interval [c,c][-c,c], are preserved and not masked by the clipping.
+At inference time, the clipping function is removed: $h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x)\right)$.
+The model, having learned to constrain pre-activations during training, is expected to remain within the approximation interval without explicit clipping at inference.
+Regularization Warm-up.
+While activation regularization and pre-activation clipping are both necessary for training stability, applying the full regularization strength from the outset can lead to numerical instability, particularly in larger models.
+In the early epochs, many pre-activation values lie outside the target interval [c,c][-c,c] across several layers.
+The clip-range penalty adds a contribution from every such layer, so the penalty term becomes very large.
+In extreme cases, this can cause the total loss to become numerically undefined.
+To address this issue, the regularization strength $\zeta$ is progressively increased over the initial epochs, which is implemented through a regularization warm up schedule [11].
+Let $T_{warm}$ denote the total number of warm-up epochs.
+For each epoch $t$, we define a time-dependent regularization weight $\zeta_{t}$:
+$\zeta_{t}\;=\;\begin{cases}\alpha_{t}\cdot\zeta&\text{if }t\leq T_{warm}\\ \zeta&\text{if }t>T_{warm}\end{cases},$
+where $\{\alpha_{t}\}_{t=1}^{T_{warm}}$ is a fixed sequence of scaling factors satisfying $0<\alpha_{1}<\cdots<\alpha_{T_{warm}}<1$.
+In practice, we empirically find that setting $T_{warm}=4$ provides a stable convergence.
+We use a predefined sequence of scaling factors: $\alpha=\left\{\frac{1}{100},\frac{1}{50},\frac{1}{10},\frac{1}{5}\right\}$, which produces the epoch-wise regularization strength: $\zeta_{t}\;\in\;\left\{\frac{\zeta}{100},\frac{\zeta}{50},\frac{\zeta}{10},\frac{\zeta}{5},\zeta,\zeta,\dots\right\}$.
+This warm-up schedule ensures that the regularization penalty is introduced progressively in the initial training epochs without suffering from exploding loss values.
+Figure 1 shows an overview of the training process with polynomial approximation.
+Figure 1: Overview of the two-stage training pipeline of a neural network with polynomial approximation.
+1 A ReLU network is approximated using quantization-aware polynomial fitting to obtain an initial polynomial network.
+2 The polynomial network is then trained through regularized polynomial network training, consisting of activation regularization, pre-activation clipping, and warm-up scheduling, to ensure stable convergence.
 4 Structural Optimizations
@@
 Our first optimization is node fusing, which merges consecutive computational nodes into a single equivalent function.
-This reduces both number of operations and multiplicative depth.
+This reduces both the number of operations and multiplicative depth.
 Formally, two sequential nodes $F(G(x))$ are replaced with $H(x)$ such that $F(G(x))=H(x)$, with $H(x)$ more efficient to evaluate.
-Since our method uses batch normalization, recall its polynomial form $B(x)=b_{1}x+b_{0},\;b_{1}=\frac{\gamma}{\sigma},\;b_{0}=\beta_{b}-b_{1}\mu$, where $\mu,\sigma$ are the input mean and standard deviation, and $\gamma,\beta_{b}$ are learnable parameters.
-We outline four common fusion cases.
-See §E.1 for full derivations.
+Following this principle, we present four cases where compositions are fused into a single, more efficient operation, as illustrated in Figure 2.
+Since all cases involve batch normalization, recall its polynomial form:
+$B(x)=b_{1}x+b_{0}$
+with $b_{1}=\frac{\gamma}{\sigma},\;b_{0}=\beta_{b}-b_{1}\mu$, where $\mu$, $\sigma$ are the input mean and standard deviation, and $\gamma$, $\beta_{b}$ are learnable parameters.
+We outline the fusion cases below.
+See §B.1 for full derivations.
 Case 1: $P(B(x))\mapsto P(x)$.
@@
 Case 2: $B(C(x))\mapsto C(x)$.
-When batch normalization follows a convolution $C(\cdot)$, the two can be fused (Markuš, 2018).
-A convolution is given by $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$, where $w_{i}$ are the convolution weights and $\beta_{c}$ is the bias.
+Batch normalization can be fused into a preceding convolution [26].
+A convolution is given by $C(x)=\sum_{i}w_{i}x_{i}+\beta_{c}$, where $w_{i}$ are the weights and $\beta_{c}$ is the bias.
 The fused form becomes a single convolution $B(C(x))=\sum_{i}\omega_{i}x_{i}+\alpha$, with fused coefficients $\omega_{i}=b_{1}w_{i},\;\alpha=b_{1}\beta_{c}+b_{0}$.
@@
 Specifically, $P(B_{X}(x)+B_{Y}(y))$.
-Here, $B_{X}$ and $B_{Y}$ are independent batch normalization layers to input $x$ and $y$.
-This structure can be fused into a single quadratic bivariate polynomial, which we call polyskip: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2}b_{Y1}^{2},\;d_{XY}=2c_{2}b_{X1}b_{Y1}$, $d_{X}=b_{X1}(2c_{2}(b_{X0}+b_{Y0})+c_{1}),\;d_{Y}=b_{Y1}(2c_{2}(b_{X0}+b_{Y0})+c_{1})$, $d_{0}=c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}$.
+Here, $B_{X}$ and $B_{Y}$ are independent batch normalization layers to inputs $x$ and $y$, respectively.
+This structure can be fused into a single quadratic bivariate polynomial $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2}b_{Y1}^{2},\;d_{XY}=2c_{2}b_{X1}b_{Y1}$, $d_{X}=b_{X1}(2c_{2}(b_{X0}+b_{Y0})+c_{1}),\;d_{Y}=b_{Y1}(2c_{2}(b_{X0}+b_{Y0})+c_{1})$, $d_{0}=c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}$.
 Case 4: $P(B_{X}(x)+y)\mapsto S(x,y)$.
 Some skip connections use identity shortcuts, where the input $y$ is added directly to the batch-normalized branch $B_{X}(x)$ before applying the activation.
-This structure can be fused into a quadratic bivariate polynomial: $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$,with coefficients $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2},\;d_{XY}=2c_{2}b_{X1}$, $d_{X}=b_{X1}(2c_{2}b_{X0}+c_{1}),\;d_{Y}=2c_{2}b_{X0}+c_{1}$, $d_{0}=c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}$.
-Node fusing collapses batch normalization, activations, convolutions, and skip connections into single fused polynomials, eliminating redundant nodes and reducing multiplicative depth.
+This structure can be fused into a quadratic bivariate polynomial $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$, with $d_{X2}=c_{2}b_{X1}^{2},\;d_{Y2}=c_{2},\;d_{XY}=2c_{2}b_{X1}$, $d_{X}=b_{X1}(2c_{2}b_{X0}+c_{1}),\;d_{Y}=2c_{2}b_{X0}+c_{1}$, $d_{0}=c_{2}b_{X0}^{2}+c_{1}b_{X0}+c_{0}$.
+Node fusing collapses batch normalizations, activation functions, convolutional layers, and skip connections into single fused polynomials, eliminating redundant nodes and reducing multiplicative depth.
+This technique is applied after model training but before inference, transforming the model structure into a functionally equivalent, more efficient form.
+Figure 2: Illustration of node fusing cases. Nodes are batch normalization ( B ), convolution ( C ), univariate ( P ) and bivariate ( S ) polynomial activations, and addition ( + ).
 4.2 Weight Redistribution
 Our second optimization is weight redistribution, a technique that redistributes weights across the network while maintaining functional equivalence.
-The goal is to reduce the multiplicative depth of certain functions by one.
-It specifically targets average pooling, batch normalization, and polynomial activations.
+The goal is to reduce the multiplicative depth of selected operations by one.
+It specifically targets average pooling, batch normalization, and polynomial activation.
 For average pooling, redistribution eliminates the normalization step by setting the divisor to one.
-For polynomial functions, the highest-order coefficient is set to one.
-For example, the polynomial activation $c_{2}x^{2}+c_{1}x+c_{0}$ is transformed into $x^{2}+p_{1}x+p_{0}$, reducing its multiplicative depth from two to one, the theoretical minimum for an activation function.
+For polynomial functions, coefficients are normalized so that the highest-order coefficient is set to one.
+For example, the polynomial activation $c_{2}x^{2}+c_{1}x+c_{0}$ is transformed into $x^{2}+\bar{c}_{1}x+\bar{c}_{0}$, reducing its multiplicative depth from two to one, the theoretical minimum for an activation function.
 These transformations alone break model equivalence.
-To maintain it, other nodes in the network must compensate for the change.
-We refer to the nodes initiating weight redistribution as donors, and the nodes adjusting their parameters to preserve equivalence as receivers.
-We represent the network as a directed graph and traverse it, identifying all eligible donor nodes.
-For each donor, redistribution can be applied either forward or backward, affecting receivers among the donor’s successors or predecessors, respectively.
+To maintain it, other nodes in the network must compensate for the redistributed weights.
+We refer to the nodes that initiate redistribution as donors, and to nodes that adjust their parameters to restore equivalence as receivers.
+We represent the network as a directed graph and traverse it to identify all eligible donor nodes.
+For each donor, redistribution can be applied either forward or backward, affecting receivers among the donor’s successors or predecessors, respectively, as illustrated in Figure 3.
 Update Forward.
@@
 Our goal is to construct a normalized function $\bar{F}(x)$ such that $\upsilon\bar{F}(x)=F(x)$, where $F(x)$ is the original donor function and $\upsilon$ is the update term to be propagated to receivers.
+Average Pooling.
 The average pooling operation is given by $\mu(x)=k^{-1}\sum_{i}{x_{i}}$, where $k$ is the kernel size.
 In the normalized function we set $k=1$, reducing the operation to $\bar{\mu}(x)=\sum_{i}x_{i}$.
-Thus, for the equality to be valid, we have $\upsilon=k^{-1}$.
+Thus, for the equality to be true, we have $\upsilon=k^{-1}$.
+Polynomial Functions.
 Batch normalization and polynomial activation are polynomial functions.
 For a univariate polynomial $P(x)=\sum_{i=0}^{d}{c_{i}x^{i}}$ of degree $d$, and a normalized polynomial $\bar{P}(x)=x^{d}+\sum_{i=0}^{d-1}{\bar{c}_{i}x^{i}}$, where $\bar{c}_{i}$ are the normalized coefficients, we have $\upsilon=c_{d}$ and $\bar{c}_{i}=c_{i}\upsilon^{-1}$ for $\upsilon\bar{P}(x)=P(x)$ to hold.
-The extension to the bivariate polynomial activation is analogous and deferred to §E.2.
+The extension to the bivariate polynomial activation is analogous and deferred to §B.2.
 Receivers.
@@
 Updating the donor to $\bar{F}(\cdot)$ without compensating the receiver results in $G(\upsilon^{-1}F(\cdot))$, breaking model equivalence.
-To restore it, we modify the receiver so that $\bar{G}(\bar{F}(\cdot))=\bar{G}(\upsilon^{-1}F(\cdot))=G(F(\cdot))$.
+To restore it, we modify the receiver to $\bar{G}(\cdot)$ so that $\bar{G}(\bar{F}(\cdot))=\bar{G}(\upsilon^{-1}F(\cdot))=G(F(\cdot))$.
 Receivers fall into two categories: kernel functions and polynomial functions.
-Kernel functions share the form $K(x)=\sum_{i}w_{i}x_{i}+\beta$.
-To maintain model equivalence, we must update the kernel function such that $\bar{K}(\upsilon^{-1}x)=K(x)$.
-For that, we must set $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta$.
-Polynomial functions update as $\bar{P}(\upsilon^{-1}x)=P(x)$, which implies that $\bar{c}_{i}=c_{i}\upsilon^{i}$.
-For the bivariate case, in $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$ the index $i$ corresponds to the exponent of $x$, while in $\bar{S}(x,\upsilon^{-1}y)=S(x,y)$ it corresponds to the exponent of $y$.
+Kernel Functions.
+Convolution, linear, and average pooling share the form $K(x)=\sum_{i}w_{i}x_{i}+\beta$ ($w_{i}=k^{-1}$ and $\beta=0$ for average pooling).
+To maintain model equivalence, we must update the kernel function such that $\bar{K}(\upsilon^{-1}x)=K(x)$ is valid.
+For that, we set $\bar{w}_{i}=w_{i}\upsilon$ and keep $\bar{\beta}=\beta$.
+Polynomial Functions.
+Batch normalization and polynomial activation update as $\bar{P}(\upsilon^{-1}x)=P(x)$, implying that $\bar{c}_{i}=c_{i}\upsilon^{i}$.
+For the bivariate polynomial activation, in $\bar{S}(\upsilon^{-1}x,y)=S(x,y)$ the index $i$ corresponds to the exponent of $x$, while in $\bar{S}(x,\upsilon^{-1}y)=S(x,y)$ it corresponds to the exponent of $y$.
 Update Backward.
@@
 Receivers.
-Receivers must satisfy $\bar{G}(x)=\upsilon G(x)$ to preserve model equivalence.
-Kernel receivers are updated as $\bar{K}(x)=\upsilon K(x)$, which implies that $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta\upsilon$.
-And polynomial receivers update according to $\bar{P}(x)=\upsilon P(x)$; therefore, $\bar{c}_{i}=c_{i}\upsilon$.
+To preserve model equivalence, receivers must satisfy $\bar{G}(x)=\upsilon G(x)$.
+Kernel receivers update as $\bar{K}(x)=\upsilon K(x)$, which implies that $\bar{w}_{i}=w_{i}\upsilon$ and $\bar{\beta}=\beta\upsilon$.
+And polynomial receivers update according to $\bar{P}(x)=\upsilon P(x)$; hence, $\bar{c}_{i}=c_{i}\upsilon$.
+Figure 3: Illustration of weight redistribution cases.
+$D$ represents a donor and $R$ a receiver.
+The number of bars ($\bar{\cdot}$) on a $D$ or $R$ represents the number of updates the node received.
+$\mathbf{f}$ represents a forward and $\mathbf{b}$ a backward update.
+Weight redistribution is performed after model training, specifically immediately following node fusing, and prior to inference.
+It reduces multiplicative depth, while maintaining model equivalence.
+Figure 3 illustrates common cases for CNNs, including residual connections, where redistribution must be applied consistently along all converging or diverging branches.
+Refer to §B.2 for detailed derivations for all cases.
 4.3 Tower Reuse
-In CKKS, the Residue Number System (RNS) moduli are typically chosen as primes close to the scaling factor $\Delta$ (see §A for an overview of LFHE).
-This choice is motivated by the fact that, after each homomorphic multiplication, the scaling factors of the operands are multiplied, producing a ciphertext with scale $\Delta^{2}$.
-To control this growth, the rescale operation reduces the scale back to $\Delta$ by removing one modulus from the modulus chain.
-However, rescaling introduces approximation errors: the further the dropped modulus deviates from $\Delta$, the larger the error.
-These errors may accumulate and amplify through subsequent homomorphic operations.
-A larger $\Delta$ is desirable for two reasons: (i) it provides higher precision for CKKS encoding and (ii) the RNS moduli are relatively closer to the scaling factor, reducing approximation errors.
-The drawback is that a larger $\Delta$ requires larger RNS moduli, which in turn increases the ciphertext modulus $Q=\prod_{i=0}^{\mathcal{L}}q_{i}$, where $\mathcal{L}$ is the number of levels and $q_{i}$ are the individual RNS primes.
-Since the security level depends on the ratio $N/Q$ (with $N$ the polynomial degree), a larger $Q$ reduces security.
-One way to increase security is to increase $N$, but this slows computations due to larger polynomials.
+In CKKS, the RNS moduli are typically selected as primes close to the scaling factor $\Delta$.
+This choice is motivated by the fact that the scaling factors of operands multiply after each homomorphic multiplication, yielding a ciphertext with scale $\Delta^{2}$.
+To prevent exponential growth of the scale, the rescale operation is used to reduce the scale back to $\Delta$, consuming one modulus from the modulus chain in the process.
+For this reason, the number of moduli depends on the application’s multiplicative depth $\delta$.
+Since homomorphic operations are carried out across all moduli, their number directly impacts computational efficiency.
+Rescaling also introduces approximation errors: the further the dropped modulus from $\Delta$, the larger the error.
+This issue is especially relevant for small scaling factors, as fewer RNS primes lie relatively close to $\Delta$.
+These errors can accumulate and amplify through subsequent homomorphic operations, increasing the likelihood of incorrect results.
+Using a large $\Delta$ has drawbacks, as it requires larger RNS moduli and therefore increases the ciphertext modulus $Q=\prod_{i=0}^{\mathcal{L}}q_{i}$, where $\mathcal{L}$ is the number of levels and $q_{i}$ are the RNS primes, which reduces the security level.
+One can increase security by increasing the polynomial degree $N$.
+However, it is undesirable because operating on larger polynomials slows computation.
 Thus, for a fixed security level and polynomial degree, there is a maximum allowable $Q$.
-Consequently, for an application-defined multiplicative depth $\delta$, this imposes a bound on the maximum scaling factor $\Delta$.
+Consequently, for an application-defined $\delta$, this imposes a bound on the maximum $\Delta$.
 In practice, when $\delta$ is large, as is common in deep learning models, the resulting $\Delta$ is small.
-This has two adverse effects, opposite to the benefits of a large $\Delta$: lower numerical precision and larger approximation errors, as moduli $q_{i}$ are relatively farther from $\Delta$.
 Proposed Solution.
-We introduce a more general method for determining the RNS moduli by introducing sublevels within each level.
+We introduce tower reuse, a more general method for determining the RNS moduli by introducing sublevels within each level.
 Instead of enforcing $q_{i}\approx\Delta$, we allow $q_{i}\approx\Delta^{\ell}$, where $\ell$ denotes the number of sublevels.
 Rescaling is performed only when a ciphertext scale exceeds the sublevel capacity of the modulus to be dropped.
-Formally, $\lambda(x)>\lambda(q_{i})\implies\downarrow x$, where $\downarrow$ denotes the rescale operation, and $\lambda(x)=\left\lfloor\log_{\Delta}{\Delta_{x}}\right\rceil$ returns the sublevel of a ciphertext, plaintext, or modulus: with $\Delta_{x}$ referring to the scale of $x$ and $\Delta$ being the default scale defined by the encryption parameters.
-See §E.3 for an example of this technique.
-This method allows smaller $\Delta$ values while maintaining larger $q_{i}$, mitigating rescaling errors.
-Moreover, fewer effective levels (moduli) reduce error amplification, and since homomorphic operations are applied across all moduli, reducing their number improves computational efficiency.
+Formally, $\lambda(x)>\lambda(q_{i})\implies\downarrow x$, where $x$ represents a ciphertext, $\downarrow$ denotes the rescale operation, and $\lambda(x)=\left\lfloor\log_{\Delta}{\Delta_{x}}\right\rceil$ returns the sublevel of a ciphertext, plaintext, or modulus: with $\Delta_{x}$ referring to the scale of $x$ and $\Delta$ being the default scaling factor defined by the encryption parameters.
+Example.
+Suppose following an activation function, there is a convolution.
+Let modulus $q_{i}\approx\Delta^{2}$, thus $\lambda(q_{i})=2$.
+Consider an input $x$ with scaling factor $\Delta_{x}=\Delta$, giving $\lambda(x)=1$.
+Applying a quadratic activation $y=x^{2}+c_{1}x+c_{0}$ yields $\lambda(y)=2$, with $\lambda(c_{1})=1$ and $\lambda(c_{0})=2$.
+Next, a convolution $z=\sum_{i}\omega_{i}y_{i}+\alpha$ is performed.
+Since $\lambda(y)=2$, by setting $\lambda(\omega)=1$, we obtain $\lambda(z)=3$.
+As $\lambda(z)>\lambda(q_{i})$, a rescale is triggered.
+The rescale operation drops $q_{i}$ from the modulus chain, reducing the level $\Lambda(\cdot)$ by one and resetting the sublevel to one:
+$\Lambda(\downarrow z)=\Lambda(z)-1,\qquad\lambda(\downarrow z)=\lambda(z)-\lambda(q_{i})=1.$
+In terms of $\Delta$, a ciphertext $z$ with $\Delta_{z}=\Delta^{3}$ rescaled by $q_{i}\approx\Delta^{2}$ results in output scale $\Delta_{\downarrow z}\approx\Delta$.
+Tower reuse is an automatic inference-time technique governed by $q_{i}$, $\Delta$, and the network structure.
+It improves computational efficiency and reduces approximation errors.
+The efficiency gains arise from decreasing the number of RNS moduli by a factor of $\ell$, since the method allows $\ell$ levels of multiplication to be performed before rescaling, and from the corresponding reduction in the number of required rescale operations.
+The reduction in approximation error is likewise due to the fewer rescale operations, and also to the fact that the method permits the use of smaller $\Delta$ values together with larger $q_{i}$, making the moduli relatively closer to $\Delta^{\ell}$.
 4.4 Impact on RNS Levels
-The combination of our polynomial activation function with the structural optimizations proposed in this section reduces the number of RNS levels $\mathcal{L}$ required for multiplication compared to PILLAR (Diaa et al., 2024), which previously achieved the lowest $\mathcal{L}$: from 87 to 18 for ResNet-18, from 97 to 20 for ResNet-20, and from 157 to 32 for ResNet-32.
-A complete analysis of the contribution of each technique to the reduction in required levels for ResNet models is provided in §E.4.
-4 Codesign Techniques
+The combination of our polynomial activation function with the structural optimizations proposed in this section reduces the number of RNS levels $\mathcal{L}$ required for multiplication compared to PILLAR [11], which previously achieved the lowest $\mathcal{L}$: from 79 to 16 for VGG-16, 87 to 18 for ResNet-18, 97 to 20 for ResNet-20, and 157 to 32 for ResNet-32.
+Level Analysis.
+Table 1 summarizes the number of levels $\mathcal{L}$ required for multiplication across VGG and ResNet models under various optimization techniques.
+To illustrate, we focus on ResNet-18, which comprises 17 convolutional layers, batch normalization layers, and activation functions, along with a single average pooling and linear layer:
+  • $P_{4}$.
+Prior work with the lowest $\mathcal{L}$ [11] uses a degree-4 polynomial activation, giving $\mathcal{L}=87$: convolutions, batch normalizations, average pooling, and linear layers each contribute $\mathcal{L}=1$, and the activation contributes $\mathcal{L}=3$.
+  • $P_{2}$. By replacing the activation function with a degree-2 polynomial, as described in §3, the required level for activations drops to 2, reducing the model $\mathcal{L}$ to 70.
+  • $P_{2}F$. Employing node fusing (§4.1) on $P_{2}$ removes all batch normalizations, which lowers $\mathcal{L}$ to 53.
+  • $P_{2}R$.
+Applying weight redistribution (§4.2) to $P_{2}$ instead yields $\mathcal{L}=35$.
+The highest-order coefficients in batch normalizations and polynomial activations are normalized to one, simplifying them to $x+b_{0}$ ($\mathcal{L}=0$) and $x^{2}+c_{1}x+c_{0}$ ($\mathcal{L}=1$); the average-pooling divisor is likewise normalized to one ($\mathcal{L}=0$).
+These simplifications give $\mathcal{L}=35$.
+  • $P_{2}FR$. Combining node fusing and weight redistribution retains $\mathcal{L}=35$, but reduces homomorphic operations, since node fusing eliminates batch normalizations.
+  • $P_{2}FRT$.
+Applying tower reuse (§4.3) to $P_{2}FR$ lets a nonlinear polynomial activation and subsequent linear kernel share a level, halving the number of levels to $\mathcal{L}=18$, following $\mathcal{L}_{P_{2}FRT}=\lceil\frac{\mathcal{L}_{P_{2}FR}}{2}\rceil$ for VGG and ResNet models.
+In summary, by combining a lower-degree polynomial activation with structural optimizations, we reduce levels by nearly a factor of five compared to prior work.
+Table 1: Number of RNS levels required for homomorphic multiplications in VGG and ResNet models.
+$P_{4}$: Degree-4 polynomial activation used in PILLAR [11].
+$P_{2}$: Degree-2 polynomial activation (§3).
+$P_{2}F$: $P_{2}$ with node fusing (§4.1).
+$P_{2}R$: $P_{2}$ with weight redistribution (§4.2).
+$P_{2}FR$: $P_{2}$ with node fusing and weight redistribution.
+$P_{2}FRT$: $P_{2}FR$ with tower reuse (§4.3).
+Model $P_{4}$ $P_{2}$ $P_{2}F$ $P_{2}R$ $P_{2}FR$ $P_{2}FRT$
+VGG-16 79 64 51 31 31 16
+ResNet-18 87 70 53 35 35 18
+ResNet-20 97 78 59 39 39 20
+ResNet-32 157 126 95 63 63 32
+5 Co-design Techniques
 5.1 Data Layout
-Data layout defines how model inputs and weights are mapped into FHE ciphertext and plaintext slots during encoding. In this work, we adopt the HW layout (Dathathri et al., 2019), which assigns
-Figure 1: Illustration of HW layout for convolution on a $3\times 3\times 3$ input and two $3\times 2\times 2$ filters, configured with padding 0 and stride 1.
-each input channel to a separate ciphertext and fills its slots with spatial values indexed by two–dimensional coordinates.
-Unlike the naive layout (Gilad-Bachrach et al., 2016), where each ciphertext holds only one value and incurs significant overhead, HW layout is more efficient, packing $h\cdot w$ values per ciphertext, with $h$ and $w$ denoting input height and width.
-Figure 1 shows this for a $3\times 3\times 3$ input with two $3\times 2\times 2$ filters (stride 1, no padding), illustrating slot arrangements and gaps that may appear after convolution or other kernel operations.
+Data layout defines how model inputs and weights are mapped into FHE ciphertext and plaintext slots during encoding.
+In this work, we adopt the HW layout [9], which assigns each input channel to a separate ciphertext and fills its slots with spatial values indexed by two–dimensional coordinates.
+Unlike the naive layout [13], where each ciphertext holds only one value and incurs significant overhead, HW layout is more efficient, packing $h\cdot w$ values per ciphertext, with $h$ and $w$ denoting input height and width.
+Figure 4 shows this for a $3\times 4\times 4$ input with a $3\times 2\times 2$ filters (stride 1, no padding), illustrating slot arrangements and gaps that may appear after convolution or other kernel operations.
 These gaps could be removed by remapping, but that requires additional multiplications and increases multiplicative depth.
@@
 HW layout does not always fully utilize slots, especially when input channels have far fewer values than available slots.
-More compact layouts such as CHW layout (Dathathri et al., 2019) and its variants (Lee et al., 2022) address this by mapping multiple channels into one ciphertext, which reduces ciphertext counts and decreases the number of multiplications and additions, ultimately leading to faster inference.
+More compact layouts such as CHW layout [9] and its variants [20] address this by mapping multiple channels into one ciphertext, which reduces ciphertext counts and decreases the number of multiplications and additions, ultimately leading to faster inference.
 However, PPML relies on polynomial approximations that reduce accuracy compared to plaintext ReLU models.
@@
 To offset the added latency, we introduce a parameter clustering technique.
+Figure 4: Illustration of HW layout for convolution on a $3\times 4\times 4$ input and a $3\times 2\times 2$ filter, configured with padding 0 and stride 1.
 5.2 Clustering of Convolution Parameters
-Convolution layers account for most parameters in deep neural networks, and their weight handling under FHE is costly since each weight must be repeatedly encoded to match varying ciphertext layouts and levels.
-As these depend on kernel position and layer structure, the same scalar weight is often redundantly encoded.
-Pre-encoding all weight–layout combinations would lead to prohibitive memory usage, while on-demand encoding reduces memory but adds heavy computational overhead.
-To mitigate this, weights are limited to a small fixed set of representative values.
+Convolution layers account for the majority of parameters in deep neural networks, and their handling under FHE is particularly costly.
+Each convolution weight must be encoded repeatedly to match varying ciphertext layouts and levels, which depend on kernel position and layer structure.
+As a result, the same scalar weight is often redundantly encoded many times.
+Pre-encoding all possible weight–layout combinations would lead to prohibitive memory usage, while encoding weights on-demand reduces memory usage but introduces heavy computational cost.
+To mitigate this overhead, we restrict weights to a small, fixed set of representative values.
 Each representative is encoded once into a plaintext codebook, and during inference, the appropriate plaintext is retrieved rather than recomputed.
-Time and memory then scale with codebook size instead of the number of weights.
-Clustering provides a way to build this representative set by grouping similar weights and replacing each with its closest centroid, bounding the number of unique encodings.
-Formally, let a polynomial network contain $L$ convolution layers.
-The weight tensor of layer $l$ is $W^{(l)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$, where $O_{l}$ is the number of output channels, $I_{l}$ the number of input channels, and $H_{l}\times W_{l}$ the kernel size.
-Flattening all convolution weights gives $\theta=(\theta_{1},\dots,\theta_{P})\in\mathbb{R}^{P},$ where $P$ is the total number of parameters.
+Consequently, both runtime and memory scale with the codebook size rather than with the total number of weights.
+Clustering provides a mechanism for constructing such a representative set by grouping similar weights and replacing each with its nearest centroid, thereby bounding the number of unique encodings.
+Formally, consider a polynomial network with $L$ convolution layers.
+The weight tensor of layer $l$ is denoted by $W^{(l)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$, where $O_{l}$ is the number of output channels, $I_{l}$ the number of input channels, and $H_{l}\times W_{l}$ the kernel size.
+Flattening all convolution weights gives $\theta=(\theta_{1},\dots,\theta_{P})\in\mathbb{R}^{P},$ where $P$ is the total number of convolution parameters.
 To compress the model, each $\theta_{j}$ is replaced by its nearest codebook value from $\mathcal{C}=\{c_{1},\dots,c_{k}\}\subset\mathbb{R}$, where $k\ll P$ is the number of representatives in the codebook.
@@
 Quantization is applied elementwise, so each convolutional kernel retains its shape while its values are drawn from the codebook.
-The choice of clustering strategy determines how the codebook is constructed and how assignments are made.
+The choice of clustering strategy determines how the codebook is constructed and how assignments are performed.
 Full Clustering.
@@
 All convolution parameters, regardless of layer or kernel position, are quantized to the same shared set of representative values.
-This produces a uniform quantization scheme with easy implementation and a small codebook size.
+This produces a uniform quantization scheme with simple implementation and a small codebook size.
 Limitation.
-While full clustering reduces the number of distinct weight values, it does not eliminate repeated plaintext encodings.
-Each weight must still match the ciphertext layout and level of its layer, so the same scalar value appearing in different layers must be re-encoded.
-Even within a single convolution, all weights share the same level, but weights from different filter columns map to different layouts and require separate encodings.
+While full clustering reduces the number of distinct weight values, it does not eliminate redundant plaintext encodings.
+Each weight must still match the ciphertext layout and level of its layer, so the same scalar values appearing in different layers must be re-encoded.
+Even within a single convolution layer, although all weights share the same level, weights in different filter columns map to different layouts and thus require separate encodings.
 Only weights in the same column across layer filters align in layout and level.
-We define this group as a slice, which can share plaintext encoding.
-Thus, full clustering reduces codebook size but not redundancy from mismatched levels and layouts.
+We refer to such a group as a slice, which can share plaintext encoding.
+Thus, full clustering reduces codebook size but fails to address redundancy arising from mismatched levels and layouts.
 Slice Clustering.
-To address the redundancy remaining under full clustering, we refine quantization to the slice level.
-Since only weights within the same slice can share plaintext encodings, clustering is applied per slice rather than across all parameters.
-For a given layer $l$ and spatial position $s\in\{1,\dots,W_{l}\}$, the slice is $S^{(l)}_{s}=\{W^{(l)}_{o,i,h,s}|o=1,\dots,O_{l};\ i=1,\dots,I_{l};\ h=1,\dots,H_{l}\},$ which can be viewed as a vector in $\mathbb{R}^{O_{l}I_{l}H_{l}}$ (see §F.1 for an illustration).
-Each slice $S^{(l)}_{s}$ is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}\subset\mathbb{R}$, and quantization is applied elementwise: $\tilde{W}^{(l)}_{o,i,h,s}=\arg\min_{c\in\mathcal{C}^{(l)}_{s}}d(W^{(l)}_{o,i,h,s},c).$ This ensures clustering respects FHE structure while adapting to the local weight distribution of each slice.
-The additional storage from slice-specific codebooks is modest compared to the resulting gains in accuracy and efficiency.
+To address the redundancy remaining under full clustering, we refine quantization at the slice level.
+Since only weights within the same slice can share plaintext encodings, clustering is applied independently to each slice rather than across all parameters.
+For a given layer $l$ and spatial position $s\in\{1,\dots,W_{l}\}$, we define the slice $S^{(l)}_{s}=\{W^{(l)}_{o,i,h,s}|o=1,\dots,O_{l};\ i=1,\dots,I_{l};\ h=1,\dots,H_{l}\},$ which can be viewed as a vector in $\mathbb{R}^{O_{l}I_{l}H_{l}}$.
+Each slice $S^{(l)}_{s}$ is clustered independently using its own codebook $\mathcal{C}^{(l)}_{s}\subset\mathbb{R}$, and quantization is applied elementwise: $\tilde{W}^{(l)}_{o,i,h,s}=\arg\min_{c\in\mathcal{C}^{(l)}_{s}}d(W^{(l)}_{o,i,h,s},c).$ This ensures clustering respects FHE structure while adapting to the local weight distribution of each slice.
+Figure 5 illustrates slice clustering for a single convolution layer.
+Convolution filters are decomposed into vertical slices along the kernel width dimension $W_{l}$.
+Each colored slice $S^{(l)}_{s}$ is clustered independently using its own codebook.
+Each three-dimensional block represents a convolution filter associated with one output channel.
+The three axes correspond to kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth).
+The collection of such filters forms the output channels $O_{l}$.
+Slices are taken as vertical strips along the width dimension $W_{l}$, and the figure highlights them with distinct colors.
+The blue strip corresponds to slice $S^{(l)}_{1}$, the yellow strip to slice $S^{(l)}_{2}$, and the pink strip to slice $S^{(l)}_{3}$.
+Each slice $S^{(l)}_{s}$ groups together all weights $W^{(l)}_{o,i,h,s}$} at a fixed width position $s$ across output channels $O_{l}$, input channels $I_{l}$, and kernel height $H_{l}$, and is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}$.
+This illustration emphasizes that clustering is performed slice by slice, rather than across the entire filter volume.
+Figure 5: Illustration of single model slice clustering. Convolution filters are decomposed into slices along the kernel width, and each slice is clustered independently with its own codebook.
 5.3 Ensemble of Polynomial Network
 Inference with a single polynomial network $g(x)$ often shows high variance across training runs.
-Let $g_{(m)}(x)\in\mathbb{R}^{K}$ be the logits from the $m$-th independently trained instance.
-For a fixed input, these logits can differ significantly, especially near decision boundaries, due to (1) approximation error from the fixed-point polynomial activation $p_{d}(\cdot)$ and (2) training stochasticity such as random weight initialization and mini-batch ordering.
-As a result, different $g_{(m)}$ may predict different classes for the same input.
+Let $g_{(m)}(x)\in\mathbb{R}^{K}$ denote the logits by the $m$-th independently trained instance.
+For a fixed input, these logits can differ significantly, especially near decision boundaries, due to (1) approximation error introduced by the fixed-point polynomial activation $p_{d}(\cdot)$ and (2) training stochasticity such as random weight initialization and mini-batch ordering.
+As a result, different $g_{(m)}$ instances may predict different classes for the same input.
 To reduce this variance, we use an ensemble of $M$ polynomial networks $\{g_{(m)}\}_{m=1}^{M}$.
-Each has the same architecture and activation $p_{d}(\cdot)$ but is trained independently with different seeds and mini-batch orders.
-For input $x$, the ensemble output is the average of logits: $\overline{g}(x)=\frac{1}{M}\sum_{m=1}^{M}g_{(m)}(x)$ with the predicted class $\hat{y}=\arg\max_{k\in\{1,\dots,K\}}\overline{g}(x)_{k}.$ Averaging smooths training noise and polynomial approximation errors, producing more stable and accurate predictions.
-Each $g_{(m)}$ is trained with the regularized loss $\mathcal{L}_{\mathcal{B}}$ (Eq.
+All ensemble members share the same architecture and activation $p_{d}(\cdot)$, but are trained independently with different random seeds and mini-batch orders.
+Given an input $x$, the ensemble output is defined as the average of logits: $\overline{g}(x)=\frac{1}{M}\sum_{m=1}^{M}g_{(m)}(x)$ with the predicted class $\hat{y}=\arg\max_{k\in\{1,\dots,K\}}\overline{g}(x)_{k}.$ Logit averaging smooths both training noise and polynomial approximation errors, producing more stable and accurate predictions.
+Each $g_{(m)}$ is trained using the same regularized loss $\mathcal{L}_{\mathcal{B}}$ (Eq.
 1), ensuring compatibility with fixed-point polynomial inference under FHE.
-Crucially, this ensemble design adds no extra computation or memory overhead.
-As noted in §4.1, our HW layout leaves some ciphertext slots unused.
-We fill these slots with weights from different models, while all ensemble members share the same ciphertexts for inputs and activations.
-This reuse of ciphertext–plaintext structures avoids redundant ciphertexts and repeated encodings, making ensemble inference practical in the FHE framework.
+Crucially, this ensemble design adds no additional computation or memory overhead.
+As described in §5.1, our HW layout leaves a subset of ciphertext slots unused.
+We populate these slots with weights from different ensemble members, while all models share the same ciphertexts for inputs and activations.
+This reuse of ciphertext–plaintext structures avoids redundant ciphertexts and repeated encodings, making ensemble inference practical within the FHE framework.
 Limits of Ensemble with Clustering.
-While ensemble inference reuses unused ciphertext slots, its combination with parameter clustering creates a challenge.
-In slice clustering, each slice $S_{s}^{(l)}$ with $O_{l}\times I_{l}\times H_{l}$ weights is represented by $k$ centroids in codebook $\mathcal{C}_{s}^{(l)}$, reducing plaintext encodings to $k$ per slice.
-In an ensemble, however, each plaintext must encode parameters from all $M$ models.
-As centroids are unlikely to align across models, the count of unique encodings per slice grows as $\mathcal{O}(k^{M})$, canceling the gains of clustering and making the naive combination impractical.
+While ensemble inference reuses unused ciphertext slots, naively combining it with parameter clustering introduces a challenge.
+In slice clustering, each slice $S_{s}^{(l)}$ containing $O_{l}\times I_{l}\times H_{l}$ weights is represented by $k$ centroids in codebook $\mathcal{C}_{s}^{(l)}$, reducing the number of plaintext encodings to $k$ per slice.
+In an ensemble setting, however, each plaintext must encode parameters from all $M$ models.
+Since centroids are unlikely to align across independently trained models, the number of distinct encodings per slice grows as $\mathcal{O}(k^{M})$, negating the benefits of clustering and making the naive combination impractical.
 Slice Ensemble Clustering.
-To address the inefficiency of independent clustering, we extend slice clustering to ensembles by enforcing shared representatives across models.
-We cluster weights jointly at the same kernel position so that all models use a common codebook.
-Formally, the ensemble has $M$ polynomial networks.
-For convolution layer $l$, model $m$ has weights $W^{(l,m)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$.
-For kernel position $s\in\{1,\dots,W_{l}\}$, we extract slices: $S^{(l,m)}_{s}=\{W^{(l,m)}_{o,i,h,s}|o=1,\dots,O_{l};i=1,\dots,I_{l};h=1,\dots,H_{l}\}.$ Stacking slices across $M$ models produces $X^{(l)}_{s}=[S^{(l,1)}_{s}\;S^{(l,2)}_{s}\cdots S^{(l,M)}_{s}]\in\mathbb{R}^{N_{s}\times M},$ where $N_{s}=O_{l}I_{l}H_{l}$.
-Each row is an $M$-dimensional vector for the same weight coordinate across models (see §F.2 for an illustration).
-We cluster rows in $\mathbb{R}^{M}$: $\min_{\mathcal{C}^{(l)}_{s}\subset\mathbb{R}^{M}}\sum_{j=1}^{N_{s}}\min_{c\in\mathcal{C}^{(l)}_{s}}d(X^{(l)}_{s,j},c),$ with codebook $\mathcal{C}^{(l)}_{s}=\{c_{1},\dots,c_{k}\}$.
-Each $X^{(l)}_{s,j}$ is replaced by its nearest centroid, producing quantized slices $\tilde{X}^{(l)}_{s,j}\in\mathcal{C}^{(l)}_{s}$.
-Quantized weights $\tilde{W}^{(l,m)}$ are reconstructed by reshaping slices.
-This shared clustering aligns weights across models, mapping small variations at the same coordinate to a common centroid.
-It thus requires fewer distinct encodings, reducing codebook size and inference time while retaining the accuracy gains of ensemble.
+To address the inefficiency of independent clustering, we extend slice clustering to ensemble models by enforcing a shared set of representatives across all ensemble members.
+Specifically, we cluster weights jointly at the same kernel position so that all models use a common codebook.
+Consider an ensemble of $M$ polynomial networks.
+For convolutional layer $l$, model $m$ has weights $W^{(l,m)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$.
+For each kernel width position $s\in\{1,\dots,W_{l}\}$, we extract slices: $S^{(l,m)}_{s}=\{W^{(l,m)}_{o,i,h,s}|o=1,\dots,O_{l};i=1,\dots,I_{l};h=1,\dots,H_{l}\}.$ Stacking slices from all $M$ models produces $X^{(l)}_{s}=[S^{(l,1)}_{s}\;S^{(l,2)}_{s}\cdots S^{(l,M)}_{s}]\in\mathbb{R}^{N_{s}\times M},$ where $N_{s}=O_{l}I_{l}H_{l}$.
+Each row of $X^{(l)}_{s}$ is an $M$-dimensional vector corresponding to the same weight coordinate across all models.
+We cluster these rows in $\mathbb{R}^{M}$ by solving: $\min_{\mathcal{C}^{(l)}_{s}\subset\mathbb{R}^{M}}\sum_{j=1}^{N_{s}}\min_{c\in\mathcal{C}^{(l)}_{s}}d(X^{(l)}_{s,j},c),$ where $\mathcal{C}^{(l)}_{s}=\{c_{1},\dots,c_{k}\}$ is the shared codebook.
+Each row $X^{(l)}_{s,j}$ is replaced by its nearest centroid, producing quantized slices $\tilde{X}^{(l)}_{s,j}\in\mathcal{C}^{(l)}_{s}$.
+Quantized weights $\tilde{W}^{(l,m)}$ are then reconstructed by reshaping the corresponding slices.
+This shared clustering aligns weights across ensemble members, mapping small per-model variations at the same coordinate to a common centroid.
+It thus requires fewer distinct encodings, reducing codebook size and inference time while retaining the accuracy gains of ensembling.
+Figure 6 illustrates how slice clustering is extended from a single model to an ensemble of polynomial networks.
+Figure 6: Illustration of slice-wise clustering for an ensemble. Slices at the same kernel width position from different models are stacked and clustered jointly using a shared codebook.
+Each three-dimensional block in the figure represents the convolution filters of a single model instance, spanning kernel height $H_{l}$, kernel width $W_{l}$, and input channels $I_{l}$.
+The collection of such filters across output channels $O_{l}$ forms the layer weights $W^{(l,m)}$ for model $m$.
+For a fixed layer $l$, slices are extracted as vertical strips along the width dimension $W_{l}$, with colors indicating different slices $S^{(l,m)}_{s}$ within each model.
+Unlike the single-model case, however, we now align slices from multiple models at the same width position $s$ and stack them together.
+This produces a slice matrix $X^{(l)}_{s}\in\mathbb{R}^{N_{s}\times M}$, where each row corresponds to one weight coordinate across all $M$ models, and $N_{s}=O_{l}I_{l}H_{l}$.
+Clustering is then performed jointly in $\mathbb{R}^{M}$ across models, producing a shared codebook $\mathcal{C}^{(l)}_{s}$ for slice $s$.
+The illustration emphasizes that ensemble clustering is not applied to each model independently.
+Instead, slices are stacked across models at the same spatial position, and clustering is performed jointly.
 6 Experimental Results
-5.1 Experimental Setup
-We trained ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using PyTorch 2.4.1.
-Each experiment was repeated 10 times, and we report mean accuracy and private inference time with standard deviations.
-Private inference used a C++20 framework we developed, built on Microsoft SEAL 4.1 (SEAL, ) for leveled FHE operations and GMP 6.2.1 for multiprecision arithmetic.
-The framework includes a tool that automatically converts trained models for C++ inference.
-We use a machine with two AMD EPYC 64-core processors, 2 TB memory across 32 DDR4 DIMMs at 2933 MT/s, on Ubuntu 22.04.5 LTS with kernel 5.15.0-151-generic (x86_64) to run all experiments, including related work.
-For model training (see §2), we used hyperparameters $c=2$ and $\zeta=0.001$.
-An ablation study for both is provided in §G.
-We set $b=10$, aligned with PILLAR (Diaa et al., 2024), as it provides accurate polynomial approximation after quantization-aware fitting.
-For all clustering strategies, we used the $\ell_{2}$ norm as the distance metric, with k-means as the clustering algorithm (Lloyd, 1982).
-Details on FHE encryption parameters are in §H.1.
-Additional results, including peak memory, are reported in §H.3.
-We compare our approach against non-interactive PPML baselines, MPCNN (Lee et al., 2022) and AutoFHE (Ao & Boddeti, 2024), in terms of accuracy and inference time.
-For fairness, we use their publicly available implementations (git, 2024a; 2022) with specified encryption parameters.
+6.1 Evaluation of Training Methodology
+Training Setup.
+We evaluate our polynomial approximation strategy on the CIFAR-10, CIFAR-100, and Tiny-ImageNet (referred to as Tiny) datasets, which are widely adopted in related work on privacy-preserving neural network inference.
+Our experiments use ResNet-18, ResNet-20, and ResNet-32 architectures, and additionally include VGG-16 to assess generality beyond residual networks.
+Each experiment is repeated 10 times with different random seeds, and we report mean classification accuracy along with standard deviations.
+For model training (see §3), we set the clipping parameter to $c=2$ and the penalty strength to $\zeta=0.001$; the impact of both hyperparameters is analyzed subsequently.
+We fix the fixed-point precision to $b=10$, following PILLAR [11], as it provides accurate polynomial approximation under quantization-aware fitting.
+For all clustering strategies, we use the $\ell_{2}$ norm as the distance metric and employ k-means clustering [22].
+Impact of Clipping Parameter ($c$).
+We evaluate the effect of the clipping interval [c,c][-c,c] on model performance.
+The choice of $c$ determines the trade-off between preserving the activation distribution and truncating extreme values.
+If $c$ is too small, informative activations are excessively clipped, reducing representational capacity; conversely, if $c$ is too large, the clipping mechanism provides little stabilization, allowing unstable activations to propagate.
+Results are reported for ResNet-18 on CIFAR-10, which we present without loss of generality.
+Figure 7(a) shows the distribution of pre-activation inputs to ReLU across the network.
+The values are largely concentrated within the interval [2,2][-2,2], with only a small fraction outside this range, supporting the choice of restricting pre-activations to [2,2][-2,2].
+Figure 7(b) shows the classification accuracy obtained under different clipping intervals.
+Among the tested ranges, [2,2][-2,2] consistently achieves the highest mean accuracy across runs.
+Smaller intervals degraded performance by discarding a substantial portion of activations, while larger intervals reduce stability and result in lower accuracy.
+This trend is consistent across other dataset–architecture combinations.
+Based on these observations, we adopt [2,2][-2,2] as the clipping range in all experiments.
+Impact of Penalty Strength ($\zeta$).
+We analyze the impact of the penalty strength $\zeta$ on model performance.
+The parameter $\zeta$ controls the magnitude of the regularization term that penalizes pre-activations outside the target interval [c,c][-c,c].
+Figure 7(c) reports results ResNet-18 on CIFAR-10.
+The highest mean accuracy is obtained for $\zeta=10^{-3}$.
+Smaller values fail to provide sufficient regularization and lead to degraded accuracy, whereas larger values reduce performance by overly constraining the model.
+This trend is consistent across other dataset–architecture combinations.
+Based on these observations, we set $\zeta=10^{-3}$ in all experiments.
+Figure 7: Effects of activation clipping and regularization for ResNet-18 on CIFAR-10.
+(a) Pre-Activation value distribution across all layers, showing concentration within [-2, 2].
+(b) Accuracy across clipping ranges, with [-2, 2] performing best.
+(c) Accuracy versus penalty strength $\zeta$, with best performance at $\zeta=10^{-3}$; smaller values under-regularize the model, while larger values overly constrain it and reduce accuracy.
+Comparison to Related Work.
+We compare the training methodology of our approach against PILLAR, which represents the lowest-degree polynomial-approximation method previously shown to train deep neural networks reliably.
+We evaluate PILLAR using its open-source implementation [3].
+PILLAR was originally designed in the context of interactive PPML, which differs from our non-interactive setting, and its repository does not provide any implementation for private inference.
+For these reasons, we restrict the comparison to classification accuracy and do not evaluate inference time.
+Table 2 reports classification accuracies (mean $\pm$ standard deviation) across all evaluated datasets and architectures using ReLU, PAPER (our method), and two versions of PILLAR.
+The original PILLAR method employs a degree-4 polynomial activation.
+For a fair comparison, we additionally adapt this implementation to use a degree-2 polynomial, allowing a direct comparison with PAPER at the same polynomial degree while keeping the original training procedure unchanged.
+As shown in Table 2, when averaged across all datasets and architectures, the adapted degree-2 variant of PILLAR exhibits a substantial degradation in accuracy, with an average drop of 7.99% relative to ReLU and notably high variance, particularly on more challenging datasets such as CIFAR-100 and Tiny-ImageNet.
+The original degree-4 variant of PILLAR is more stable but still underperforms ReLU, with an average accuracy gap of 6.54%.
+In contrast, PAPER with degree-2 polynomial activation remains much closer to ReLU, incurring an average accuracy drop of only 1.84% across all datasets and architectures.
+Importantly, PAPER demonstrates consistently stable training behavior across all evaluated settings, avoiding the large run-to-run fluctuations seen in both variants of PILLAR.
+These results demonstrate that while PILLAR either degrades significantly when reduced to a degree-2 configuration or remains less accurate in its intended degree-4 configuration, PAPER enables accurate and stable training with quadratic polynomial activations.
+Overall, these results indicate that the accuracy gap typically associated with low-degree polynomial approximations arises primarily from limitations in training methodology, and can be substantially mitigated through the regularization techniques introduced in PAPER.
+Table 2: Classification accuracy (%, mean $\pm$ standard deviation) on CIFAR and Tiny-ImageNet datasets for VGG and ResNet models using ReLU, PAPER (degree-2), and two versions of PILLAR.
+Method VGG-16 ResNet-18 ResNet-20 ResNet-32
+CIFAR-10
+ReLU 92.57 $\pm$ 0.36 94.57 $\pm$ 0.23 94.84 $\pm$ 0.25 95.23 $\pm$ 0.30
+PAPER 90.41 $\pm$ 0.46 93.73 $\pm$ 0.27 94.06 $\pm$ 0.26 93.86 $\pm$ 0.24
+PILLAR ($d=2$) 88.32 $\pm$ 0.19 91.78 $\pm$ 0.13 90.43 $\pm$ 0.40 90.40 $\pm$ 1.73
+PILLAR ($d=4$) 90.76 $\pm$ 0.18 93.26 $\pm$ 0.15 90.98 $\pm$ 0.29 92.24 $\pm$ 0.27
+Method ResNet-18 ResNet-20 ResNet-32 ResNet-32
+CIFAR-100 Tiny
+ReLU 76.22 $\pm$ 0.44 74.96 $\pm$ 0.36 76.58 $\pm$ 0.40 61.37 $\pm$ 0.36
+PAPER 74.96 $\pm$ 0.51 73.89 $\pm$ 0.37 73.93 $\pm$ 0.53 56.78 $\pm$ 0.42
+PILLAR ($d=2$) 71.92 $\pm$ 0.28 62.30 $\pm$ 3.96 66.38 $\pm$ 1.78 40.85 $\pm$ 0.68
+PILLAR ($d=4$) 74.81 $\pm$ 0.22 62.55 $\pm$ 7.51 66.10 $\pm$ 1.12 43.29 $\pm$ 0.83
+Note: The original PILLAR method uses a degree-4 polynomial [11]; the degree-2 variant is our adaptation of its open-source code for comparison.
+6.2 Evaluation of Private Inference
+Private Inference Setup.
+Private inference uses an in-house C++20 framework, built on Microsoft SEAL 4.1 [32] for leveled FHE operations and GMP 6.2.1 for multiprecision arithmetic.
+The framework includes a tool to automatically convert trained models for C++ inference.
+All experiments, including related work, were run on a machine with two AMD EPYC 64-core processors, 2 TB memory across 32 DDR4 DIMMs at 2933 MT/s, on Ubuntu 22.04.5 LTS.
+For PAPER, we set $N=2^{16}$ for ResNet-32 and $N=2^{15}$ for the other models, as these are the smallest polynomial degrees enabling correct inference with $\approx$128-bit security.
+Details on FHE encryption parameters are in §C.
+We compare our approach against state-of-the-art non-interactive PPML baselines, MPCNN [20] and AutoFHE [5], in terms of accuracy and inference time, using their publicly available implementations [2, 1] as is.
 Both baselines implement ResNet-20 on CIFAR-10 and ResNet-32 on CIFAR-10 and CIFAR-100, all with $N=2^{16}$.
-We also trained plaintext models using standard ReLU activations to serve as a reference for accuracy.
-In this section, we mainly report results for ResNet-20 and ResNet-32, since these allow direct comparison with baselines.
-Results for ResNet-18 are provided in §H.2.
-5.2 Summary of Results
-Accuracy of Polynomial Approximation.
-Training with the polynomial approximation described in §2 shows that, on average across CIFAR-10 and CIFAR-100, our degree-2 polynomial achieves accuracy only about 1.3% lower than ReLU, while PILLAR with degree-4 polynomial is about 5.4% lower.
-A degree-2 adaptation of PILLAR shows an average drop of 6.5%.
-These results demonstrate that our method can use low-degree polynomials effectively, while PILLAR does not reach comparable accuracy even with higher degrees.
-A detailed comparison is provided in §G.3.
-[…]
+We also train plaintext models using standard ReLU activations to serve as reference for accuracy.
+Figure 8 summarizes private inference time versus accuracy across several CNN models and datasets, comparing clustering methods, ensembles, and baselines.
 Figure 8: Accuracy vs.
-private inference time for ResNet-20/32 on CIFAR-10/100.
+private inference time for VGG and ResNet models on CIFAR and Tiny-ImageNet datasets.
 Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model).
-We compare Standard ensembles (no clustering), Slice ensembles (slice-wise clustering), and Full Clustering.
+We compare Standard ensembles (no clustering), Slice ensembles (Slice Clustering), and Full Clustering.
 Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering.
 Baselines include AutoFHE (numbers = bootstrapping layers) and MPCNN.
-The green dashed line shows plaintext ReLU accuracy.
+The green line shows plaintext ReLU accuracy.
+[…]
 Ensemble Evaluation.
-We evaluate ensembles with $M=\{1,2,4\}$ models.
-Figure 2 (middle row) shows accuracy and inference time for two approaches: Standard-$M$ (ensembles without clustering) and Slice-$M$ (ensembles with slice clustering as in §4.3).
+In Figure 8 (rows 2 and 5), we evaluate ensembles with $M=\{1,2,4\}$ models.
+The figure depicts accuracy and inference time for two approaches: Standard-$M$ (ensembles without clustering) and Slice-$M$ (ensembles with slice clustering as in §5.3).
 Inference time does not increase with ensemble size since all models fit in one ciphertext, requiring the same number of homomorphic operations.
@@
 Since each centroid represents an $M$-dimensional space, higher $M$ can require more centroids to capture model parameters.
-Thus, Slice-4 can underperform Slice-2 at a fixed $k$, and Slice-2 needs more centroids to match Standard-2 accuracy.
-Despite this, Slice-$M$ provides notable advantages.
-Slice-2 mostly outperforms Standard-2 and reaches accuracy comparable to Standard-4.
-This suggests clustering acts as an implicit regularizer, mitigating overfitting.
-Comparison with Related Work.
-Figure 2 (bottom row) presents Pareto fronts of Slice Clustering in terms of accuracy and private inference time, compared with AutoFHE and MPCNN.
-For reference, we also report Standard-4 and plaintext ReLU accuracy.
+Thus, Slice-4 can underperform Slice-2 at a fixed $k$, and Slice-2 needs more centroids to match Standard-2 accuracy than either method with $M=1$.
+Despite this, Slice-$M$ is advantageous as it produces models with higher accuracy.
+For instance, Slice-2 outperforms Standard-2 and reaches accuracy comparable to Standard-4.
+This suggests clustering acts as an implicit regularizer that mitigates overfitting.
+Accuracy of Polynomial Approximation.
+Figure 8 (rows 3 and 6) presents Pareto fronts for Slice Clustering in terms of accuracy and private inference time, and compares them against AutoFHE and MPCNN.
+For reference, the figure also reports Standard-4 and plaintext ReLU accuracy.
+Beyond the gains obtained from our training methodology, Slice Ensemble Clustering5.3) further improves accuracy by ensembling multiple models within a single ciphertext.
+For the simpler CIFAR-10 dataset, Slice Clustering achieves accuracy ranging from 1.1% below to 0.3% above ReLU.
+For the more complex CIFAR-100 and Tiny-ImageNet datasets, our approach consistently surpasses ReLU by 0.7–2.6%.
+These results demonstrate that our method enables the effective use of low-degree polynomial activations without sacrificing accuracy.
+Comparison to Related Work.
 For ResNet-20, where our models operate with $N=2^{15}$, Slice Clustering achieves consistently faster inference than related work (up to $4\times$).
@@
 Inference time does not change for our approach for different datasets, whereas AutoFHE slows down because it requires a high number of bootstrapping operations to maintain accuracy.
-The main factors affecting our inference time relative to related work are: The main factors affecting our inference time relative to related work are: (i) the polynomial activation (§2) and structural optimizations (§3) reduce both multiplicative depth and number of RNS moduli, allowing execution under LFHE without bootstrapping; (ii) our clustering technique (§4.2) further reduces the number of unique parameters, thereby limiting plaintext encodings per convolution; (iii) for ResNet-20 (and ResNet-18), we operate at a smaller polynomial degree, which directly accelerates homomorphic operations; (iv) our choice of data layout (§4.1) deliberately sacrifices some latency to better exploit ciphertext slots for accuracy, but above optimizations offset this overhead.
+The main factors affecting our inference time relative to related work are: (i) the polynomial activation (§3) and structural optimizations (§4) reduce both multiplicative depth and number of RNS moduli, allowing execution under LFHE without bootstrapping; (ii) our clustering technique (§5.2) further reduces the number of unique parameters, thereby limiting plaintext encodings per convolution; (iii) for ResNet-20, we operate at a smaller polynomial degree, which directly accelerates homomorphic operations; (iv) our choice of data layout (§5.1) deliberately sacrifices some latency to better exploit ciphertext slots for accuracy, but the above optimizations offset this overhead.
 In terms of accuracy, our approach consistently outperforms related work thanks to our training strategy (§3).
-In fact, it closely matches plaintext ReLU models, effectively eliminating the accuracy gap typically observed in polynomial approximations of ResNets.
+In fact, it closely matches plaintext ReLU models, effectively eliminating the accuracy gap typically observed in polynomial approximations of CNNs.
+Scalability.
+The training methodology and optimization techniques exhibit scalability along multiple dimensions: they (i) are applicable to different CNN architectures, as demonstrated with VGG and ResNet models; (ii) extend to deeper networks, allowing more layers before bootstrapping, as shown with ResNet-32, whereas prior LFHE works were limited to AlexNet and VGG-11 [12]; (iii) support datasets with both a small (CIFAR-10) and a large number of classes (CIFAR-100); and (iv) accommodate higher-resolution images (Tiny-ImageNet) with no inference slowdown provided the number of ciphertexts for data encoding remains unchanged.
 7 Conclusion
 […]
-Appendix A Leveled Fully Homomorphic Encryption
-Homomorphic Encryption is a special type of encryption that allows meaningful computations to be performed directly on encrypted data.
-In this work, we employ the Cheon-Kim-Kim-Song (CKKS) (Cheon et al., 2017) encryption scheme, specifically the Residue Number System (RNS) variant (Cheon et al., 2018), which supports approximate arithmetic over encrypted real or complex numbers.
-CKKS is an FHE scheme based on the Ring-Learning With Errors (RLWE) problem (Lyubashevsky et al., 2013).
-The scheme operates across three domains: message $\mathcal{M}=\mathbb{C}$, plaintext $\mathcal{P}=R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$, and ciphertext $\mathcal{C}=R_{Q}\times R_{Q}$ spaces.
-The encoding function $encode(\cdot)=\mathcal{M}^{N/2}\mapsto\mathcal{P}$ maps a vector of $N/2$ complex numbers into a degree-$N$ polynomial in the plaintext space.
-The encryption function $encrypt(\cdot)=\mathcal{P}\mapsto\mathcal{C}$ transforms a plaintext into a ciphertext.
-In RNS-CKKS, the ciphertext modulus $Q$ is decomposed into $\mathcal{L}+1$ smaller primes $Q=\prod_{i=0}^{\mathcal{L}}{q_{i}}$, where $\mathcal{L}$ denotes the level.
-A polynomial in $R_{Q}$ is represented as $\mathcal{L}+1$ polynomials in $R_{q_{i}}\forall i\in[0,\mathcal{L}]\cap\mathbb{Z}$, which allows computations using native 64-bit integer arithmetic rather than costly arbitrary-precision operations.
-CKKS supports three homomorphic operations: addition, multiplication, and rotation.
-Additions and multiplications act element-wise on the encoded vector, similar to SIMD vector operations.
-The rotation operation rotates the vector of $N/2$ encoded messages by a step $s\in\mathbb{Z}_{N/2}$.
-During encoding, coefficients are scaled by a factor $\Delta$, rounded to the nearest integers, and reduced modulo $Q$, making CKKS a fixed-point FHE scheme.
-Multiplying two encoded messages scales the result by $\Delta^{2}$, requiring a rescale operation to restore the scale to $\Delta$.
-Rescaling removes one RNS modulus, reducing the level by 1.
-It also accumulates approximation errors, since the RNS modulus being discarded and the scaling factor $\Delta$ are not equal.
-When the level reaches zero, bootstrapping can reset the ciphertext to a higher level, enabling unbounded computation.
-However, bootstrapping is computationally expensive and avoided when possible.
-FHE without bootstrapping is called Levelled Fully Homomorphic Encryption (LFHE).
-In LFHE, the encryption parameters $(N,Q)$ are chosen to provide the desired security and sufficient levels for a predefined arithmetic circuit with known multiplicative depth.
-To the best of our knowledge, this work is the first to apply LFHE to deep CNN models like ResNet-20 and ResNet-32.
-Appendix B Challenges of Using Polynomial Activation
-ReLU, defined as $ReLU(x)=\max(0,x)$, is one of the most widely used activation functions in deep learning due to its simplicity and computational efficiency.
-Its piecewise linear structure introduces the non-linearity necessary for neural networks to capture complex patterns in data.
-While ReLU is straightforward to implement in plaintext settings, it poses significant challenges in FHE environments due to its reliance on conditional branching.
-FHE can evaluate only polynomial functions as it supports only additions and multiplications, thus, any FHE representation of ReLU must be a polynomial approximation.
-Evidently, the higher degree the polynomial, the more precise the approximation.
-However, the multiplicative depth of the activation functions grows logarithmically to the degree of the polynomial used for approximation.
-Precisely, the multiplicative depth can be computed as $\delta=\lceil\log_{2}{d}+1\rceil$, where $d$ denotes the polynomial degree and the +1+1 refers to the coefficient multiplication.
-In the literature, the lowest multiplicative depth for a polynomial approximation of ReLU is achieved by PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial of the form $\sum_{i=0}^{4}{c_{i}x^{i}}$ with $\delta=3$.
-The theoretical minimum multiplicative depth for an activation function is one, achieved with a polynomial of the form $x^{2}+c_{1}x+c_{0}$.
-Polynomials like this have been effectively used in shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), but failed to work for deeper models due to the escaping activation problem (Garimella et al., 2021).
-In addition, polynomial approximations are inherently limited to bounded intervals of the input domain.
-When substituted directly for ReLU during model training, these approximations often lead to a significant drop in model accuracy (Garimella et al., 2021; Hussain et al., 2021; Diaa et al., 2024).
-This section analyzes two primary causes of this degradation: escaping activations, where intermediate layer outputs fall outside the polynomial’s approximation interval and coefficient truncation, which results from representing polynomial coefficients in fixed-point format to ensure compatibility with FHE arithmetic.
-B.1 Escaping Activation Problem
-Let $p_{d}(x)=\sum_{k=0}^{d}a_{k}x^{k}$ denote a degree-$d$ polynomial approximation of the ReLU function, where each coefficient $a_{k}\in\mathbb{R}$ for all $k\in\{0,\dots,d\}$.
-These coefficients are typically obtained by minimizing the least-squares error between the polynomial and the ReLU function over a finite set of real-valued sample points $\{x_{i}\}_{i=1}^{N}\subset[-c,c]$, for some parameter $c>0$:
-$\min_{a_{0},\ldots,a_{d}}\;\sum_{i=1}^{N}\left(ReLU(x_{i})\;-\;p_{d}(x_{i})\right)^{2}$
-The resulting polynomial $p_{d}(x)$ provides a close approximation to ReLU within the interval [c,c][-c,c].
-However, outside the interval, the polynomial behavior differs significantly from that of ReLU.
-ReLU exhibits piecewise linear growth:
-$ReLU(x)\;=\;\begin{cases}0,&\text{if }x\leq 0\\ x,&\text{if }x>0\end{cases}\quad\Rightarrow\quad ReLU(x)\;=\;\mathcal{O}(x)$
-In contrast, the growth of $p_{d}(x)$ for $d\geq 2$ is polynomial, dominated asymptotically by the highest-degree term:
-$p_{d}(x)\;=\;a_{d}x^{d}\;+\;a_{d-1}x^{d-1}\;+\;\cdots\;+\;a_{1}x\;+\;a_{0}\;=\;\mathcal{O}(x^{d})$
-This fundamental mismatch in growth rates gives rise to the problem of escaping activations, first identified by Garimella et al.
-(2021).
-As inputs propagate through the layers of a neural network, the intermediate values passed into the polynomial activation can escape the intended interval [c,c][-c,c], entering regions where $p_{d}(x)$ no longer approximates ReLU accurately.
-As a consequence, the network can produce excessively large activations, which in turn cause exploding weights and rapid degradation of model performance unless specific modifications are made to the training procedure to contain them.
-B.2 Coefficient Truncation
-In privacy-preserving inference based on FHE, all computations are carried out using fixed-point arithmetic with limited precision.
-This constraint restricts the range and resolution of values that can be accurately represented, affecting both the domain of activation function inputs and the magnitudes of polynomial approximation coefficients.
-The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are typically derived via floating-point least-squares fitting.
-To enable fixed-point evaluation under FHE, these coefficients are quantized using:
-$\tilde{a}_{k}\;=\;\frac{\lfloor a_{k}\cdot 2^{b}\rceil}{2^{b}}$
-where $b\in\mathbb{N}$ represents the number of fractional bits in the fixed-point format and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer.
-The least-squares fitting procedure often produces small-magnitude coefficients, particularly for higher-degree monomials.
-If any coefficient satisfies $\left|a_{k}\right|<2^{-(b+1)}$, the quantized value $\tilde{a}_{k}$ becomes zero.
-This effectively discards the corresponding monomial $x^{k}$ from the approximation.
-This phenomenon, referred to as coefficient truncation (Diaa et al., 2024), changes the shape of the polynomial and can cause significant deviation from the intended ReLU behavior, even within the designated approximation interval [c,c][-c,c].
-Appendix C Why the Penalty Function Works?
+Ethical Considerations
+Our work focuses on improving the computational efficiency and accuracy of privacy-preserving machine learning using leveled FHE.
+All experiments use publicly available benchmark datasets such as CIFAR-10, CIFAR-100, and Tiny-ImageNet.
+The research does not involve human subjects, personal data, or interactions with live systems.
+The primary stakeholders are end users who need strong protection for sensitive data, organizations that provide encrypted inference services, and the research community.
+Our techniques reduce reliance on costly bootstrapping and allow deeper neural networks to run efficiently on encrypted inputs.
+This provides a clear societal benefit by making practical privacy-preserving machine learning more accessible.
+The method operates strictly within the security guarantees of the underlying cryptographic scheme and does not introduce new vulnerabilities or mechanisms that could enable data misuse or surveillance.
+It focuses solely on improving the performance of established secure computation methods.
+All implementation details and model configurations are included to support reproducibility.
+Because the research does not involve human subjects, sensitive information, or activities that require institutional oversight, no IRB review is required.
+Appendix A Why the Penalty Function Works?
 Lemma 1 (Pre-activation Update Decomposition).
@@
 $\ell(x)=\ell_{\mathrm{CE}}\left(g(x),y\right)\;+\;\zeta\;R\left(z_{p}^{(l)}\right),\quad R\left(z_{p}^{(l)}\right)\;=\;\left\|d^{(l)}\right\|_{2}$
-The first term is the standard cross‑entropy loss computed from the network output $g(x)$.
+The first term is the standard cross‑entropy loss from the network output $g(x)$.
 The second term is a layer-specific penalty that acts only on the pre-activations at layer $l$, penalizing values that fall outside the clipping interval.
@@
 $\;=\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}$
-Putting it all together, we obtain the total gradient-based update for $W^{(l)}$
+Putting it all together, we obtain the total update for $W^{(l)}$
 $\Delta W^{(l)}\;=\;-\eta\left(g^{(l)}\;+\;\zeta\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}}\right)\;\left[h^{(l-1)}_{p_{d}}(x)\right]^{\top}$
 By definition, the change in pre-activation $z_{p}^{(l)}$ after one gradient descent step is given by
-$\Delta z_{p}^{(l)}$ $\;=\;\overline{z}_{p}^{(l)}\;-\;z_{p}^{(l)}$
-$\;=\;\overline{W}^{(l)}\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
+$\Delta z_{p}^{(l)}$ $\;=\;\overline{z}_{p}^{(l)}\;-\;z_{p}^{(l)}\;=\;\overline{W}^{(l)}\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
 $\;=\;\left(W^{(l)}\;+\;\Delta W^{(l)}\right)\;h^{(l-1)}_{p_{d}}(x)\;-\;W^{(l)}\;h^{(l-1)}_{p_{d}}(x)$
@@
 $\;=\;\underbrace{\left(-\eta\right)\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;g^{(l)}}_{\displaystyle\Delta z_{\rm CE}^{(l)}}+\underbrace{\left(-\eta\right)\zeta\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\frac{d^{(l)}}{\|d^{(l)}\|_{2}}}_{\displaystyle\Delta z_{\rm pen}^{(l)}}$
-we obtain the penalty components $\Delta z_{\rm CE}^{(l)}$ and $\Delta z_{\rm pen}^{(l)}$ as mentioned in the lemma. ∎
+we obtain the penalty components $\Delta z_{\rm CE}^{(l)}$ and $\Delta z_{\rm pen}^{(l)}$. ∎
 Lemma 2 (Clipping Gradient Pullback).
@@
 $\boxed{\left\langle\Delta z_{\rm pen}^{(l)},\;d^{(l)}\right\rangle=-\;\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\left\|d^{(l)}\right\|_{2}\;<\;0}$
-whenever $\left\|d^{(l)}\right\|_{2}\neq 0$. Therefore, $\Delta z_{\mathrm{pen}}^{(l)}$ “pulls back” each element of $z_{p}^{(l)}$ lying outside [c,c][-c,c] towards the clipping interval.
-The clipping residual $d^{(l)}$ is defined as the difference between the current pre‑activation $z_{p}^{(l)}$ and its clipped version.
+whenever $\left\|d^{(l)}\right\|_{2}\neq 0$. Therefore, $\Delta z_{\mathrm{pen}}^{(l)}$ “pulls back" each element of $z_{p}^{(l)}$ lying outside [c,c][-c,c] towards the clipping interval.
+The clipping residual $d^{(l)}$ is defined as the difference between the pre‑activation $z_{p}^{(l)}$ and its clipped version.
 $\Delta z_{\rm pen}^{(l)}$ is the component in $\Delta z_{p}^{(l)}$ induced by the clip-range penalty.
-Taking the inner product of $\Delta z_{\rm pen}^{(l)}$ with the clipping residual $d^{(l)}$ gives
+Taking the inner product of $\Delta z_{\rm pen}^{(l)}$ with $d^{(l)}$ gives
 $\left\langle\Delta z_{\rm pen},\;d^{(l)}\right\rangle$ $\;=\;\left\langle-\;\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{d^{(l)}}{\left\|d^{(l)}\right\|_{2}},\;d^{(l)}\right\rangle$
 $\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\left\langle\frac{d^{(l)}}{\|d^{(l)}\|_{2}},\;d^{(l)}\right\rangle$
-$\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{1}{\left\|d^{(l)}\right\|_{2}}\;\langle d^{(l)},\,d^{(l)}\rangle$
 $\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\frac{1}{\left\|d^{(l)}\right\|_{2}}\,\left\|d^{(l)}\right\|_{2}^{2}$
 $\;=\;-\eta\;\zeta\;\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}\;\left\|d^{(l)}\right\|_{2}$
-Each factor on the right-hand side: $\eta$ (learning rate), $\zeta$ (regularization parameter), $\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$ (squared norm of the preceding hidden representation) and $\left\|d^{(l)}\right\|_{2}$ (magnitude of the clipping residual) is strictly positive whenever any pre‑activation lies outside [c,c][-c,c].
+Each factor on the right-hand side: $\eta$ (learning rate), $\zeta$ (regularization parameter), $\left\|h^{(l-1)}_{p_{d}}(x)\right\|_{2}^{2}$ (squared norm of the preceding hidden representation) and $\left\|d^{(l)}\right\|_{2}$ (magnitude of the clipping residual) are strictly positive whenever any pre‑activation lies outside [c,c][-c,c].
 Hence the entire product is strictly negative
@@
 ∎
-Appendix D Training Stability during Polynomial Training
-D.1 Pre-Activation Clipping
-During early training stages, network weights remain close to their random initialization, and the optimization process does not immediately constrain the pre-activation values to lie within the target approximation interval [c,c][-c,c].
-As a result, unbounded pre-activations may arise before the model learns to keep them within range, leading to training instability that can degrade model behavior beyond recovery.
-To prevent this, we restrict pre-activations to the interval [c,c][-c,c] before evaluating the polynomial activation, which is achieved through a clipping strategy $\mathrm{clip}(z;[-c,c])$ (Diaa et al., 2024):
-$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(\mathrm{clip}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x);[-c,c]\right)\right).$
-Importantly, this clipping operation is applied only during training to stabilize learning.
-It is performed after computing the clip-range penalty to ensure that gradients from the regularization term, which encourages the network to keep pre-activations within the approximation interval [c,c][-c,c], are preserved and not masked by the clipping.
-At inference time, the clipping function is removed:
-$h^{(l)}_{p_{d}}(x)\;=\;p_{d}\left(W^{(l)}h^{(l-1)}_{p_{d}}(x)\right).$
-The model, having learned to constrain pre-activations during training, is expected to remain within the approximation interval without explicit clipping at inference.
-D.2 Regularization Warm-up
-While activation regularization and pre-activation clipping are both necessary for training stability, applying the full regularization strength from the outset can lead to numerical instability, particularly in larger models.
-In the early epochs, many pre-activation values lie outside the target interval [c,c][-c,c] across several layers.
-The clip-range penalty adds a contribution from every such layer, so the penalty term becomes very large.
-In extreme cases, this can cause the total loss to become numerically undefined.
-To address this issue, the regularization strength $\zeta$ is progressively increased over the initial training epochs, which is implemented through a regularization warm up schedule (Diaa et al., 2024).
-Let $T_{warm}$ denote the total number of warm-up epochs.
-For each epoch $t$, we define a time-dependent regularization weight $\zeta_{t}$ as:
-$\zeta_{t}\;=\;\begin{cases}\alpha_{t}\cdot\zeta&\text{if }t\leq T_{warm}\\ \zeta&\text{if }t>T_{warm}\end{cases},$
-where $\{\alpha_{t}\}_{t=1}^{T_{warm}}$ is a fixed sequence of scaling factors satisfying $0<\alpha_{1}<\cdots<\alpha_{T_{warm}}<1$.
-In practice, we empirically find that setting $T_{warm}=4$ provides a stable convergence.
-We use a simple predefined sequence of scaling factors: $\alpha=\left\{\frac{1}{100},\frac{1}{50},\frac{1}{10},\frac{1}{5}\right\}$, which produces the epoch-wise regularization strength: $\zeta_{t}\;\in\;\left\{\frac{\zeta}{100},\frac{\zeta}{50},\frac{\zeta}{10},\frac{\zeta}{5},\zeta,\zeta,\dots\right\}$.
-This warm-up schedule ensures that the regularization penalty is introduced progressively in the initial training epochs without suffering from exploding loss values.
-Appendix E Additional Details on Structural Optimizations
-E.1 Detailed Derivations of Node Fusing
+Appendix B Details on Structural Optimizations
+B.1 Derivations of Node Fusing
 Batch normalization in linear form is given by
@@
 $P(z)$ $=c_{2}z^{2}+c_{1}z+c_{0}$
-$=c_{2}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)^{2}+c_{1}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)+c_{0}.$
+$=c_{2}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)^{2}$
+$\quad+c_{1}\big(b_{X1}x+b_{Y1}y+b_{X0}+b_{Y0}\big)+c_{0}.$
 Expanding the square term
-$z^{2}$ $=b_{X1}^{2}x^{2}+b_{Y1}^{2}y^{2}+2b_{X1}b_{Y1}xy+2b_{X1}(b_{X0}+b_{Y0})x+2b_{Y1}(b_{X0}+b_{Y0})y+(b_{X0}+b_{Y0})^{2}.$
+$z^{2}$ $=b_{X1}^{2}x^{2}+b_{Y1}^{2}y^{2}+2b_{X1}b_{Y1}xy+2b_{X1}(b_{X0}+b_{Y0})x$
+$\quad+2b_{Y1}(b_{X0}+b_{Y0})y+(b_{X0}+b_{Y0})^{2}.$
 Substituting and grouping terms by monomials
 $P(z)$ $=\underbrace{c_{2}b_{X1}^{2}}_{d_{X2}}x^{2}+\underbrace{c_{2}b_{Y1}^{2}}_{d_{Y2}}y^{2}+\underbrace{2c_{2}b_{X1}b_{Y1}}_{d_{XY}}xy$
-$\quad+\underbrace{\big[b_{X1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{X}}x+\underbrace{\big[b_{Y1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{Y}}y$
+$\quad+\underbrace{\big[b_{X1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{X}}x$
+$\quad+\underbrace{\big[b_{Y1}\big(2c_{2}(b_{X0}+b_{Y0})+c_{1}\big)\big]}_{d_{Y}}y$
 $\quad+\underbrace{\big[c_{2}(b_{X0}+b_{Y0})^{2}+c_{1}(b_{X0}+b_{Y0})+c_{0}\big]}_{d_{0}}.$
@@
 Hence $S(x,y)=d_{X2}x^{2}+d_{Y2}y^{2}+d_{XY}xy+d_{X}x+d_{Y}y+d_{0}$ with the coefficients above.
-E.2 Detailed Derivations of Weight Redistribution
-E.2.1 Update Forward
+B.2 Derivations of Weight Redistribution
+B.2.1 Update Forward
 Donors. Average Pooling. Start with $\mu(x)=k^{-1}\sum_{i}{x_{i}}$ and its normalized form $\bar{\mu}(x)=\sum_{i}x_{i}$. We must find $\upsilon$ such that $\upsilon\bar{\mu}(x)=\mu(x)$ is valid
@@
 $\upsilon\bar{c}_{ij}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
-$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-1}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
+$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-1}\qquad\begin{aligned} &\forall i\in\{0,\dots,d-1\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
 Receivers.
@@
 $\bar{c}_{ij}\upsilon^{-i}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
-$\bar{c}_{ij}$ $=c_{ij}\upsilon^{i}\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
-E.2.2 Update Backward
+$\bar{c}_{ij}$ $=c_{ij}\upsilon^{i}\qquad\begin{aligned} &\forall i\in\{0,\dots,d\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
+B.2.2 Update Backward
 Donors. Average Pooling. Start with $\mu(x)=k^{-1}\sum_{i}{x_{i}}$ and its normalization $\bar{\mu}(x)=\sum_{i}x_{i}$. We must determine $\upsilon$ such that $\bar{\mu}(\upsilon x)=\mu(x)$ holds
@@
 $\bar{c}_{ij}\upsilon^{i}x^{i}y^{j}$ $=c_{ij}x^{i}y^{j}$
-$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-i}\quad\forall i\in\{0,\dots,d-1\},\forall j\in\{0,\dots,d-i\}.$
+$\bar{c}_{ij}$ $=c_{ij}\upsilon^{-i}\qquad\begin{aligned} &\forall i\in\{0,\dots,d-1\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
 Receivers.
@@
 $\bar{c}_{ij}x^{i}y^{j}$ $=\upsilon c_{ij}x^{i}y^{j}$
-$\bar{c}_{ij}$ $=c_{ij}\upsilon\quad\forall i\in\{0,\dots,d\},\forall j\in\{0,\dots,d-i\}.$
-E.3 Example of Tower Reuse
-Suppose following an activation function, there is a convolution.
-Let modulus $q_{i}\approx\Delta^{2}$, thus $\lambda(q_{i})=2$.
-Consider an input $x$ with scaling factor $\Delta_{x}=\Delta$, giving $\lambda(x)=1$.
-Applying a quadratic activation $y=x^{2}+c_{1}x+c_{0}$ yields $\lambda(y)=2$, with $\lambda(c_{1})=1$ and $\lambda(c_{0})=2$.
-Next, a convolution $z=\sum_{i}\omega_{i}y_{i}+\alpha$ is performed.
-Since $\lambda(y)=2$, by setting $\lambda(\omega)=1$, we obtain $\lambda(z)=3$.
-As $\lambda(z)>\lambda(q_{i})$, a rescale is triggered.
-The rescale operation drops $q_{i}$ from the modulus chain, reducing the level $\Lambda(\cdot)$ by one and resetting the sublevel to one:
-$\Lambda(\downarrow z)=\Lambda(z)-1,\qquad\lambda(\downarrow z)=\lambda(z)-\lambda(q_{i})=1.$
-In terms of $\Delta$, a ciphertext $z$ with $\Delta_{z}=\Delta^{3}$ rescaled by $q_{i}\approx\Delta^{2}$ results in output scale $\Delta_{\downarrow z}\approx\Delta$.
-E.4 Level Analysis
-Table 1: Number of levels required for multiplications in ResNet models.
-$P_{4}$: Degree-4 polynomial activation used in PILLAR (Diaa et al., 2024).
-$P_{2}$: Degree-2 polynomial activation (§2).
-$P_{2}F$: $P_{2}$ with node fusing (§3.1).
-$P_{2}R$: $P_{2}$ with weight redistribution (§3.2).
-$P_{2}FR$: $P_{2}$ with node fusing and weight redistribution.
-$P_{2}FRT$: $P_{2}FR$ with tower reuse (§3.3).
-Model $P_{4}$ $P_{2}$ $P_{2}F$ $P_{2}R$ $P_{2}FR$ $P_{2}FRT$
-ResNet-18 87 70 53 35 35 18
-ResNet-20 97 78 59 39 39 20
-ResNet-32 157 126 95 63 63 32
-Table 1 summarizes the number of levels $\mathcal{L}$ required for multiplication across different ResNet variants under various optimization techniques.
-To illustrate, we focus on ResNet-18, which comprises 17 convolutional layers, batch normalization layers, and activation functions, along with a single average pooling and linear layer:
-  • $P_{4}$: Prior work achieving the lowest $\mathcal{L}$ (Diaa et al., 2024) uses a degree-4 polynomial as activation function.
-This results in $\mathcal{L}=87$, with each convolution, batch normalization, average pooling, and linear layer contributing $\mathcal{L}=1$, and the polynomial activation $\mathcal{L}=3$.
-  • $P_{2}$: By replacing the activation function with a degree-2 polynomial, as described in §2, the required level for activations drops to 2, reducing the model $\mathcal{L}$ to 70.
-  • $P_{2}F$: Applying node fusing (§3.1) to $P_{2}$ removes all batch normalizations, lowering $\mathcal{L}$ to 53.
-  • $P_{2}R$: Alternatively, applying weight redistribution (§3.2) to $P_{2}$ yields $\mathcal{L}=35$.
-This is achieved because the highest-order coefficients in batch normalization and polynomial activations are normalized to one, simplifying them to $x+b_{0}$ ($\mathcal{L}=0$) and $x^{2}+c_{1}x+c_{0}$ ($\mathcal{L}=1$), respectively.
-Moreover, the divisor in average pooling is also normalized to one, giving it $\mathcal{L}=0$.
-These simplifications collectively result in $\mathcal{L}=35$.
-  • $P_{2}FR$: Combining node fusing and weight redistribution retains $\mathcal{L}=35$.
-However, it requires fewer homomorphic operations overall, since node fusing eliminates batch normalizations entirely.
-  • $P_{2}FRT$: Finally, applying tower reuse (§3.3) to $P_{2}FR$ allows a non-linear polynomial activation followed by a linear kernel function to share the same level.
-This halves the total levels to $\mathcal{L}=18$, following $\mathcal{L}_{P_{2}FRT}=\lceil\frac{\mathcal{L}_{P_{2}FR}}{2}\rceil$ for ResNet models.
-In summary, our techniques reduce levels by nearly a factor of five compared to prior work.
-Appendix F Illustration of Slice-wise Clustering
-F.1 Single Model Slice Clustering
-Figure 3 illustrates how convolution filters are divided into slices along the width axis $W_{l}$, with each colored slice $S^{(l)}_{s}$ clustered independently using its own codebook.
-Each 3D block in the figure corresponds to a convolution filter associated with one output channel.
-The three axes correspond to kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth).
-The collection of such filters forms the output channels $O_{l}$.
-Slices are taken as vertical strips along the width axis $W_{l}$, and the figure highlights them with distinct colors.
-The blue strip corresponds to slice $S^{(l)}_{1}$, the yellow strip to slice $S^{(l)}_{2}$, and the pink strip to slice $S^{(l)}_{3}$.
-Each slice $S^{(l)}_{s}$ groups together all weights $W^{(l)}_{o,i,h,s}$} across output channels $O_{l}$, input channels $I_{l}$, and kernel height $H_{l}$ at a fixed width position $s$, and is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}$.
-This illustration emphasizes that clustering is performed slice by slice, rather than across the entire filter volume.
-Figure 3: Illustration of single model slice-wise clustering.
-F.2 Ensemble Slice Clustering
-Figure 4 illustrates how slice-wise clustering is extended from a single model to an ensemble of polynomial networks.
-Figure 4: Illustration of ensemble slice-wise clustering.
-Each 3D block corresponds to the convolution filters of one model instance, spanning kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth).
-The collection of such filters across output channels $O_{l}$ forms the layer weights $W^{(l,m)}$ for model $m$.
-For a fixed layer $l$, slices are again taken as vertical strips along the width axis $W_{l}$, with colors indicating different slices $S^{(l,m)}_{s}$ within each model.
-Unlike the single-model case, however, we now align slices from multiple models at the same width position $s$ and stack them together.
-This produces a slice matrix $X^{(l)}_{s}\in\mathbb{R}^{N_{s}\times M}$, where each row corresponds to one weight coordinate across all $M$ models, and $N_{s}=O_{l}I_{l}H_{l}$.
-Clustering is performed jointly in $\mathbb{R}^{M}$, producing a shared codebook $\mathcal{C}^{(l)}_{s}$ for slice $s$.
-The illustration emphasizes that ensemble clustering is not applied to each model independently.
-Instead, slices are stacked across models at the same spatial position, and clustering is performed jointly.
-Appendix G Ablation Study
-G.1 Ablation on Clipping Range (Value of $c$)
-We conduct an ablation study to evaluate the effect of the clipping interval [c,c][-c,c] on model performance.
-The choice of $c$ determines how much of the activation distribution is preserved versus truncated: if $c$ is too small, informative activations are excessively clipped, which can harm representational capacity, while if $c$ is too large, the clipping mechanism provides little stabilization benefit, allowing unstable activations to propagate.
-Figure 5 shows the results for CIFAR-10 with ResNet-18.
-Figure 5: Results of the clipping range ablation study on CIFAR-10 with ResNet-18.
-(a) Distribution of pre-activation inputs to ReLU across the entire network, showing concentration within [2,2][-2,2].
-(b) Test accuracy under different clipping intervals, where [2,2][-2,2] produces the best performance.
-Figure 5(a) presents the distribution of pre-activation inputs to ReLU across the entire network, where the values are concentrated within the interval [2,2][-2,2] and only a small fraction lies outside this range, supporting the choice of restricting pre-activations to [2,2][-2,2].
-Figure 5(b) reports the accuracy obtained under different clipping intervals.
-While several ranges were tested, [2,2][-2,2] consistently achieved the highest mean accuracy across runs.
-Smaller intervals degraded performance by removing too many activations, whereas larger intervals reduced stability and resulted in lower accuracy.
-The same trend holds across other dataset–architecture combinations.
-Based on these observations, we adopt [2,2][-2,2] as the clipping range in all our experiments.
-G.2 Ablation on Penalty Strength (Value of $\zeta$)
-We conduct an ablation study to evaluate the effect of the penalty strength $\zeta$ on model performance.
-The parameter $\zeta$ controls the magnitude of the regularization term that penalizes pre-activations outside the target interval [c,c][-c,c].
-Figure 6 shows the results for CIFAR-10 with ResNet-18.
-Figure 6: Results of the penalty strength ablation study on CIFAR-10 with ResNet-18, showing test accuracy across different values of $\zeta$.
-The best performance is achieved at $\zeta=10^{-3}$; smaller values fail to regularize effectively, while larger values start to decrease accuracy.
-The highest accuracy is obtained for $\zeta=10^{-3}$.
-Values smaller than $10^{-3}$ do not provide sufficient regularization and result in poor accuracy, while values greater than $10^{-3}$ start decreasing performance, indicating that excessive penalization constrains the model too strongly.
-The same trend holds across other dataset–architecture combinations.
-Based on these results, we set $\zeta=10^{-3}$ in all our experiments.
-G.3 Comparison with PILLAR
-We evaluate PILLAR using its open-source implementation (git, 2024b).
-PILLAR is designed for interactive PPML, which is a different use case from our non-interactive setting, and its repository does not provide code for private inference.
-For these reasons, we compare only in terms of accuracy and do not evaluate inference time.
-Table 2 reports accuracies (mean $\pm$ standard deviation) for ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using ReLU, PAPER (our method), and two versions of PILLAR.
-The original PILLAR is defined with a degree-4 polynomial.
-For a fair comparison, we additionally adapted this implementation to use a degree-2 polynomial, allowing a direct comparison to PAPER.
-On average, the adapted degree-2 version shows an accuracy drop of 6.5% with very high variance, particularly on CIFAR-100.
-The original degree-4 version is more stable but still falls short of ReLU by about 5.4% on average.
-In contrast, PAPER with degree-2 consistently stays much closer to ReLU, with only about a 1.3% drop on average.
-Importantly, PAPER remains stable across both datasets and all ResNet variants, avoiding the large fluctuations seen in PILLAR.
-These results demonstrate that while PAPER can exploit low-degree polynomials effectively, PILLAR either collapses when reduced to degree 2 or, in its intended degree-4 form, does not reach comparable accuracy.
-Table 2: Classification accuracies (mean $\pm$ standard deviation) on CIFAR-10 and CIFAR-100 for ResNet-18, ResNet-20, and ResNet-32 using ReLU, PAPER (degree-2), and two versions of PILLAR.
-The original PILLAR method is degree-4 (Diaa et al., 2024), while the degree-2 variant is our adaptation of its open-source implementation for comparison.
-Model Dataset ReLU PAPER PILLAR ($d=2$) PILLAR ($d=4$)
-ResNet-18 CIFAR-10 94.57 $\pm$ 0.23 93.73 $\pm$ 0.27 91.78 $\pm$ 0.13 93.26 $\pm$ 0.15
-CIFAR-100 76.22 $\pm$ 0.44 74.96 $\pm$ 0.51 71.92 $\pm$ 0.28 74.81 $\pm$ 0.22
-ResNet-20 CIFAR-10 94.84 $\pm$ 0.25 94.06 $\pm$ 0.26 90.43 $\pm$ 0.40 90.98 $\pm$ 0.29
-CIFAR-100 74.96 $\pm$ 0.36 73.89 $\pm$ 0.37 62.30 $\pm$ 3.96 62.55 $\pm$ 7.51
-ResNet-32 CIFAR-10 95.23 $\pm$ 0.30 93.86 $\pm$ 0.24 90.40 $\pm$ 1.73 92.24 $\pm$ 0.27
-CIFAR-100 76.58 $\pm$ 0.40 73.93 $\pm$ 0.53 66.38 $\pm$ 1.78 66.10 $\pm$ 1.12
-Appendix H Additional Experimental Results
-H.1 Additional Details of FHE Parameters
+$\bar{c}_{ij}$ $=c_{ij}\upsilon\qquad\begin{aligned} &\forall i\in\{0,\dots,d\},\\ &\forall j\in\{0,\dots,d-i\}.\end{aligned}$
+Appendix C Details on FHE Parameters
 For each model, we detail the encryption parameters used in our experiments.
@@
 The moduli are ordered as $q_{0},\dots,q_{L},P$, where $q_{0}$ holds the output, $q_{1},\ldots,q_{L}$ are for rescaling, and $P$ is for modulus expansion in relinearization.
-FHE terminology is provided in §A.
+FHE terminology is provided in §2.1.
+VGG-16: $N=2^{15}$, $\Delta=2^{25}$, $\log_{2}{Q}=6\cdot 50+10\cdot 49+26+54=870$.
+Moduli: {0x4000000120001, 0x3ffffffdf0001, 0x40000001b0001, 0x3ffffffd20001, 0x4000000270001, 0x3ffffffcd0001, 0x20000000b0001, 0x20000001a0001, 0x1fffffff50001, 0x1ffffffea0001, 0x20000003b0001, 0x1ffffffd40001, 0x1ffffffba0001, 0x20000005e0001, 0x1ffffffb40001, 0x20000006d0001, 0x3ff0001, 0x3fffffffd60001}.
 ResNet-18: $N=2^{15}$, $\Delta=2^{22}$, $\log_{2}{Q}=18\cdot 44+23+54=869$.
@@
 Moduli: {0x10000000060001, 0xffffffff00001, 0x10000000180001, 0xfffffffe40001, 0x10000000200001, 0xfffffffe20001, 0x100000003e0001, 0xfffffffbe0001, 0x10000000500001, 0xfffffffa60001, 0x100000006e0001, 0xfffffff820001, 0x100000007e0001, 0xfffffff480001, 0x10000000960001, 0xfffffff280001, 0x10000000c80001, 0x10000000d80001, 0xffffffed60001, 0x10000000ec0001, 0xffffffec40001, 0x10000000fc0001, 0xffffffeb00001, 0x100000010e0001, 0xffffffe9e0001, 0x10000001380001, 0xffffffe9a0001, 0x100000016a0001, 0xffffffe940001, 0x10000001bc0001, 0xffffffe6a0001, 0x10000001be0001, 0x8020001, 0x3fffffffd60001}.
-H.2 Results on ResNet18
-Figure 7: Accuracy vs.
-private inference time for ResNet-18 on CIFAR-10/100.
-Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model).
-We compare Standard ensembles (no clustering), Slice ensembles (slice-wise clustering), and Full Clustering.
-Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering.
-The green dashed line shows plaintext ReLU accuracy.
-Figure 7 shows the trade-offs between accuracy and inference time for ResNet-18 on CIFAR-10 and CIFAR-100.
-Results show similar traits to ResNet-20, leading to similar conclusions.
-In the clustering plots (left column), both Full and Slice Clustering reduce inference time compared to Standard.
-At small centroid counts ($k\leq 32$), accuracy drops noticeably due to a lack of representative centroids.
-Slice Clustering consistently achieves higher accuracy than Full Clustering, benefiting from its finer-grained parameter representation.
-When the centroid count is sufficiently large ($k\geq 64$), Slice Clustering reaches the accuracy of Standard while still reducing latency substantially.
-The middle column plots compare Standard and Slice Clustering ensembles with $M\in\{1,2,4\}$.
-Standard gains accuracy as $M$ increases, while Slice Clustering shows more variable behavior because each centroid encodes an $M$-dimensional parameter space; e.g.
-Slice-4 underperforms relative to Slice-2 for small $k$.
-Slice Clustering provides a regularization benefit and, with larger centroid counts, Slice-2 achieves higher accuracy than Standard-2 and nearly matches Standard-4.
-The right column presents the Pareto front of Slice Clustering and compares it to Standard-4.
-Slice Clustering shifts the curve toward lower latency while staying close to Standard in accuracy.
-On both datasets, Slice Clustering fully recovers plaintext ReLU accuracy and sometimes even surpasses it, demonstrating that accuracy can be preserved while reducing inference time.
-H.3 Experimental Results
-Tables 3, 4, and 5 report the accuracy, inference time, and peak memory usage of ResNet-18, ResNet-20, and ResNet-32, respectively, on the CIFAR-10 and CIFAR-100 datasets using the Standard, Full Clustering, and Slice Clustering methods.
-For Full Clustering and Slice Clustering, we varied the number of centroids as $k=2^{i}\ \forall i\in\{1,\dots,10\}$.
-For Standard and Slice Clustering, we additionally evaluated ensemble models with $M\in{1,2,4}$.
-In the case of Slice Clustering with $M\geq 2$, we extended the centroid range to $k=2^{i}\ \forall i\in\{1,\dots,12\}$ in ResNet-18 and ResNet-20.
-We did not extend this configuration to ResNet-32 due to system memory limitations (§5.1).
-Table 3: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-18.
-Standard Standard
-M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-1 93.73 0.27 3925 51 292 1 74.96 0.51 3935 79 338
-2 94.34 0.16 3930 54 314 2 76.48 0.28 3874 40 355
-4 94.69 0.12 3840 44 347 4 77.43 0.24 3929 54 390
-Full Clustering (M = 1) Full Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 10.16 0.72 404 16 197 2 1.00 0.07 414 25 240
-4 10.14 0.89 406 18 200 4 1.02 0.12 413 17 242
-8 9.91 1.48 420 31 202 8 1.08 0.17 426 23 244
-16 15.89 7.51 427 27 204 16 2.01 1.10 430 25 244
-32 70.18 14.19 439 28 204 32 32.33 13.66 441 23 246
-64 91.42 1.25 481 31 204 64 68.01 4.00 484 34 246
-128 93.24 0.29 530 34 205 128 72.62 1.04 554 41 247
-256 93.53 0.26 619 36 209 256 73.98 0.68 646 44 250
-512 93.65 0.25 719 40 221 512 74.50 0.70 773 53 263
-1024 93.69 0.28 856 34 245 1024 74.65 0.69 926 58 286
-Slice Clustering (M = 1) Slice Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 9.79 0.76 418 32 198 2 1.07 0.17 426 16 239
-4 12.66 3.02 409 23 200 4 2.76 1.17 413 13 242
-8 81.28 6.04 412 23 203 8 42.21 13.71 420 24 244
-16 92.25 0.68 420 19 204 16 69.28 2.37 443 25 246
-32 93.36 0.33 456 28 204 32 72.85 0.92 466 27 245
-64 93.62 0.27 504 30 204 64 74.31 0.77 520 31 246
-128 93.70 0.26 586 30 205 128 74.61 0.73 602 36 248
-256 93.70 0.28 696 34 214 256 74.68 0.71 724 39 256
-512 93.73 0.27 855 58 232 512 74.69 0.73 874 49 274
-1024 93.72 0.27 1016 57 266 1024 74.71 0.74 1066 66 308
-Slice Clustering (M = 2) Slice Clustering (M = 2)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 92.59 0.45 520 43 222 64 68.93 1.94 524 10 263
-128 93.85 0.15 597 38 222 128 74.47 0.59 622 23 263
-256 94.35 0.10 695 22 230 256 75.79 0.43 734 51 273
-512 94.50 0.09 854 28 250 512 76.35 0.18 871 37 292
-1024 94.60 0.08 1045 61 284 1024 76.67 0.12 1049 23 326
-2048 94.73 0.07 1164 36 394 2048 76.82 0.14 1202 72 435
-4096 94.78 0.04 1268 42 643 4096 76.90 0.08 1308 18 681
-8192 94.77 0.05 1482 59 1148 8192 76.97 0.07 1440 13 1174
-Slice Clustering (M = 4) Slice Clustering (M = 4)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 18.38 4.80 526 27 256 64 1.73 0.41 538 26 297
-128 35.79 8.02 621 45 258 128 5.31 0.94 615 27 299
-256 66.88 5.19 728 51 265 256 29.69 2.16 750 47 308
-512 84.28 2.38 866 28 285 512 57.88 2.20 885 16 326
-1024 91.47 0.46 1035 69 319 1024 68.98 1.13 1057 46 360
-2048 93.10 0.21 1211 57 428 2048 73.96 0.38 1195 64 469
-4096 94.11 0.14 1298 17 679 4096 75.93 0.22 1263 24 716
-8192 94.55 0.11 1509 22 1183 8192 76.66 0.12 1355 17 1208
-Table 4: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-20.
-Standard Standard
-M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-1 94.06 0.26 1886 30 279 1 73.89 0.37 1902 26 305
-2 94.71 0.11 1884 41 300 2 76.59 0.25 1874 27 320
-4 95.04 0.10 1886 32 330 4 77.97 0.13 1904 30 351
-Full Clustering (M = 1) Full Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 10.08 0.26 242 10 175 2 1.03 0.07 246 6 208
-4 10.08 0.32 246 6 190 4 1.08 0.23 246 9 211
-8 10.17 0.89 250 12 192 8 1.17 0.24 252 7 212
-16 14.44 6.34 260 13 192 16 2.05 1.58 258 8 213
-32 78.18 8.99 264 7 193 32 32.59 14.07 272 16 214
-64 92.46 0.95 288 10 193 64 68.18 3.24 290 11 214
-128 93.77 0.37 334 12 194 128 71.56 1.19 334 15 214
-256 94.00 0.26 407 22 198 256 73.39 0.49 392 9 218
-512 94.02 0.26 475 18 212 512 73.75 0.35 473 20 232
-1024 94.05 0.26 553 25 248 1024 73.85 0.35 563 28 264
-Slice Clustering (M = 1) Slice Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 10.05 0.41 249 18 186 2 1.12 0.22 254 10 207
-4 14.59 4.33 244 7 190 4 2.11 0.62 255 13 210
-8 85.17 2.66 247 7 192 8 49.94 5.99 264 21 212
-16 93.20 0.45 262 6 193 16 69.60 1.71 262 7 214
-32 93.87 0.32 276 9 194 32 72.98 0.45 292 12 214
-64 94.01 0.28 314 9 194 64 73.68 0.38 323 13 214
-128 94.01 0.28 366 14 194 128 73.85 0.38 377 18 215
-256 94.04 0.27 442 10 205 256 73.89 0.34 452 17 226
-512 94.05 0.26 536 26 226 512 73.89 0.36 538 10 247
-1024 94.05 0.26 612 19 288 1024 73.88 0.36 626 21 309
-Slice Clustering (M = 2) Slice Clustering (M = 2)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 92.52 0.56 312 9 209 64 68.98 0.97 321 14 230
-128 94.34 0.10 370 8 211 128 74.31 0.44 380 16 230
-256 94.81 0.08 443 20 219 256 76.08 0.25 464 7 241
-512 94.93 0.08 531 31 242 512 76.65 0.19 541 22 263
-1024 95.08 0.07 611 14 304 1024 76.99 0.11 614 22 324
-2048 95.11 0.06 679 7 458 2048 77.16 0.13 690 23 476
-4096 95.08 0.04 814 60 768 4096 77.12 0.12 746 17 781
-8192 95.08 0.03 885 37 1389 8192 77.23 0.09 897 26 1390
-Slice Clustering (M = 4) Slice Clustering (M = 4)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 12.79 2.88 332 14 240 64 1.72 0.48 341 13 261
-128 25.95 4.81 389 7 242 128 3.86 0.96 407 10 262
-256 57.23 3.93 455 21 251 256 19.97 2.42 463 19 271
-512 83.18 1.59 534 10 274 512 52.26 2.81 546 8 294
-1024 91.45 0.46 603 18 335 1024 68.37 0.63 631 14 356
-2048 93.89 0.26 648 9 489 2048 73.73 0.53 673 11 508
-4096 94.65 0.16 722 12 799 4096 76.44 0.31 765 9 813
-8192 94.90 0.10 815 17 1421 8192 77.62 0.17 855 16 1421
-Table 5: Accuracy, inference time, and peak memory for CIFAR-10 and CIFAR-100 on ResNet-32.
-Standard Standard
-M Accuracy (%) Inf. Time (s) Memory M Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-1 93.86 0.24 9405 109 1332 1 73.93 0.53 8865 132 1401
-2 94.69 0.18 8854 80 1394 2 76.79 0.34 8902 152 1418
-4 94.99 0.13 8813 63 1441 4 78.34 0.23 8907 141 1465
-Full Clustering (M = 1) Full Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 10.12 0.22 1578 63 842 2 0.98 0.07 1544 5 889
-4 10.13 0.21 1538 3 867 4 0.98 0.07 1577 20 908
-8 10.12 0.22 1588 79 879 8 0.97 0.07 1540 24 914
-16 10.35 1.35 1570 8 893 16 1.03 0.21 1547 27 929
-32 37.13 16.87 1578 3 898 32 3.91 2.63 1565 13 942
-64 86.61 3.69 1703 80 901 64 52.92 7.49 1686 44 945
-128 92.82 0.47 2131 27 902 128 70.26 1.24 1819 5 947
-256 93.70 0.32 2164 151 916 256 73.18 0.38 2013 33 954
-512 93.84 0.27 2495 164 987 512 73.75 0.37 2381 70 1020
-1024 93.86 0.26 2898 174 1155 1024 73.89 0.33 2822 53 1159
-Slice Clustering (M = 1) Slice Clustering (M = 1)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-2 10.12 1.07 1605 57 857 2 0.99 0.08 1559 11 894
-4 10.00 0.99 1573 24 879 4 1.00 0.09 1526 27 920
-8 34.61 12.19 1601 65 893 8 9.67 5.64 1586 11 936
-16 89.08 2.44 1559 6 901 16 60.50 4.61 1576 8 942
-32 93.12 0.49 1761 90 902 32 71.75 0.80 1679 19 946
-64 93.75 0.29 1833 54 905 64 73.49 0.36 1803 9 947
-128 93.84 0.26 2077 111 905 128 73.83 0.33 2027 30 946
-256 93.86 0.27 2488 185 950 256 73.91 0.33 2379 83 993
-512 93.85 0.25 2849 128 1060 512 73.93 0.33 2815 60 1104
-1024 93.86 0.24 3317 192 1373 1024 73.94 0.53 3313 57 1417
-Slice Clustering (M = 2) Slice Clustering (M = 2)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 85.12 4.61 1831 69 927 64 52.29 6.60 1791 17 968
-128 92.75 0.58 2069 79 933 128 70.20 1.26 2014 44 972
-256 94.06 0.21 2388 156 977 256 74.51 0.42 2340 52 1019
-512 94.59 0.12 2788 167 1087 512 76.30 0.25 2809 38 1128
-1024 94.76 0.12 3093 11 1399 1024 76.77 0.24 3238 64 1440
-Slice Clustering (M = 4) Slice Clustering (M = 4)
-k Accuracy (%) Inf. Time (s) Memory k Accuracy (%) Inf. Time (s) Memory
-$\mu$ $\sigma$ $\mu$ $\sigma$ (GB) $\mu$ $\sigma$ $\mu$ $\sigma$ (GB)
-64 9.27 0.65 1845 89 979 64 0.92 0.00 1817 22 1019
-128 14.20 2.55 2096 100 983 128 0.92 0.01 2031 26 1022
-256 30.01 9.08 2409 149 1030 256 1.79 0.74 2367 50 1068
-512 69.18 5.72 2842 193 1136 512 16.68 6.66 2823 47 1178
-1024 88.24 1.02 3379 64 1450 1024 47.96 5.14 3495 15 1492